### v1

# How Does A Text Preprocessing Pipeline Affect Ontology Syntactic Matching?

Abstract
The generic text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many ontology matching (OM) systems. However, the lack of standardisation in text preprocessing creates diversity in mapping results. In this paper, we investigate the effect of the text preprocessing pipeline on OM tasks at syntactic levels. Our experiments on 8 Ontology Alignment Evaluation Initiative (OAEI) track repositories with 49 distinct alignments indicate: (1) Tokenisation and Normalisation are currently more effective than Stop Words Removal and Stemming/Lemmatisation; and (2) The selection of Lemmatisation and Stemming is task-specific. We recommend standalone Lemmatisation or Stemming with post-hoc corrections. We find that (3) Porter Stemmer and Snowball Stemmer perform better than Lancaster Stemmer; and that (4) Part-of-Speech (POS) Tagging does not help Lemmatisation. To repair less effective Stop Words Removal and Stemming/Lemmatisation used in OM tasks, we propose a novel context-based pipeline repair approach that significantly improves matching correctness and overall matching performance. We also discuss the use of text preprocessing pipeline in the new era of large language models (LLMs).
Index Terms:
ontology matching, text preprocessing
I Introduction
Ontology matching (OM), also known as ontology alignment, is essential to enable interoperability between heterogeneous ontologies. An OM process usually takes two ontologies as input, discovers mappings between entities, and produces a set of correspondences [1]. A classical OM system usually contains syntactic, lexical, and semantic matching. This multilayer architecture has been implemented in several successful OM systems, such as LogMap [2][3], AgreementMakerLight (AML) [4][5], and FCA-Map [6][7].
Syntactic matching captures “anchor mappings”, providing the foundation for the latter lexical and semantic matching. There are many strategies to extract syntactic information from an ontology entity, including the older approach of Bag-of-Words (e.g. TF-IDF [8]), popular word embedding models (e.g. Word2Vec [9]), and state-of-the-art language models (e.g. BERT [10]). Despite the diversity of the models used, they all apply text preprocessing for cleaning the text data before fitting it into the model. Fig. 1 shows an example of using a classical text preprocessing pipeline to process the ontology entity “cmt:reviewerBiddingStartedBy”. The text preprocessing pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation [11]. Tokenisation is the process of breaking the text into the smallest units (i.e. tokens). We use whitespace to split the tokens in the example. Normalisation is the process of transforming these different tokens into a single canonical form. Stop Words Removal is the process of removing filler words that usually carry little meaning and can be omitted in most cases. Stemming/Lemmatisation is used to deal with the grammatical variation of words, applying rules to find the simplest common form of the word. This helps to capture the key information from the text and therefore improves efficiency.
Figure 1: Text preprocessing pipeline used in syntactic matching.
While a number of OM systems use the text preprocessing pipeline for syntactic OM, few studies explain why a specific method is selected for a certain OM task. Our study tackles the challenge in two ways. Firstly, we conduct a comprehensive experimental analysis of the text preprocessing pipeline in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline in OM tasks at syntactic levels. In each phase, a text preprocessing method is evaluated for its correctness and completeness. Secondly, we propose a novel context-based pipeline repair approach for syntactic OM. The method offers a customised way to fine-tune the text preprocessing pipeline for each domain-specific OM task and shows promising results for repairing false mappings. We also discuss the non-deterministic nature of using LLMs for text preprocessing and highlight the importance of utilising the classical text preprocessing pipeline. Specifically, this paper makes the following contributions:
  • (Section IV) We categorise the text preprocessing pipeline used in syntactic OM into two phases. We find a significant improvement using Phase 1 text preprocessing methods. In contrast, Phase 2 text preprocessing methods are currently less effective. We compare the performance of (1) Stemming and Lemmatisation, (2) different stemming methods (Porter, Snowball, and Lancaster), and (3) Lemmatisation and Lemmatisation + Part-of-Speech (POS) Tagging. We find that inappropriate stop words removal, over-stemming, and over-lemmatisation are common on 8 Ontology Alignment Evaluation Initiative (OAEI) [12] track repositories with 49 distinct alignments.
  • (Section V) We propose a simple and intuitive context-based pipeline repair method. This method is evaluated on the same OM tasks we analysed, showing promising results to improve the correctness of syntactic OM when inserted in the pipeline repair before Phase 2 text preprocessing methods.
  • (Section VI) We discuss the use of text preprocessing pipeline in the era of LLMs. We find LLMs cannot perform syntactic matching without text preprocessing, but they can be used as an alternative to repair the false mappings.
The remainder of the paper is organised as follows: Section II and Section III review the related work and the preliminaries of OM task formulation. Section IV analyses the text preprocessing pipeline used in OM. Section V proposes the context-based pipeline repair approach and experimentally validates its performance. Section VI demonstrates the pros and cons of using LLMs for text processing in OM. Section VII and Section VIII summarise the findings and limitations of this work. Section IX concludes the paper.
II Related Work
Syntactic matching considers only the meaning of the entity’s name or label, ignoring its lexical and structural context in an ontology [13]. Correct syntactic matches are often implicit and usually require extra human observation and domain knowledge. Text preprocessing is introduced to automate this process. The use of text preprocessing in syntactic matching can be traced back to the early stages of OM systems, having been initially developed and widely used as a basic component to generate linguistic-based mappings. AgreementMaker [14] has a normaliser to unify the textual information of the entities. SAMBO [15] uses the Porter Stemmer for each word to improve the similarity measure for terms with different prefixes and suffixes. In RiMOM [16], the context information of each entity is viewed as a document. The text in each document is preprocessed with tokenisation, stop words removal, and stemming. Recently, machine learning (ML) models have emerged for modern OM systems. While text preprocessing remains useful, its role is more focused on normalising the text that becomes the input to the model. For example, BERTMap [17] uses BERT’s inherent WordPiece tokeniser [18] to build the subword of each entity. A more recent approach DeepOnto [19] extends the normalisation to axioms using EL embedding models [20]. The ML extension of LogMap [21] reuses the seed mappings of the traditional system, where each entity is split into its component English word and the mapping is based on their similarity.
However, to the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have been conducted to evaluate the effect of text preprocessing on syntactic OM.
III Preliminaries: OM task formulation
Given a source ontology $O_{s}$ and a target ontology $O_{t}$, OM establishes mappings between pairs of entities drawn from each of two ontologies. A correspondence (i.e. one instance of mappings) is defined as a 4-tuple $(e_{1},e_{2},r,c)$, where $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. $r$ is the relationship between two matched entities $e_{1}$ and $e_{2}$, and $c\in[0,1]$ is the confidence for each correspondence. The relationship $r$ in OM tasks can be equivalence ($\equiv$), subsumption ($\subseteq$), disjointness ($\perp$), or other more complex relationships. In this paper, we focus only on the equivalence relationship ($\equiv$) and evaluate the effect of text preprocessing on syntactic matching to produce the “anchor mappings” on which to base any subsequent lexical and semantic matching. We address only equivalence because subsumption and disjointness are typically dealt with in a later semantic and structural matching phase of OM. An alignment (A) is a set of candidate correspondences generated by tools, while a Reference (R) is a set of gold standard correspondences verified by domain experts (i.e. the ground truth alignment).
IV Analysis of Text Preprocessing Pipeline
In this section, we illustrate the procedure used to analyse the text preprocessing pipeline in ontology syntactic matching and our evaluation results of the effect of different text preprocessing in the OAEI repositories.
IV-A Method
Fig. 2 shows the method used to analyse the text preprocessing pipeline. Alignment (A) is generated via the text preprocessing pipeline. Firstly, for both $O_{s}$ and $O_{t}$, we retrieve the entities from the named classes (i.e. owl:Class) and named properties (i.e. object properties owl:ObjectProperty and data type properties owl:DatatypeProperty). For those ontologies where the names of the concepts are not textual (e.g. a numerical identifier), instead, we retrieve the meaningful text from entity labels (i.e. rdfs:label). Then, we apply the text preprocessing pipeline method $f(.)$ on each entity $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. If $f(e_{1})=f(e_{2})$, we store the correspondence in the corresponding alignment file. Finally, we compare the generated Alignment (A) with Reference (R) to evaluate the performance of the text preprocessing pipeline on OM tasks.
Figure 2: Method used to analyse the text preprocessing pipeline. B-Base Entity without Text Preprocessing, T-Tokenisation, N-Normalisation, R-Stop Words Removal, S/L-Stemming/Lemmatisation.
  1. (1)
    Selected OAEI track repositories: The Ontology Matching Evaluation Toolkit (MELT) [22] is a powerful framework for OM evaluation. We retrieve 17 OAEI normal track repositories stored in the MELT public repository. The repositories are categorised as schema matching or instance matching. 12 of the 17 track repositories are schema-matching and applicable for evaluating OM. Only 8 of these track repositories are selected because the other 4 have noisy data or miss required files. Specifically, the knowledgegraph repository contains both schema and instance matching. The complex repository is primarily focused on detecting complex correspondences, while the multifarm repository is an extension of the conference repository to multilingualism. The laboratory repository lacks reference files at the time of writing. Table I shows the details of the selected track repositories. We consider only the equivalence mappings contained in the reference files. Each track repository may contain more than one alignment corresponding to different pairs of ontologies. For example, the largebio track repository has 6 reference files, pairing the whole and small versions of FMA [23] and NCI [24], FMA and SNOMED [25], and SNOMED and NCI, respectively. The number of references for each track repository is given in the table. There are 49 distinct alignments evaluated in this study.
    Name Domain Number of References anatomy Human and Mouse Anatomy 1 biodiv Biodiversity and Ecology 9 commonkg Common Knowledge Graphs 3 conference Conference 24 food Food Nutritional Composition 1 largebio Biomedical 6 mse Materials Science & Engineering 3 phenotype Disease and Phenotype 2
    TABLE I: Selected OAEI track repositories.
    Compound words are frequently used in ontology naming conventions. For example, compound words “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO [26] or the Snake case “art_gallery” used in Wikidata [27]. Fig. 3 shows the proportion of compound words in the selected OAEI track repositories. For comparisons between entities with compound words, one approach is to use their tokens, where “XY” is equivalent to “YX” because they share the same concepts X and Y. However, our comparison considers both concepts and their order, so that “XY” is not equivalent to “YX”. Although concurrent “XY” and “YX” are rare in the matching process, we assume that it could happen and so we take account of the order of concepts in compound word names to ensure a fair comparison in our experiments.
    Figure 3: The proportion of compound words.
  2. (2)
    Selected subtasks of pipeline methods: There are a variety of subtasks that can be used in a general text preprocessing pipeline, but not all of them are applicable to OM tasks. We select the following subtasks in each text preprocessing method: (1) Tokenisation: includes word tokenisation only. We do not use sentence tokenisation (also known as segmentation) because text retrieved from the entity’s name or label is commonly short. (2) Normalisation: includes lowercasing, HTML tags removal, separator formatting, and punctuation removal. Other subtasks that may potentially change the word semantics, such as removal of special characters and numbers, are excluded. (3) Stop Words Removal: includes the most common English stop words defined in the Natural Language Toolkit (NLTK) [28]. (4) Stemming/Lemmatisation: Stemming methods include Porter Stemmer, Snowball Stemmer, and Lancaster Stemmer. Lemmatisation uses the NLTK Lemmatiser [28] based on WordNet [29], and the word categorisation uses POS Tagging.
  3. (3)
    Selected OM evaluation measures: In information retrieval, there are four primitive measures: true positive (TP), false positive (FP), false negative (FN), and true negative (TN). In the context of OM, evaluation compares an alignment (A) returned by the OM system with a gold standard reference (R). Fig. 4 illustrates that the four primitive measures in OM can be interpreted as $TP=A\cap R$, $FP=A-R$, $FN=R-A$, and $TN=(C\times C^{\prime})-(A\cup R)$, where $C\times C^{\prime}$ refers to all possible correspondences $\in\{O_{s},O_{t}\}$.
    Figure 4: OM evaluation measures [30].
    Accuracy (Acc), Specificity (Spec), Precision (Prec), Recall (Rec), and $F_{\beta}$ Score ($F_{\beta}$) are the most common evaluation measures based on TP, FP, FN, and TN. In the context of OM, since $C\times C^{\prime}$ is extremely large (the Cartesian product of $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$), TN is often much larger than TP, FP, and FN. This means that Accuracy (Acc) and Specificity (Spec) are close to 1, and they have no statistically significant difference across different alignments. We note that Precision (Prec) and Recall (Rec) contribute equally to $F_{\beta}$. Therefore, we choose Precision (Prec), Recall (Rec), and F1 Score ($\beta=1$) in this study. They are defined as:
    $\footnotesize Prec=\frac{|A\cap R|}{|A|}\quad Rec=\frac{|A\cap R|}{|R|}\quad F_{1}=\frac{2}{Prec^{-1}+Rec^{-1}}$ (1)
IV-B Results
  1. (1)
    Comparison of pipeline methods: Fig. 5 summarises the comparison of the text preprocessing methods Tokenization (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L). The result indicates that the vast majority of correct correspondences are found by T and N. We do not see R and S/L playing a prime role in the OM tasks.
    Figure 5: Comparison of the generic text preprocessing pipeline: Base Entity without Text Preprocessing (B), Tokenisation (T), Normalisation (N), Stop Words Removal (R), Stemming/Lemmatisation (S/L). The methods are always applied sequentially in the pipeline. Three horizontal lines inside each violin plot show three quartiles: Q1, median, and Q3. The extension of the curves beyond 0% and 100% is an artefact of the seaborn [31] violin plot. (a) Precision: The median increases with T and N but decreases with R and S/L. After T, the shape of the distribution is unchanged by R and S/L. (b) Recall: After T, the median increases slightly with each of N, R, and S/L. The shape of the distribution does not change after N. (c) F1 Score: the median increases with T and N but then decreases with R and S/L. The shape of the distribution does not change after N.
  2. (2)
    Stemming vs. Lemmatisation: Fig. 6 compares Stemming (S) and Lemmatisation (L) in 49 alignments after Tokenisation (T) and Normalisation (N) have been applied. L is commonly better than S.
    Figure 6: Comparison of Stemming (S) and Lemmatisation (L). (a) Precision: The median after L is greater than that achieved by S. The shape of the distribution is slightly different. (b) Recall: The median and the shape of the distribution are identical after S and after L. (c) F1 Score: The median after L is greater than for S. The shape of the distribution is identical.
  3. (3)
    Porter Stemmer vs. Snowball Stemmer vs. Lancaster Stemmer: Fig. 7 compares Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL) in 49 alignments. SP and SS have been found to perform better than SL.
    Figure 7: Comparison of different stemmers: Porter Stemmer (SP), Snowball Stemmer (SS) and Lancaster Stemmer (SL). (a) Precision: The median number in SP and SS is greater than that in SL, and there is no difference between SP and SS. The shape of the distribution is identical. (b) Recall: The median number and the shape of the distribution are identical in SP, SS, and SL. (c) F1 Score: The median number in SP and SS is greater than that in SL, and there is no difference between SP and SS. The shape of the distribution is identical.
  4. (4)
    Lemmatisation vs Lemmatisation + POS Tagging: Fig. 8 summarises the comparison of Lemmatisation (L) and Lemmatisation + POS Tagging (LT) in 49 alignments. The result indicates that POS Tagging does not help with Lemmatisation in Precision, Recall, and overall F1 Score in the total of 49 alignments we analysed.
    Figure 8: Comparison of Lemmatisation (L) and Lemmatisation + POS Tagging (LT). For each of (a) Precision, (b) Recall and (c) F1 Score, the median and the shape of the distribution are identical for L and LT.
IV-C Discussion
For syntactic OM, the text preprocessing pipeline can be categorised into two phases. Phase 1 text preprocessing pipeline methods contain Normalisation and Tokenisation, whereas Stop Words Removal and Stemming/Lemmatisation are assigned to Phase 2. Phase 1 text preprocessing pipeline methods do not change word semantics; instead, they only change syntactic features such as formatting and typography. Conversely, in Phase 2 of the text preprocessing pipeline, words are changed to better capture semantic similarity, such as removing prefixes and suffixes or substituting common etymological roots.
  1. (1)
    Phase 1 text preprocessing methods: We observe that matching performance usually increases with each Phase 1 text preprocessing method, indicating the benefit of these “rules of thumb” for syntactic OM. If two terms can be matched using a heuristic or intuitive technique, there is no need to leverage the more complex Phase 2 methods. However, Phase 1 methods could also benefit from customisation in OM tasks. Some traditional methods originating for natural language processing (NLP) are not useful for OM and need to be adjusted. For example, sentence segmentation (i.e.splitting long text into sentences) is not applicable for ontology entities because they are generally short text fragments and do not need such operations. Word tokenisation (i.e. breaking text into single words) could be rewritten to detect the abbreviations that are common practice in ontology concept names and to tackle the conflicting use of camel case and snake case.
  2. (2)
    Phase 2 text preprocessing methods: We observe no benefit from Phase 2 text preprocessing methods. In some cases, Stop Words Removal and Stemming/Lemmatisation may even hamper the mapping. There are plausible explanations for this behaviour arising from the nature of ontologies as distinct from natural language text, as follows:
    1. (a)
      Stop Words Removal: “AND”, “OR”, and “NOT” are stop words in English and usually carry little useful information, whereas these words express logical operations in ontology entities and therefore may carry important semantics.
    2. (b)
      Stemming vs. Lemmatisation: Based on our experiments, Lemmatisation is better than Stemming. While lemmatisation tends to avoid generating FPs, it may also miss some implicit TPs. The complexity of finding a missing TP is equal to the size of $(C^{\prime})$. On the other hand, stemming is more aggressive in finding TPs, but the aggression can lead to more FPs as well. The complexity of finding FPs is equal to the smaller size of $(A)$. The workload of discovering FPs after stemming is generally much lighter than detecting missing TPs after lemmatisation, but this may also depend on the accuracy of post-hoc corrections that can be performed.
    3. (c)
      Porter Stemmer vs. Snowball Stemmer vs. Lancaster Stemmer: Porter Stemmer and Snowball Stemmer (also known as Porter 2 Stemmer) have been found to perform better than Lancaster Stemmer, and we cannot see a significant difference between Porter and Snowball. Although The Lancaster Stemmer is more aggressive in detecting more TPs, it does not lead to performance improvement as more FPs are matched.
    4. (d)
      Lemmatisation vs. Lemmatisation + POS Tagging: Lemmatisation + POS Tagging is generally expected to have better results than Lemmatisation alone because tagging can help detect more precise root words. However, we do not observe such a performance improvement when using POS Tagging in our study. The reason may be that ontology classes are usually nouns or gerunds, and in such cases, we could expect the simpler grammatical assumption to have similar results.
V Context-based Pipeline Repair
Experimental results in the last section demonstrate that only Phase 1 text processing (i.e. Tokenisation and Normalisation) helps with syntactic OM. The use of Phase 2 text preprocessing (i.e. Stop Words Removal and Stemming/Lemmatisation) does not improve performance and may even have negative impacts. In this section, we explore the underlying cause of this phenomenon. Furthermore, we introduce an intuitive context-based pipeline repair approach and demonstrate its effectiveness on the same datasets we analysed.
V-A Motivation
  1. (1)
    Phase 1 text preprocessing methods (Tokenisation and Normalisation) do not change the text meaning. For example, isReviewing is equivalent to is_reviewing. This means that applying Phase 1 methods only helps detect TPs, while the number of FPs remains unchanged. For this reason, Phase 1 methods always have a positive effect on Precision, Recall, and overall F1 Score.
    $\footnotesize Prec\uparrow=\frac{|A\cap R|}{|A|}=\frac{TP}{TP+FP}=1-\frac{FP}{TP\uparrow+FP}$ (2)
    $\footnotesize Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$ (3)
    $\footnotesize F_{1}\uparrow=\frac{2}{Prec\uparrow^{-1}+Rec\uparrow^{-1}}$ (4)
    In most cases, OM only requires minor preprocessing using Phase 1 text preprocessing methods. This is because entity names in the ontology are often compound words that do not occur in natural language, but they are partially formalised by agreement. There have been well-defined conventions established over the years, e.g. singularity, positive names, nouns for classes and verbs for properties [32] [33].
  2. (2)
    Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are actually relaxations of matching rules in OM tasks. Moving through the text preprocessing pipeline tends to detect more TPs and FPs in the derived alignment (A). For example, isReviewing and isReviewedBy may be object properties with distinctly different meaning, but removing the common stop words “is” and “by”, and using Stemming/Lemmatisation to retrieve the same root word “Review”, could cause a false match. For this reason, Recall is always increasing, but Precision and overall F1 Score are less reliable.
    $\footnotesize Prec\ ?=\frac{|A\cap R|}{|A|}=\frac{TP\uparrow}{TP\uparrow+FP\uparrow}$ (5)
    $\footnotesize Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$ (6)
    $\footnotesize F_{1}\ ?=\frac{2}{Prec\ ?^{-1}+Rec\uparrow^{-1}}$ (7)
    If we define $\Delta TP$ and $\Delta FP$ as the increase in TP and FP for a preprocessing method, then the threshold to increase Precision and F1 Score is:
    $\scriptsize\frac{TP+\Delta TP}{TP+\Delta TP+FP+\Delta FP}>\frac{TP}{TP+FP}\Rightarrow\frac{\Delta TP}{\Delta FP}>\frac{TP}{FP}$ (8)
    In our experiments, we actually observe a reduction in Precision and overall F1 Score. This means that Phase 2 methods produce more FPs than TPs, and this proportion is less than the original number of $TP/FP$. So performance does not improve, unless the benefit of each TP is considered more valuable than the disbenefit of each FP. This could apply, for example, if we are expecting a post-hoc correction phase where removing FPs is considered to be an easier human task than adding missing TPs.
V-B Method
Phase 2 text preprocessing pipeline methods (Stop Words Removal and Stemming/Lemmatisation) have been shown to be less effective in OM tasks, caused mainly by FPs. We propose a pipeline repair approach that aims to differentiate FPs and therefore improve Precision and F1 Score. As illustrated in Fig. 9: (1) Phase 1 methods shift Alignment (A) towards Reference (R). The number of TPs increases, while FPs decrease. (2) Phase 2 methods expand Alignment (A). The number of TPs increases, but FPs also increase. (3) Our context-based pipeline repair approach collapses Alignment (A). The number of FPs significantly decreases, with a slight decrease in TPs.
Figure 9: Mechanism of the context-based pipeline repair.
One critical step in our approach is to retrieve a reserved word set that may cause FPs after the text preprocessing, and these words will be excluded before the text preprocessing. The selection criteria are based on two widely agreed assumptions: (1) there are no duplicate entities within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies (we call our approach context-based here because pairs of words may have the same or different meanings in different contexts). Based on these two assumptions, we propose a simple and intuitive Algorithm 1 to retrieve the reserved word set for context-based pipeline repair. For both $O_{s}$ and $O_{t}$, the algorithm first iterates on all pairs of entities $e_{i},e_{j}$ in each of them. For a specific text preprocessing method $f(.)$, if $f(e_{i})=f(e_{j})$, we retrieve the different words between $e_{i}$ and $e_{j}$ and store them in the reserved word set. If a word appears in the reserved set, the text preprocessing pipeline skips the operation for this word. To simplify the reserved word set, we also remove the words where $f(w)=w$ from the final set because either skipping or keeping these words in the reserved word set would not change the mapping results.
Algorithm 1 Finding the reserved words
0:  Source Ontology $O_{s}$, Target Ontology $O_{t}$, Text Preprocessing Pipeline Method $f(x)$
0:  $Reserved\_Word\_Set$
  /* Phase 1: Find duplicates in $O_{s}$ and $O_{t}$/
  /* Phase 1.1: Find duplicates in $O_{s}$/
  for $Entity\ e_{i},e_{j}\in O_{s}$ do
     if $f(e_{i})=f(e_{j})$ then
        $Reserved\_Word\_Set\leftarrow differ(e_{i},e_{j})$
     end if
  end for
  /* Phase 1.2: Find duplicates in $O_{t}$/
  /* Same procedure applies …*/
  /* Phase 2: Find duplicates in $Reserved\_Word\_Set$/
  for $Word\ w\in Reserved\_Word\_Set$ do
     if $f(w)=w$ then
        $Reserved\_Word\_Set\rightarrow w$
     end if
  end for
  return  $Reserved\_Word\_Set$
An example of generating and using a simple reserved word set is illustrated below. (1) Two object properties was_a_member_of and has_members from a single ontology have the same result “member” via the traditional text preprocessing. Because there are no duplicate entities within a single ontology, we use a reserved word set in our proposed pipeline repair approach to determine that they are distinguishing entities. The initial step (i.e. Phase 1 of Algorithm 1) is to add [“was”, “a”, “member”, “of”, “has”, “members”] to the reserved word set so that these two object properties would not have the same text preprocessing results. was_a_member_of preprocessed with skipping the reserved word set is “was a member of”, while has_members preprocessed with the reserved word set is “has members”. The revision step (i.e. Phase 2 of Algorithm 1) is to check whether there are duplicates in the reserved set. We can observe that the word “member” is a duplicate because it is the same before and after text preprocessing. Removing this word from the reserved word set still makes the two object properties different. Therefore, the final reserved word set is [“was”, “a”, “of”, “has”, “members”]. (2) The generated reserved word set can be used to repair false mappings between entities within the same domain context but coming from different ontologies. For example, we expect that the two object properties has_a_steering_committee and was_a_steering_committee_of are non-identical. While a false mapping may occur when they both have the same result “steer committe” after the traditional text preprocessing, using the reserved word set can repair this false mapping. As the words “has”, “a”, “was”, and “of” are listed as reserved words, these two named properties are preprocessed as “has a steer committe” and “was a steer committe of”, respectively.
V-C Evaluation
We apply our context-based pipeline repair approach to the same OAEI track repositories and alignments as above. Fig. 10 compares with and without context-based pipeline repair in 8 track repositories with 49 distinct alignments. The text preprocessing pipeline methods implemented include all the Phase 2 text preprocessing methods: Stop Words Removal (R), Porter Stemmer (SP), Snowball Stemmer (SS), Lancaster Stemmer (SL), Lemmatisation (L), and Lemmatisation + POS Tagging (LT). We can see that our context-based repair approach can significantly improve the Precision. As a trade-off, it may cause a slight decrease in Recall, but the overall F1 Score is still increasing in the majority of the alignments. For example, this repair approach applied to the MaterialInformation-EMMO alignment using Lancaster Stemmer improved Precision and overall F1 score by 8.95% and 2.92%, with only a decrease of 3.17% in Recall.
Figure 10: Context-based pipeline repair use in Phase 2 text preprocessing methods (the total number of each category can appear to be less than 49 when data points overlap).
VI Text Preprocessing in the era of LLMs
With the recent advances in LLMs, it becomes questionable whether we still need the classic text processing pipeline with the availability of powerful LLMs. In this section, we demonstrate the non-deterministic nature of LLMs in OM text preprocessing and showcase the potential use of LLMs in OM text preprocessing pipeline repair.
VI-A LLMs Used in OM Text Preprocessing
We observe that some LLMs can perform syntactic matching without preprocessing, but their robustness remains questionable. We conduct a study on the use of LLMs for syntactic matching with and without text preprocessing. The experiment settings are described as follows:
  1. (1)
    We select the same example described in Section IV-A. The compound word “art gallery” can have different naming conventions in different ontologies. For example, the Pascal case “ArtGallery” used in YAGO and the Snake case “art_gallery” used in Wikidata.
  2. (2)
    We choose 10 LLMs from 4 different families. These include two OpenAI models (gpt-3.5-turbo and gpt-4-turbo) [34], three Anthropic Claude models (claude-3-opus, claude-3-sonnet, and claude-3-haiku) [35], three Mistral AI models (mistral-large, mistral-medium, and mistral-small) [36], and two open-source Meta Llama models (llama-3-70b and llama-3-8b) [37]. We use LangChain [38] to build the chats with OpenAI, Mistral AI, and Anthropic Claude models, and Ollama [39] to access Meta Llama models. The versions of LLMs used in the experiment are listed in Table II. All the model temperatures are set to 0 to minimise the random results generated by LLMs.
    Series LLMs Versions GPT gpt-4-turbo gpt-4-turbo-2024-04-09 gpt-3.5-turbo gpt-3.5-turbo-0125 Claude claude-3-opus claude-3-opus-20240229 claude-3-sonnet claude-3-sonnet-20240229 claude-3-haiku claude-3-haiku-20240307 Mistral mistral-large mistral-large-2402 mistral-medium mistral-medium-2312 mistral-small mistral-small-2402 Llama llama-3-70b date accessed: 2024-06-01 llama-3-8b date accessed: 2024-06-01
    TABLE II: Versions of LLMs used in the experiment.
  3. (3)
    We generate the prompt template as “Is ArtGallery <KEYWORD> to/with art_galley?” (without text preprocessing) and “Is art galley <KEYWORD> to/with art galley?” (with text preprocessing). We use only T and N to preprocess the text because R and S/L are shown to be less effective in the main content of the paper. Furthermore, R and S/L needs are not evident in the original naming of the ontology entities. The <KEYWORD> is a placeholder for a collection of words that can be used to describe an equivalence relationship. In this study, we experiment with three common words namely “identical”, “interchangeable”, and “equivalent”.
  4. (4)
    Considering the complexity of prompts, we also test an additional case where the prompts have a self-reflection phase (i.e. add “Write a short explanation” to the prompt). We use parentheses to indicate that the different responses generated by the prompts have a self-reflection phase.
Table III shows the results of using LLMs for syntactic matching without text preprocessing. We expect the LLMs to generate “Yes” answers for this TP sample. However, the majority of the LLMs output the “No” answers (marked with the red colour in the table) without text preprocessing.
LLMs <KEYWORD> “identical” “interchangeable” “equivalent” gpt-4-turbo No No No gpt-3.5-turbo No No No claude-3-opus No Yes Yes claude-3-sonnet No No No claude-3-haiku No No No mistral-large No No Yes mistral-medium No No No mistral-small No Yes Yes (No) llama-3-70b No No No llama-3-8b Yes Yes Yes
TABLE III: Using LLMs for syntactic matching without text preprocessing. Prompt: “Is ArtGallery <KEYWORD> to/with art_gallery? Answer yes or no. (Write a short explanation.)”
Table IV shows the results of using LLMs for syntactic matching with text preprocessing. We can see that the error rate can be significantly reduced, but some LLMs still continue to produce incorrect results. The result also shows that LLMs with a larger number of parameters (e.g. llama-3-70b) and prompts with a self-reflection phase (labelled in parentheses) offer “too much of a good thing”. They do not help with syntactic matching and may even produce more errors in OM tasks. Although this preliminary study does not cover all possible triggers used in prompt engineering, it is adequate to highlight the importance of proper preprocessing before fitting the text into the LLMs. We believe that the classical programming-based text preprocessing pipeline is more stable and reliable than using LLM’s implicit text preprocessing or prompt-based text preprocessing for OM tasks.
LLMs <KEYWORD> “identical” “interchangeable” “equivalent” gpt-4-turbo Yes Yes Yes gpt-3.5-turbo Yes Yes Yes claude-3-opus Yes Yes Yes claude-3-sonnet Yes (No) No Yes (No) claude-3-haiku No No Yes (No) mistral-large Yes Yes Yes mistral-medium No Yes Yes mistral-small Yes Yes Yes llama-3-70b Yes Yes (No) Yes (No) llama-3-8b Yes Yes Yes
TABLE IV: Using LLMs for syntactic matching with text preprocessing. Prompt: “Is art gallery <KEYWORD> to/with art gallery? Answer yes or no. (Write a short explanation.)”
VI-B LLMs Used in OM Text Preprocessing Pipeline Repair
Using LLMs for OM text preprocessing is not without merit. With strong background knowledge, they can be used to facilitate several subtasks in OM. For example, LLMs can be used as a repair tool for the text preprocessing pipeline, as an alternative to the approach we proposed in Section V-B. The slight difference is that the context-based pipeline repair approach is an ad hoc repair prior to syntactic matching, while the LLM-based approach is a post hoc repair after syntactic matching. We conduct a preliminary study on LLMs used in OM text preprocessing pipeline repair. The experiment settings are described as follows:
  1. (1)
    The experiment is set up on the largebio track repository, a biomedical track repository requires higher Precision with low rates of FPs. The Lancaster Stemmer always shows lower Precision across 6 pairs of alignments, meaning that the results contain a significant number of FPs.
  2. (2)
    We choose the same 10 LLMs used in the previous experiment in Appendix VI-A.
  3. (3)
    We use the prompt “Is X equivalent to Y?” and preprocess the entity names X and Y with N and T before fitting them into the LLMs. The previous experiment in Appendix VI-A shows these two settings are more suitable for OM tasks.
  4. (4)
    Considering the complexity of prompts, we also test an additional case where the prompts have a self-reflection phase (i.e. add “Write a short explanation” to the prompt).
Figure 11: The discovery rate of FPs using LLMs on the largebio track repository.
Fig. 11 shows the discovery rate of FPs using 10 different LLMs. We expect the LLMs to generate “No” answers for these FP samples. The GPT model gpt-4-tubo, the Claude model claude-3-haiku, the Mistral model mistral-medium, and the Llama model llama-3-70b achieve better performance in detecting FPs. For the GPT and Llama series, LLMs with a larger number of parameters perform better than those with a relatively smaller number of parameters. For the Claude and Mistral series, LLMs with larger parameters do not improve performance. In the test cases we analysed, they are even less effective than those with a relatively smaller number of parameters. Unlike detecting TPs, we can see a self-reflection phase improves the overall performance of detecting FPs.
VI-C Implications
LLMs were originally developed for the purpose of question-answering (QA). OM is not only a QA task, but also a knowledge-intensive task that relates to a comprehensive understanding and reasoning of domain knowledge. The studies in this section demonstrate the limitations and opportunities of using LLMs for OM tasks. From our experience, we highlight two pathways to customise LLM for a new domain task. One pathway is function calling (FC), which refers to teaching LLMs using external tools. In LLM-based text preprocessing, we could package the programming-based text preprocessing pipeline as a tool for LLMs to use. The other pathway is fine-tuning (FT), where LLMs are customised with local data. In LLM-based text preprocessing repair, we could use human-in-the-loop to validate the repaired results generated by LLMs and feed the validated data back into LLMs, so that they can better understand the context of the task.
VII Discussion
Ontologies provide formalised conceptualisations of knowledge graphs (KGs). Incorporating ontologies for KG-related NLP tasks can enhance the capability to handle complex concepts and relations, thereby improving the quality of KGs built from text. Ontologies also assist interoperability of text generated from KGs. However, aligning and integrating heterogeneous ontologies remains a challenge when using them for NLP tasks. OM aims to bridge the gap by capturing similar concepts that occur in different ontologies.
OM is a complex process that combines syntactic, lexical, and semantic matching. Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching. This paper concentrates on text preprocessing, a common practice in syntactic matching. While generic text preprocessing pipeline methods in syntactic OM are usually applied on the basis of intuition or extrapolation from other experience, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
  1. (1)
    Our study shows (i) whether to use, or not use, and (ii) how to use these text preprocessing methods. Such experimental results will benefit decision making when selecting appropriate text preprocessing methods for syntactic OM. For large-scale OM, it can significantly reduce unnecessary trial costs.
  2. (2)
    Our context-based pipeline repair approach is proposed to repair the less effective Phase 2 text preprocessing methods. Its broader value is showing how to maximise true mappings and minimise false mappings throughout the text preprocessing pipeline. From a statistical perspective, it is a local optimisation for syntactic matching that will benefit the global optimisation for OM, as syntactic matching provides “anchor mappings” for lexical and semantic matching.
  3. (3)
    The use of LLMs is not a one-size-fits-all solution for OM tasks. OM is a complex multi-level matching mixed with syntactic, lexical, and semantic matching. LLMs may not outperform traditional methods in performing these sub-matching tasks. The development and implementation of an LLM-based framework that incorporates different matching techniques is a new research direction for OM, but it is beyond the scope of this paper. We refer readers to several recent works in [40, 41, 42, 43, 44].
This study also demonstrates that the nature of text preprocessing used in NLP applications can be context-based. Not all pipeline methods are effective, and some of them may even hamper the overall performance of downstream tasks. Modifications are required before applying text preprocessing to specific NLP tasks.
VIII Limitations
OAEI track repositories are comprehensive but do not cover all real-world scenarios. Furthermore, the reference files contained in the OAEI track repositories are not “complete” gold standards and some need further development. In this study, we only deal with equivalence mappings between named classes, object properties, and data type properties. Disjointness mappings are not considered. Finding widely used baseline data with ground truth matches for subsumptions is empirically difficult. Only a few OAEI track repositories contain subsumption mappings, and they are not enough to robustly demonstrate text preprocessing used in subsumption mappings. This work has not been entered into the official OAEI campaign because of its intense focus on only the preprocessing step.
IX Conclusion
In this paper, we conduct a comprehensive study on the effect of the text preprocessing pipeline on syntactic OM. 8 OAEI track repositories with 49 distinct alignments are evaluated. Despite the importance of text preprocessing in syntactic OM, our experimental results indicate that the text preprocessing pipeline is currently ill-equipped to handle OM tasks. (1) We find that Phase 1 text preprocessing methods (Tokenisation and Normalisation) help with both matching completeness (i.e. Recall) and correctness (i.e. Precision). Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are less effective. They can improve matching completeness (i.e. Recall), but matching correctness (i.e. Precision) is relatively low. (2) We propose a novel context-based pipeline repair approach to repair the less effective Phase 2 text preprocessing methods. By using a reserved word set to reduce false positive samples detected, our approach outperforms the traditional text preprocessing pipeline, in particular, the matching correctness (i.e. Precision) and overall matching performance (i.e. F1 Score).
Our future work will focus on handling class axioms and complex relationships to evaluate the text preprocessing pipeline for OM tasks. We will also study the pipeline and pipeline repair approach working with both knowledge-based OM systems and LLM-based OM systems.
Supplemental Material Statement
The source code for this study is available from https://github.com/qzc438/ontology-nlp. The OAEI track repositories used in this study can be downloaded from https://dwslab.github.io/melt/track-repository (date accessed: 2024-06-01). The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.
Acknowledgment
The authors would like to thank Alice Richardson of the Statistical Support Network, Australian National University (ANU), for helpful advice on the statistical analysis in this paper. The authors also thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.

### v2

@@
 The method offers a customised way to fine-tune the text preprocessing pipeline for each domain-specific OM task and shows promising results for repairing false mappings.
-We also discuss the non-deterministic nature of using LLMs for text preprocessing and highlight the importance of utilising the classical text preprocessing pipeline.
+Thirdly, we discuss the non-deterministic nature of LLMs for text preprocessing and highlight the importance of retaining the classical text preprocessing pipeline.
 Specifically, this paper makes the following contributions:
@@
 the ground truth alignment).
-IV Analysis of Text Preprocessing Pipeline
-In this section, we illustrate the procedure used to analyse the text preprocessing pipeline in ontology syntactic matching and our evaluation results of the effect of different text preprocessing in the OAEI repositories.
+IV Text Preprocessing Pipeline: Analysis
+In this section, we discuss the role of the text preprocessing pipeline in ontology syntactic matching and report our experiments studying the effect of different text preprocessing options over the OAEI repositories.
 IV-A Method
@@
 V Context-based Pipeline Repair
-Experimental results in the last section demonstrate that only Phase 1 text processing (i.e.
+Experimental results in Section IV demonstrate that only Phase 1 text processing (i.e.
 Tokenisation and Normalisation) helps with syntactic OM.
@@
 Stop Words Removal and Stemming/Lemmatisation) does not improve performance and may even have negative impacts.
-In this section, we explore the underlying cause of this phenomenon.
-Furthermore, we introduce an intuitive context-based pipeline repair approach and demonstrate its effectiveness on the same datasets we analysed.
-V-A Motivation
+In this section, we explore the underlying cause of this behaviour.
+Furthermore, we introduce an intuitive context-based pipeline repair approach and demonstrate its effectiveness on the same datasets.
+V-A Causal analysis
   1. (1)
@@
 This could apply, for example, if we are expecting a post-hoc correction phase where removing FPs is considered to be an easier human task than adding missing TPs.
-V-B Method
+V-B Repair Method
 Phase 2 text preprocessing pipeline methods (Stop Words Removal and Stemming/Lemmatisation) have been shown to be less effective in OM tasks, caused mainly by FPs.
@@
 VI Text Preprocessing in the era of LLMs
-With the recent advances in LLMs, it becomes questionable whether we still need the classic text processing pipeline with the availability of powerful LLMs.
-In this section, we demonstrate the non-deterministic nature of LLMs in OM text preprocessing and showcase the potential use of LLMs in OM text preprocessing pipeline repair.
+With the recent advances in LLMs, we ask whether we still need the classic text processing pipeline.
+In this section, we demonstrate the non-deterministic nature of LLMs in OM text preprocessing and showcase the benefits of LLMs in a repair function inside the OM text preprocessing pipeline.
 VI-A LLMs Used in OM Text Preprocessing
@@
 The source code for this study is available from https://github.com/qzc438/ontology-nlp.
-The OAEI track repositories used in this study can be downloaded from https://dwslab.github.io/melt/track-repository (date accessed: 2024-06-01).
+The OAEI track repositories used in this study can be downloaded from OAEI MELT at https://dwslab.github.io/melt/track-repository (date accessed: 2024-06-01).
 The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.
@@
 The authors also thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.
+Appendix A Supplementary Material
+Appendix B Extended Experiment Details
+B-A Analysis of Text Preprocessing Pipeline
+Fig. 121314, and 15 show the details of the experiment to compare the text preprocessing pipeline in syntactic OM.
+Figure 12: Base without Preprocessing (B) vs. Tokenisation (T).
+Figure 13: Tokenisation (T) vs. Normalisation (N).
+Figure 14: Normalisation (N) vs. Stop Words Removal (R).
+Figure 15: Stop Words Removal (R) vs. Stemming/Lemmatisation (S/L).
+  1. (1)
+    For Phase 1 text preprocessing methods (Tokenisation and Normalisation), most of the data points located above the equivalent line in Precision, Recall, and F1 Score indicate that they help syntactic OM.
+  2. (2)
+    For Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation): Some data points are located above the equivalent line in Recall, but the majority of data points located below the equivalent line in Precision and F1 Score indicate that they do not help syntactic OM.
+Fig.
+16 shows the details of the experiment to compare Stemming (S) and Lemmatisation (L).
+Most of the data points located above the L=S line in Precision and F1 Score indicate that using Lemmatisation is better than Stemming in syntactic OM (assuming that the post hoc correction is excluded).
+Figure 16: Stemming (S) vs. Lemmatisation (L).
+Fig. 1718, and 19 show the details of the experiment to compare different stemmers.
+Figure 17: Porter Stemmer (SP) vs. Snowball Stemmer (SS).
+Figure 18: Porter Stemmer (SP) vs. Lancaster Stemmer (SL).
+Figure 19: Snowball Stemmer (SS) vs. Lancaster Stemmer (SL).
+  1. (1)
+    For Porter Stemmer vs. Snowball Stemmer, all data points located in the equivalent line indicate that there is no difference using Porter Stemmer and Snowball Stemmer in syntactic OM.
+  2. (2)
+    For Porter/Snowball Stemmer vs.
+Lancaster Stemmer, most of the data points the data points located above the equivalent line in Precision and F1 Score indicate that Porter/Snowball Stemmer is more effective than Lancaster Stemmer in syntactic OM.
+Fig 20 shows the details of the experiment to compare Lemmatisation vs.
+Lemmatisation + POS Tagging.
+All data points located in the equivalent line indicate that using POS Tagging in Lemmatisation does not help syntactic OM.
+Figure 20: Lemmatisation (L) vs. Lemmatisation + Part-Of-Speech Tagging (LT).
+B-B Context-based Pipeline Repair
+Fig.
+2122232425, and 26 show the details of the experiment to consider the benefit of using context-based pipeline repair in Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation).
+  1. (1)
+    For Precision, most of the data points are located above the equivalent line, indicating that context-based pipeline repair significantly improves matching correctness.
+  2. (2)
+    For Recall, most of the data points are located in the equivalent line, and only a few data points are located below the equivalent line, indicating that context-based pipeline repair slightly reduces matching completeness.
+  3. (3)
+    For F1 Score, most of the data points are located above the equivalent line, indicating that context-based pipeline repair also improves overall matching performance.
+  4. (4)
+    Experimentally, the wider ellipse around the data points indicates that the matching performance improvement ranking in Phase 2 text preprocessing methods is Stemming (S) > Lemmatisation (L) > Stop Words Removal (R).
+  5. (5)
+    Experimentally, the wider ellipse around the data points indicates that the matching performance improvement ranking in different stemmers is Lancaster Stemmer (SL) > Porter Stemmer (SP) = Snowball Stemmer (SS).
+  6. (6)
+    Experimentally, the same ellipse around the data points indicates that the matching performance improvement ranking in lemmatisation is Lemmatisation (L) = Lemmatisation + POS Tagging (LT).
+Figure 21: Comparison of using and without using context-based repair in Stop Words Removal (R).
+Figure 22: Comparison of using and without using context-based repair in Porter Stemmer (SP).
+Figure 23: Comparison of using and without using context-based repair in Snowball Stemmer (SS).
+Figure 24: Comparison of using and without using context-based repair in Lancaster Stemmer (SL).
+Figure 25: Comparison of using and without using context-based repair in Lemmatisation (L).
+Figure 26: Comparison of using and without using context-based repair in Lemmatisation + POS Tagging (LT).

### v3

@@
 F1 Score).
+(3) We also argue the traditional text preprocessing pipeline is not obsolete in the era of LLMs, but an integral part for LLMs used in the OM tasks.
 Our future work will focus on handling class axioms and complex relationships to evaluate the text preprocessing pipeline for OM tasks.
@@
 Supplemental Material Statement
-The source code for this study is available from https://github.com/qzc438/ontology-nlp.
+The source code for this study is available from https://github.com/qzc438-research/ontology-nlp.
 The OAEI track repositories used in this study can be downloaded from OAEI MELT at https://dwslab.github.io/melt/track-repository (date accessed: 2024-06-01).
@@
 The authors also thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.
-Appendix A Supplementary Material
-Appendix B Extended Experiment Details
-B-A Analysis of Text Preprocessing Pipeline
-Fig. 121314, and 15 show the details of the experiment to compare the text preprocessing pipeline in syntactic OM.
-Figure 12: Base without Preprocessing (B) vs. Tokenisation (T).
-Figure 13: Tokenisation (T) vs. Normalisation (N).
-Figure 14: Normalisation (N) vs. Stop Words Removal (R).
-Figure 15: Stop Words Removal (R) vs. Stemming/Lemmatisation (S/L).
-  1. (1)
-    For Phase 1 text preprocessing methods (Tokenisation and Normalisation), most of the data points located above the equivalent line in Precision, Recall, and F1 Score indicate that they help syntactic OM.
-  2. (2)
-    For Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation): Some data points are located above the equivalent line in Recall, but the majority of data points located below the equivalent line in Precision and F1 Score indicate that they do not help syntactic OM.
-Fig.
-16 shows the details of the experiment to compare Stemming (S) and Lemmatisation (L).
-Most of the data points located above the L=S line in Precision and F1 Score indicate that using Lemmatisation is better than Stemming in syntactic OM (assuming that the post hoc correction is excluded).
-Figure 16: Stemming (S) vs. Lemmatisation (L).
-Fig. 1718, and 19 show the details of the experiment to compare different stemmers.
-Figure 17: Porter Stemmer (SP) vs. Snowball Stemmer (SS).
-Figure 18: Porter Stemmer (SP) vs. Lancaster Stemmer (SL).
-Figure 19: Snowball Stemmer (SS) vs. Lancaster Stemmer (SL).
-  1. (1)
-    For Porter Stemmer vs. Snowball Stemmer, all data points located in the equivalent line indicate that there is no difference using Porter Stemmer and Snowball Stemmer in syntactic OM.
-  2. (2)
-    For Porter/Snowball Stemmer vs.
-Lancaster Stemmer, most of the data points the data points located above the equivalent line in Precision and F1 Score indicate that Porter/Snowball Stemmer is more effective than Lancaster Stemmer in syntactic OM.
-Fig 20 shows the details of the experiment to compare Lemmatisation vs.
-Lemmatisation + POS Tagging.
-All data points located in the equivalent line indicate that using POS Tagging in Lemmatisation does not help syntactic OM.
-Figure 20: Lemmatisation (L) vs. Lemmatisation + Part-Of-Speech Tagging (LT).
-B-B Context-based Pipeline Repair
-Fig.
-2122232425, and 26 show the details of the experiment to consider the benefit of using context-based pipeline repair in Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation).
-  1. (1)
-    For Precision, most of the data points are located above the equivalent line, indicating that context-based pipeline repair significantly improves matching correctness.
-  2. (2)
-    For Recall, most of the data points are located in the equivalent line, and only a few data points are located below the equivalent line, indicating that context-based pipeline repair slightly reduces matching completeness.
-  3. (3)
-    For F1 Score, most of the data points are located above the equivalent line, indicating that context-based pipeline repair also improves overall matching performance.
-  4. (4)
-    Experimentally, the wider ellipse around the data points indicates that the matching performance improvement ranking in Phase 2 text preprocessing methods is Stemming (S) > Lemmatisation (L) > Stop Words Removal (R).
-  5. (5)
-    Experimentally, the wider ellipse around the data points indicates that the matching performance improvement ranking in different stemmers is Lancaster Stemmer (SL) > Porter Stemmer (SP) = Snowball Stemmer (SS).
-  6. (6)
-    Experimentally, the same ellipse around the data points indicates that the matching performance improvement ranking in lemmatisation is Lemmatisation (L) = Lemmatisation + POS Tagging (LT).
-Figure 21: Comparison of using and without using context-based repair in Stop Words Removal (R).
-Figure 22: Comparison of using and without using context-based repair in Porter Stemmer (SP).
-Figure 23: Comparison of using and without using context-based repair in Snowball Stemmer (SS).
-Figure 24: Comparison of using and without using context-based repair in Lancaster Stemmer (SL).
-Figure 25: Comparison of using and without using context-based repair in Lemmatisation (L).
-Figure 26: Comparison of using and without using context-based repair in Lemmatisation + POS Tagging (LT).

### v4

@@
-How Does A Text Preprocessing Pipeline Affect Ontology Syntactic Matching?
-Abstract
-The generic text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many ontology matching (OM) systems.
+How Does A Text Preprocessing Pipeline Affect Syntactic Ontology Matching?
+Abstract.
+The generic text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many systems for syntactic ontology matching (OM).
 However, the lack of standardisation in text preprocessing creates diversity in mapping results.
-In this paper, we investigate the effect of the text preprocessing pipeline on OM tasks at syntactic levels.
-Our experiments on 8 Ontology Alignment Evaluation Initiative (OAEI) track repositories with 49 distinct alignments indicate: (1) Tokenisation and Normalisation are currently more effective than Stop Words Removal and Stemming/Lemmatisation; and (2) The selection of Lemmatisation and Stemming is task-specific.
-We recommend standalone Lemmatisation or Stemming with post-hoc corrections.
-We find that (3) Porter Stemmer and Snowball Stemmer perform better than Lancaster Stemmer; and that (4) Part-of-Speech (POS) Tagging does not help Lemmatisation.
-To repair less effective Stop Words Removal and Stemming/Lemmatisation used in OM tasks, we propose a novel context-based pipeline repair approach that significantly improves matching correctness and overall matching performance.
-We also discuss the use of text preprocessing pipeline in the new era of large language models (LLMs).
-Index Terms:
+In this paper, we investigate the effect of the text preprocessing pipeline on syntactic OM in 8 Ontology Alignment Evaluation Initiative (OAEI) tracks with 49 distinct alignments.
+We find that Phase 1 text preprocessing (Tokenisation and Normalisation) is currently more effective than Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation).
+To repair the less effective Phase 2 text preprocessing caused by unwanted false mappings, we propose a novel context-based pipeline repair approach that employs an ad hoc check to find common words that cause false mappings.
+These words are stored in a reserved word set and applied in text preprocessing.
+The experimental results show that our approach improves the matching correctness and the overall matching performance.
+We also discuss the integration of the classical text preprocessing pipeline with modern large language models (LLMs).
+We recommend that LLMs inject the text preprocessing pipeline via function calling to avoid the tendency towards unstable true mappings produced by prompt-based LLM approaches, and use LLMs to repair false mappings generated by the text preprocessing pipeline.
 Ontology Matching, Text Preprocessing
-I Introduction
-Ontology matching (OM), also known as ontology alignment, is essential to enable interoperability between heterogeneous ontologies.
-An OM process usually takes two ontologies as input, discovers mappings between entities, and produces a set of correspondences [1].
-A classical OM system usually contains syntactic, lexical, and semantic matching.
-This multilayer architecture has been implemented in several successful OM systems, such as LogMap [2][3], AgreementMakerLight (AML) [4][5], and FCA-Map [6][7].
-Syntactic matching captures “anchor mappings”, providing the foundation for the latter lexical and semantic matching.
-There are many strategies to extract syntactic information from an ontology entity, including the older approach of Bag-of-Words (e.g.
-TF-IDF [8]), popular word embedding models (e.g.
-Word2Vec [9]), and state-of-the-art language models (e.g.
-BERT [10]).
-Despite the diversity of the models used, they all apply text preprocessing for cleaning the text data before fitting it into the model.
-Fig.
-1 shows an example of using a classical text preprocessing pipeline to process the ontology entity “cmt:reviewerBiddingStartedBy”.
-The text preprocessing pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation [11].
-Tokenisation is the process of breaking the text into the smallest units (i.e.
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 29th International Conference on Extending Database Technology; March 24–27, 2026; Tampere, Finlandisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Ontology engineering
+1. Introduction
+Ontology matching (OM) is essential for enabling interoperability between heterogeneous ontologies.
+An OM process usually takes two ontologies as input, discovers mappings between entities, and produces a set of correspondences (Euzenat and Shvaiko, 2013).
+A classical OM system usually contains syntactic matching, lexical matching, and semantic matching.
+This multilayer architecture has been implemented in several successful OM systems, such as LogMap (Jiménez-Ruiz and Cuenca Grau, 2011; Jiménez-Ruiz et al., 2011), AgreementMakerLight (AML) (Faria et al., 2013, 2014), and FCA-Map (Zhao et al., 2018; Li et al., 2021).
+While syntactic matching captures “anchor mappings” and provides the foundation for the latter lexical and semantic matching, there are many strategies for extracting syntactic information from an ontology entity.
+These include older approaches Bag-of-Words (e.g.
+TF-IDF (Sammut and Webb, 2010)), popular word embedding models (e.g.
+Word2Vec (Mikolov et al., 2013)), and state-of-the-art language models (e.g.
+BERT (Devlin et al., 2019)).
+Despite the diversity of the models used, they all apply the text preprocessing pipeline to clean the text data before fitting them into the model.
+The pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L) (Manning et al., 2008).
+Table 1 illustrates examples of using the text preprocessing pipeline to find mappings in syntactic OM.
+T is the process of breaking the text into the smallest units (i.e.
 tokens).
-We use whitespace to split the tokens in the example.
-Normalisation is the process of transforming these different tokens into a single canonical form.
-Stop Words Removal is the process of removing filler words that usually carry little meaning and can be omitted in most cases.
-Stemming/Lemmatisation is used to deal with the grammatical variation of words, applying rules to find the simplest common form of the word.
-This helps to capture the key information from the text and therefore improves efficiency.
-Figure 1: Text preprocessing pipeline used in syntactic matching.
-While a number of OM systems use the text preprocessing pipeline for syntactic OM, few studies explain why a specific method is selected for a certain OM task.
-Our study tackles the challenge in two ways.
-Firstly, we conduct a comprehensive experimental analysis of the text preprocessing pipeline in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline in OM tasks at syntactic levels.
-In each phase, a text preprocessing method is evaluated for its correctness and completeness.
+N is the process of transforming these different tokens into a single canonical form.
+R is the process of removing filler words that usually carry little meaning and can be omitted in most cases.
+S/L is used to deal with the grammatical variation of words, applying rules to find the simplest common form of the word.
+Pipeline Ontology Entity Name Output T cmt ProgramCommittee Program Committee sigkdd Program_Committee N cmt ProgramCommittee program committee conference Program_committee R conference has_a_review review ekaw hasReview S/L conference reviewes review edas isReviewing
+Table 1.
+Examples of using the text preprocessing pipeline in syntactic OM: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
+In the examples, we use whitespace to split the tokens, normalise all the words into lowercase, remove the common stop words “has” and “a”, and apply S/L to find the simplest word “review” from “reviewes” and “reviewing”.
+While a number of OM systems use the text preprocessing pipeline for syntactic OM, few studies explain why a specific text preprocessing method is selected.
+Our study tackles this challenge.
+Firstly, we conduct a comprehensive experimental analysis of the text preprocessing pipeline in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline in syntactic OM.
+In each step, text preprocessing methods are evaluated for correctness and completeness.
 Secondly, we propose a novel context-based pipeline repair approach for syntactic OM.
-The method offers a customised way to fine-tune the text preprocessing pipeline for each domain-specific OM task and shows promising results for repairing false mappings.
-Thirdly, we discuss the non-deterministic nature of LLMs for text preprocessing and highlight the importance of retaining the classical text preprocessing pipeline.
+The method offers a customised way to fine-tune the text preprocessing pipeline for each domain-specific OM and shows promising results for repairing less effective text preprocessing methods.
+Thirdly, we discuss the non-deterministic nature of modern large language models (LLMs) for text preprocessing and highlight the importance of integrating the classical text preprocessing pipeline with LLMs.
 Specifically, this paper makes the following contributions:
-  • (Section IV) We categorise the text preprocessing pipeline used in syntactic OM into two phases.
-We find a significant improvement using Phase 1 text preprocessing methods.
-In contrast, Phase 2 text preprocessing methods are currently less effective.
-We compare the performance of (1) Stemming and Lemmatisation, (2) different stemming methods (Porter, Snowball, and Lancaster), and (3) Lemmatisation and Lemmatisation + Part-of-Speech (POS) Tagging.
-We find that inappropriate stop words removal, over-stemming, and over-lemmatisation are common on 8 Ontology Alignment Evaluation Initiative (OAEI) [12] track repositories with 49 distinct alignments.
-  • (Section V) We propose a simple and intuitive context-based pipeline repair method.
-This method is evaluated on the same OM tasks we analysed, showing promising results to improve the correctness of syntactic OM when inserted in the pipeline repair before Phase 2 text preprocessing methods.
-  • (Section VI) We discuss the use of text preprocessing pipeline in the era of LLMs.
-We find LLMs cannot perform syntactic matching without text preprocessing, but they can be used as an alternative to repair the false mappings.
-The remainder of the paper is organised as follows: Section II and Section III review the related work and the preliminaries of OM task formulation.
-Section IV analyses the text preprocessing pipeline used in OM.
-Section V proposes the context-based pipeline repair approach and experimentally validates its performance.
-Section VI demonstrates the pros and cons of using LLMs for text processing in OM.
-Section VII and Section VIII summarise the findings and limitations of this work.
-Section IX concludes the paper.
-II Related Work
-Syntactic matching considers only the meaning of the entity’s name or label, ignoring its lexical and structural context in an ontology [13].
-Correct syntactic matches are often implicit and usually require extra human observation and domain knowledge.
+  • (Section 4) We compare the performance of the text preprocessing pipeline in 8 Ontology Alignment Evaluation Initiative (OAEI) (OAEI Community, [n.
+d.]) datasets with 49 distinct alignments.
+While there is a significant improvement using Tokenisation (T) and Normalisation (N) in syntactic OM, Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are currently less effective.
+We find inappropriate stop words removal, over-stemming, and over-lemmatisation.
+We also compare the performance of several text preprocessing methods: (1) Stemming and Lemmatisation, (2) different stemming methods (Porter, Snowball, and Lancaster), and (3) Lemmatisation with and without Part-of-Speech (POS) Tagging.
+  • (Section 5) We propose a simple and intuitive context-based pipeline repair method using a reserved word set to capture certain words that may lead to false mappings.
+This method is evaluated on the same OAEI datasets we analysed, showing promising results in improving the matching correctness and the overall matching performance of syntactic OM.
+  • (Section 6) We discuss the integration of the classical text preprocessing pipeline with modern LLMs.
+We recommend using text preprocessing via function calling to overcome the non-deterministic of purely prompt-based approaches in finding true mappings, while LLMs show incredible potential to repair false mappings generated by the text preprocessing pipeline.
+The remainder of the paper is organised as follows.
+Section 2 and 3 review the related work and the preliminaries of OM task formulation.
+Section 4 analyses the text preprocessing pipeline used in syntactic OM.
+Section 5 proposes the context-based pipeline repair approach and experimentally validates its performance.
+Section 6 demonstrates the integration of the classical text preprocessing pipeline with modern LLMs in syntactic OM.
+Section 7 and 8 discuss the findings and limitations.
+Section 9 concludes the paper.
+2. Related Work
+Syntactic OM considers only the meaning of the entity’s name or label, ignoring its lexical and structural context in an ontology.
+Correct syntactic matches are often implicit and usually require extra human observation and domain knowledge (Euzenat and Shvaiko, 2013).
 Text preprocessing is introduced to automate this process.
-The use of text preprocessing in syntactic matching can be traced back to the early stages of OM systems, having been initially developed and widely used as a basic component to generate linguistic-based mappings.
-AgreementMaker [14] has a normaliser to unify the textual information of the entities.
-SAMBO [15] uses the Porter Stemmer for each word to improve the similarity measure for terms with different prefixes and suffixes.
-In RiMOM [16], the context information of each entity is viewed as a document.
+The use of text preprocessing in syntactic OM can be traced back to the early stages of OM systems, which were initially developed and widely used as a basic component to generate linguistic-based mappings.
+AgreementMaker (Cruz et al., 2009) has a normaliser to unify the textual information of the entities.
+SAMBO (Lambrix and Tan, 2006) uses the Porter Stemmer for each word to improve the similarity measure for terms with different prefixes and suffixes.
+In RiMOM (Li et al., 2009), the context information of each entity is viewed as a document.
 The text in each document is preprocessed with tokenisation, stop words removal, and stemming.
@@
 While text preprocessing remains useful, its role is more focused on normalising the text that becomes the input to the model.
-For example, BERTMap [17] uses BERT’s inherent WordPiece tokeniser [18] to build the subword of each entity.
-A more recent approach DeepOnto [19] extends the normalisation to axioms using EL embedding models [20].
-The ML extension of LogMap [21] reuses the seed mappings of the traditional system, where each entity is split into its component English word and the mapping is based on their similarity.
-However, to the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have been conducted to evaluate the effect of text preprocessing on syntactic OM.
-III Preliminaries: OM task formulation
-Given a source ontology $O_{s}$ and a target ontology $O_{t}$, OM establishes mappings between pairs of entities drawn from each of two ontologies.
+For example, BERTMap (He et al., 2022) uses BERT’s inherent WordPiece tokeniser to build the subword of each entity.
+DeepOnto (He et al., 2024) extends normalisation to axioms using EL embedding models (Kulmanov et al., 2019).
+The ML extension of LogMap (Chen et al., 2021) reuses the seed mappings of the traditional system, where each entity is split into its component word and the mapping is based on their similarity.
+However, to the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have been conducted to evaluate the effect of the entire text preprocessing pipeline on syntactic OM.
+3. Preliminaries: OM Task Formulation
+Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), OM establishes mappings between pairs of entities drawn from each of the two ontologies.
 A correspondence (i.e.
@@
 $r$ is the relationship between two matched entities $e_{1}$ and $e_{2}$, and $c\in[0,1]$ is the confidence for each correspondence.
-The relationship $r$ in OM tasks can be equivalence ($\equiv$), subsumption ($\subseteq$), disjointness ($\perp$), or other more complex relationships.
-In this paper, we focus only on the equivalence relationship ($\equiv$) and evaluate the effect of text preprocessing on syntactic matching to produce the “anchor mappings” on which to base any subsequent lexical and semantic matching.
-We address only equivalence because subsumption and disjointness are typically dealt with in a later semantic and structural matching phase of OM.
+The relationship $r$ in OM tasks can be equivalence ($\equiv$), subsumption ($\subseteq$), or other more complex relationships.
+In this study, we focus only on the equivalence relationship ($\equiv$) and evaluate the effect of text preprocessing on syntactic matching to produce the “anchor mappings” on which to base any subsequent lexical and semantic matching.
+We address only equivalence because subsumption is typically dealt with in a later semantic and structural matching phase of OM.
 An alignment (A) is a set of candidate correspondences generated by tools, while a reference (R) is a set of gold standard correspondences verified by domain experts (i.e.
 the ground truth alignment).
-IV Text Preprocessing Pipeline: Analysis
-In this section, we discuss the role of the text preprocessing pipeline in ontology syntactic matching and report our experiments studying the effect of different text preprocessing options over the OAEI repositories.
-IV-A Method
-Fig.
-2 shows the method used to analyse the text preprocessing pipeline.
+4. Analysis of Text Preprocessing Pipeline
+In this section, we discuss the role of the text preprocessing pipeline in syntactic OM and report our experiments that study the effect of different text preprocessing options over the OAEI datasets.
+4.1. Experiment Setup
+Figure 1 shows the experiment setup to analyse the text preprocessing pipeline.
 Alignment (A) is generated via the text preprocessing pipeline.
-Firstly, for both $O_{s}$ and $O_{t}$, we retrieve the entities from the named classes (i.e.
-owl:Class) and named properties (i.e.
-object properties owl:ObjectProperty and data type properties owl:DatatypeProperty).
+For both $O_{s}$ and $O_{t}$, we retrieve the entities from classes (owl:Class) and properties (both owl:ObjectProperty and owl:DatatypeProperty).
 For those ontologies where the names of the concepts are not textual (e.g.
-a numerical identifier), instead, we retrieve the meaningful text from entity labels (i.e.
-rdfs:label).
-Then, we apply the text preprocessing pipeline method $f(.)$ on each entity $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$.
+a code or numerical identifier), we retrieve the meaningful text from entity labels (rdfs:label or other annotation properties) instead.
+Then, we apply the text preprocessing pipeline method $f(\cdot)$ on each entity $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$.
 If $f(e_{1})=f(e_{2})$, we store the correspondence in the corresponding alignment file.
-Finally, we compare the generated Alignment (A) with Reference (R) to evaluate the performance of the text preprocessing pipeline on OM tasks.
-Figure 2: Method used to analyse the text preprocessing pipeline. B-Base Entity without Text Preprocessing, T-Tokenisation, N-Normalisation, R-Stop Words Removal, S/L-Stemming/Lemmatisation.
+We compare the generated alignment (A) with the reference (R) to evaluate the performance of the text preprocessing pipeline on syntactic OM.
+Figure 1. Experiment setup to analyse the text preprocessing pipeline: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
   1. (1)
-    Selected OAEI track repositories: The Ontology Matching Evaluation Toolkit (MELT) [22] is a powerful framework for OM evaluation.
-We retrieve 17 OAEI normal track repositories stored in the MELT public repository.
-The repositories are categorised as schema matching or instance matching.
-12 of the 17 track repositories are schema-matching and applicable for evaluating OM.
-Only 8 of these track repositories are selected because the other 4 have noisy data or miss required files.
-Specifically, the knowledgegraph repository contains both schema and instance matching.
-The complex repository is primarily focused on detecting complex correspondences, while the multifarm repository is an extension of the conference repository to multilingualism.
-The laboratory repository lacks reference files at the time of writing.
-Table I shows the details of the selected track repositories.
-We consider only the equivalence mappings contained in the reference files.
-Each track repository may contain more than one alignment corresponding to different pairs of ontologies.
-For example, the largebio track repository has 6 reference files, pairing the whole and small versions of FMA [23] and NCI [24], FMA and SNOMED [25], and SNOMED and NCI, respectively.
-The number of references for each track repository is given in the table.
+    Selected OAEI tracks: We select 8 OAEI tracks stored in the Ontology Matching Evaluation Toolkit (MELT) (Hertling et al., 2019) public repository (retrieved January 1, 2025).
+Table 2 shows the details of the selected tracks.
+Each track may contain more than one alignment corresponding to different pairs of ontologies.
+For example, the Largebio Track has multiple reference files, pairing FMA (Rosse and Mejino, 2003) and NCI (Golbeck et al., 2003), FMA and SNOMED (Donnelly, 2006), and SNOMED and NCI, respectively.
+The number of alignments for each track is given in the table.
 There are 49 distinct alignments evaluated in this study.
-    Name Domain Number of References anatomy Human and Mouse Anatomy 1 biodiv Biodiversity and Ecology 9 commonkg Common Knowledge Graphs 3 conference Conference 24 food Food Nutritional Composition 1 largebio Biomedical 6 mse Materials Science & Engineering 3 phenotype Disease and Phenotype 2
-    TABLE I: Selected OAEI track repositories.
+    Table 2. Selected OAEI tracks and the number of alignments.
+    Name Domain Number of Alignments Anatomy Human and Mouse Anatomy 1 Biodiv Biodiversity and Ecology 9 CommonKG Common Knowledge Graphs 3 Conference Conference 24 Food Food Nutritional Composition 1 Largebio Biomedical 6 MSE Materials Science & Engineering 3 Phenotype Disease and Phenotype 2
     Compound words are frequently used in ontology naming conventions.
-For example, compound words “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO [26] or the Snake case “art_gallery” used in Wikidata [27].
-Fig.
-3 shows the proportion of compound words in the selected OAEI track repositories.
+For example, compound words such as “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO (Suchanek et al., 2007) or the Snake case “art_gallery” used in Wikidata (Vrandečić and Krötzsch, 2014).
+Figure 2 shows the proportion of compound words in the selected OAEI tracks.
 For comparisons between entities with compound words, one approach is to use their tokens, where “XY” is equivalent to “YX” because they share the same concepts X and Y.
@@
 Although concurrent “XY” and “YX” are rare in the matching process, we assume that it could happen and so we take account of the order of concepts in compound word names to ensure a fair comparison in our experiments.
-    Figure 3: The proportion of compound words.
+    Figure 2. Proportion of compound words.
   2. (2)
-    Selected subtasks of pipeline methods: There are a variety of subtasks that can be used in a general text preprocessing pipeline, but not all of them are applicable to OM tasks.
-We select the following subtasks in each text preprocessing method: (1) Tokenisation: includes word tokenisation only.
-We do not use sentence tokenisation (also known as segmentation) because text retrieved from the entity’s name or label is commonly short.
-(2) Normalisation: includes lowercasing, HTML tags removal, separator formatting, and punctuation removal.
-Other subtasks that may potentially change the word semantics, such as removal of special characters and numbers, are excluded.
-(3) Stop Words Removal: includes the most common English stop words defined in the Natural Language Toolkit (NLTK) [28].
-(4) Stemming/Lemmatisation: Stemming methods include Porter Stemmer, Snowball Stemmer, and Lancaster Stemmer.
-Lemmatisation uses the NLTK Lemmatiser [28] based on WordNet [29], and the word categorisation uses POS Tagging.
+    Selected pipeline methods: There are a variety of methods that can be used in a general text preprocessing pipeline, but not all of them are applicable to syntactic OM.
+We select the following methods in each text preprocessing step: (a) Tokenisation (T): We use full word tokenisation to separate the compound words.
+Experiments on several subword tokenisation methods (e.g.
+Byte Pair Encoding (Gage, 1994; Sennrich et al., 2016), WordPiece (Schuster and Nakajima, 2012; Devlin et al., 2019), and Unigram (Kudo, 2018; Kudo and Richardson, 2018)) show the results after de-tokenisation (i.e.
+converting the tokens to a sentence) are generally the same as full word tokenisation.
+We do not use sentence tokenisation (a.k.a.
+segmentation) because text retrieved from the entity’s name or label is commonly short.
+(b) Normalisation (N): includes lowercasing, HTML tags removal, separator formatting, and punctuation removal.
+Other methods that may potentially change the word semantics are excluded (e.g.
+removal of special characters and numbers).
+(c) Stop Words Removal (R): includes the most common English stop words defined in the Natural Language Toolkit (NLTK) (Bird et al., 2008).
+(d) Stemming/Lemmatisation (S/L): Stemming methods include Porter, Snowball, and Lancaster.
+Lemmatisation uses the NLTK Lemmatiser based on WordNet (Miller, 1995), and the word categorisation uses POS Tagging.
   3. (3)
-    Selected OM evaluation measures: In information retrieval, there are four primitive measures: true positive (TP), false positive (FP), false negative (FN), and true negative (TN).
-In the context of OM, evaluation compares an alignment (A) returned by the OM system with a gold standard reference (R).
-Fig.
-4 illustrates that the four primitive measures in OM can be interpreted as $TP=A\cap R$, $FP=A-R$, $FN=R-A$, and $TN=(C\times C^{\prime})-(A\cup R)$, where $C\times C^{\prime}$ refers to all possible correspondences $\in\{O_{s},O_{t}\}$.
-    Figure 4: OM evaluation measures [30].
+    Selected measures: There are four primitive measures in information retrieval: true positive (TP), false positive (FP), false negative (FN), and true negative (TN).
+In the context of OM, the evaluation compares an alignment (A) returned by the OM system with a gold standard reference (R) validated by domain experts.
+Figure 3 illustrates that the four primitive measures in OM can be interpreted as $TP=A\cap R$, $FP=A-R$, $FN=R-A$, and $TN=(C\times C^{\prime})-(A\cup R)$, where $C\times C^{\prime}$ refers to all possible correspondences $\in\{O_{s},O_{t}\}$.
 Accuracy (Acc), Specificity (Spec), Precision (Prec), Recall (Rec), and $F_{\beta}$ Score ($F_{\beta}$) are the most common evaluation measures based on TP, FP, FN, and TN.
@@
 Therefore, we choose Precision (Prec), Recall (Rec), and F1 Score ($\beta=1$) in this study.
-They are defined as:
-    $\footnotesize Prec=\frac{|A\cap R|}{|A|}\quad Rec=\frac{|A\cap R|}{|R|}\quad F_{1}=\frac{2}{Prec^{-1}+Rec^{-1}}$ (1)
-IV-B Results
-  1. (1)
-    Comparison of pipeline methods: Fig.
-5 summarises the comparison of the text preprocessing methods Tokenization (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
-The result indicates that the vast majority of correct correspondences are found by T and N.
-We do not see R and S/L playing a prime role in the OM tasks.
-    Figure 5: Comparison of the generic text preprocessing pipeline: Base Entity without Text Preprocessing (B), Tokenisation (T), Normalisation (N), Stop Words Removal (R), Stemming/Lemmatisation (S/L).
+They are defined as follows.
+    Figure 3. OM Evaluation (Euzenat, 2007).
+    $Precision(Prec)=\frac{|A\cap R|}{|A|}$
+    $Recall(Rec)=\frac{|A\cap R|}{|R|}$
+    $F_{1}\ Score=\frac{2}{Prec^{-1}+Rec^{-1}}$
+4.2. Results
+4.2.1. Comparison of Text Preprocessing Pipeline
+Figure 4 shows the comparison of the text preprocessing pipeline.
+These include Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
 The methods are always applied sequentially in the pipeline.
+For T and N, most of the data points above the equivalent line in precision, recall, and F1 score indicate that they can help syntactic OM.
+For R and S/L, some data points are above the equivalent line in recall, but most data points are below the equivalent line in precision and F1 score indicate that they do not play a prime role.
 Three horizontal lines inside each violin plot show three quartiles: Q1, median, and Q3.
-The extension of the curves beyond 0% and 100% is an artefact of the seaborn [31] violin plot.
-(a) Precision: The median increases with T and N but decreases with R and S/L.
+The extension of the curves beyond 0% and 100% is an artefact of the seaborn (Waskom, 2021) package.
+In the violin plots: (a) Precision: The median increases with T and N but decreases with R and S/L.
 After T, the shape of the distribution is unchanged by R and S/L.
@@
 The shape of the distribution does not change after N.
-  2. (2)
-    Stemming vs. Lemmatisation: Fig. 6 compares Stemming (S) and Lemmatisation (L) in 49 alignments after Tokenisation (T) and Normalisation (N) have been applied. L is commonly better than S.
-    Figure 6: Comparison of Stemming (S) and Lemmatisation (L).
-(a) Precision: The median after L is greater than that achieved by S.
+Figure 4.
+Comparison of text preprocessing pipeline: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
+The symbol $\varnothing$ represents the baseline without text preprocessing.
+4.2.2. Stemming vs. Lemmatisation
+Figure 5 shows the comparison of Stemming (S) and Lemmatisation (L) after Tokenisation (T) and Normalisation (N) have been applied.
+Most data points above the L = S line in precision and F1 score indicate that L is better than S in syntactic OM (assuming that post hoc correction is excluded).
+In the violin plots: (a) Precision: The median after L is greater than that achieved by S.
 The shape of the distribution is slightly different.
@@
 The shape of the distribution is identical.
-  3. (3)
-    Porter Stemmer vs.
-Snowball Stemmer vs.
-Lancaster Stemmer: Fig.
-7 compares Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL) in 49 alignments.
-SP and SS have been found to perform better than SL.
-    Figure 7: Comparison of different stemmers: Porter Stemmer (SP), Snowball Stemmer (SS) and Lancaster Stemmer (SL).
-(a) Precision: The median number in SP and SS is greater than that in SL, and there is no difference between SP and SS.
+Figure 5. Stemming (S) vs. Lemmatisation (L).
+4.2.3. Different Stemming Methods
+Figure 6 shows the comparison of different stemming methods: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL).
+For SP and SS, all data points located on the equivalent line indicate that there is no difference between SP and SS.
+For SP/SS and SL, most data points above the equivalent line in precision and F1 score indicate that SP/SS is more effective than SL.
+In the violin plots: (a) Precision: The median number of SP and SS is greater than that of SL, and there is no difference between SP and SS.
 The shape of the distribution is identical.
 (b) Recall: The median number and the shape of the distribution are identical in SP, SS, and SL.
-(c) F1 Score: The median number in SP and SS is greater than that in SL, and there is no difference between SP and SS.
-The shape of the distribution is identical.
+(c) F1 Score: The median number of SP and SS is greater than that of SL, and there is no difference between SP and SS.
+The shape of the distribution is identical in SP, SS, and SL.
+Figure 6. Comparison of different stemmers: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL).
+4.2.4. Lemmatisation with/without POS Tagging
+Figure 7 shows the comparison of Lemmatisation (L) and Lemmatisation + POS Tagging (LT).
+All data points located on the equivalent line indicate that using POS Tagging in Lemmatisation does not help syntactic OM.
+In the violin plots, for each of (a) Precision, (b) Recall, and (c) F1 Score, the median number and the shape of the distribution are identical for L and LT.
+Figure 7. Lemmatisation (L) without/with POS Tagging (LT).
+4.3. Discussion
+For syntactic OM, the text preprocessing pipeline can be categorised into two phases.
+Phase 1 text preprocessing methods contain Normalisation and Tokenisation, whereas Phase 2 preprocessing methods consist of Stop Words Removal and Stemming/Lemmatisation.
+Phase 1 text preprocessing methods do not change word semantics, but change syntactic features such as formatting and typography.
+In contrast, Phase 2 text preprocessing methods change words to better capture semantic similarity, such as removing prefixes and suffixes or substituting common etymological roots.
   1. (1)
-    Lemmatisation vs Lemmatisation + POS Tagging: Fig.
-8 summarises the comparison of Lemmatisation (L) and Lemmatisation + POS Tagging (LT) in 49 alignments.
-The result indicates that POS Tagging does not help with Lemmatisation in Precision, Recall, and overall F1 Score in the total of 49 alignments we analysed.
-    Figure 8: Comparison of Lemmatisation (L) and Lemmatisation + POS Tagging (LT).
-For each of (a) Precision, (b) Recall and (c) F1 Score, the median and the shape of the distribution are identical for L and LT.
-IV-C Discussion
-For syntactic OM, the text preprocessing pipeline can be categorised into two phases.
-Phase 1 text preprocessing pipeline methods contain Normalisation and Tokenisation, whereas Stop Words Removal and Stemming/Lemmatisation are assigned to Phase 2.
-Phase 1 text preprocessing pipeline methods do not change word semantics; instead, they only change syntactic features such as formatting and typography.
-Conversely, in Phase 2 of the text preprocessing pipeline, words are changed to better capture semantic similarity, such as removing prefixes and suffixes or substituting common etymological roots.
-  1. (1)
-    Phase 1 text preprocessing methods: We observe that matching performance usually increases with each Phase 1 text preprocessing method, indicating the benefit of these “rules of thumb” for syntactic OM.
-If two terms can be matched using a heuristic or intuitive technique, there is no need to leverage the more complex Phase 2 methods.
+    Phase 1 text preprocessing methods: We observe that matching performance usually increases with each Phase 1 text preprocessing method.
 However, Phase 1 methods could also benefit from customisation in OM tasks.
-Some traditional methods originating for natural language processing (NLP) are not useful for OM and need to be adjusted.
-For example, sentence segmentation (i.e.splitting long text into sentences) is not applicable for ontology entities because they are generally short text fragments and do not need such operations.
+Some traditional methods originating from natural language processing (NLP) are not useful for OM and need to be adjusted.
+For example, sentence segmentation (i.e.
+splitting long text into sentences) is not applicable for ontology entities because they are generally short text fragments and do not need such operations.
 Word tokenisation (i.e.
@@
   2. (2)
-    Phase 2 text preprocessing methods: We observe no benefit from Phase 2 text preprocessing methods.
-In some cases, Stop Words Removal and Stemming/Lemmatisation may even hamper the mapping.
+    Phase 2 text preprocessing methods: We observe no benefit in using Phase 2 text preprocessing methods in syntactic OM.
+In some cases, Stop Words Removal and Stemming/Lemmatisation may even hamper the matching performance.
 There are plausible explanations for this behaviour arising from the nature of ontologies as distinct from natural language text, as follows:
     1. (a)
-      Stop Words Removal: “AND”, “OR”, and “NOT” are stop words in English and usually carry little useful information, whereas these words express logical operations in ontology entities and therefore may carry important semantics.
+      Stop Words Removal (R): “AND”, “OR”, and “NOT” are stop words in English and usually carry little useful information, whereas these words express logical operations in ontology entities and may carry important semantics.
     2. (b)
-      Stemming vs.
-Lemmatisation: Based on our experiments, Lemmatisation is better than Stemming.
-While lemmatisation tends to avoid generating FPs, it may also miss some implicit TPs.
-The complexity of finding a missing TP is equal to the size of $(C^{\prime})$.
-On the other hand, stemming is more aggressive in finding TPs, but the aggression can lead to more FPs as well.
-The complexity of finding FPs is equal to the smaller size of $(A)$.
-The workload of discovering FPs after stemming is generally much lighter than detecting missing TPs after lemmatisation, but this may also depend on the accuracy of post-hoc corrections that can be performed.
+      Stemming (S) vs.
+Lemmatisation (L): While L tends to avoid generating FPs, it may also miss some implicit TPs.
+On the other hand, S is more aggressive in finding TPs, but aggression can also lead to more FPs being found.
+Based on our experiments with no post hoc correction applied, L is better than S.
     3. (c)
-      Porter Stemmer vs.
-Snowball Stemmer vs.
-Lancaster Stemmer: Porter Stemmer and Snowball Stemmer (also known as Porter 2 Stemmer) have been found to perform better than Lancaster Stemmer, and we cannot see a significant difference between Porter and Snowball.
+      Different stemming methods: Porter Stemmer (PS) and Snowball Stemmer (SS) have been found to perform better than Lancaster Stemmer (LS), and we cannot see a significant difference between Porter and Snowball.
 Although the Lancaster Stemmer is more aggressive in detecting more TPs, it does not lead to performance improvement as more FPs are matched.
     4. (d)
-      Lemmatisation vs.
-Lemmatisation + POS Tagging: Lemmatisation + POS Tagging is generally expected to have better results than Lemmatisation alone because tagging can help detect more precise root words.
+      Lemmatisation (L) vs.
+Lemmatisation + POS Tagging (LT): LT is generally expected to have better results than L alone because tagging can help detect more precise root words.
 However, we do not observe such a performance improvement when using POS tagging in our study.
 The reason may be that ontology classes are usually nouns or gerunds, and in such cases, we could expect the simpler grammatical assumption to have similar results.
-V Context-based Pipeline Repair
-Experimental results in Section IV demonstrate that only Phase 1 text processing (i.e.
-Tokenisation and Normalisation) helps with syntactic OM.
-The use of Phase 2 text preprocessing (i.e.
-Stop Words Removal and Stemming/Lemmatisation) does not improve performance and may even have negative impacts.
-In this section, we explore the underlying cause of this behaviour.
-Furthermore, we introduce an intuitive context-based pipeline repair approach and demonstrate its effectiveness on the same datasets.
-V-A Causal analysis
+  3. (3)
+    In most cases, OM only requires minor preprocessing using Phase 1 text preprocessing methods.
+This is because entity names in the ontology are often compound words that do not occur in natural language, but they are partially formalised by agreement.
+There have been well-defined conventions established over the years, e.g.
+singularity, positive names, nouns for classes and verbs for properties (Schober et al., 2007; Taylor et al., 2015).
+5. Context-Based Pipeline Repair
+The experimental results in Section 4 demonstrate that only Phase 1 text processing (Tokenisation and Normalisation) helps with syntactic OM.
+The use of Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation) does not improve performance and may even have negative impacts.
+In this section, we explore the underlying cause of this behaviour and introduce an intuitive context-based pipeline repair approach.
+5.1. Causal Analysis
   1. (1)
@@
 For example, isReviewing is equivalent to is_reviewing.
-This means that applying Phase 1 methods only helps detect TPs, while the number of FPs remains unchanged.
+This means that applying Phase 1 methods only helps to detect TPs, while the number of FPs remains unchanged.
 For this reason, Phase 1 methods always have a positive effect on precision, recall, and overall F1 score.
-    $\footnotesize Prec\uparrow=\frac{|A\cap R|}{|A|}=\frac{TP}{TP+FP}=1-\frac{FP}{TP\uparrow+FP}$ (2)
-    $\footnotesize Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$ (3)
-    $\footnotesize F_{1}\uparrow=\frac{2}{Prec\uparrow^{-1}+Rec\uparrow^{-1}}$ (4)
-    In most cases, OM only requires minor preprocessing using Phase 1 text preprocessing methods.
-This is because entity names in the ontology are often compound words that do not occur in natural language, but they are partially formalised by agreement.
-There have been well-defined conventions established over the years, e.g.
-singularity, positive names, nouns for classes and verbs for properties [32] [33].
+    (1) $Prec\uparrow=\frac{|A\cap R|}{|A|}=\frac{TP}{TP+FP}=1-\frac{FP}{TP\uparrow+FP}$
+    (2) $Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$
+    (3) $F_{1}\ Score\uparrow=\frac{2}{Prec\uparrow^{-1}+Rec\uparrow^{-1}}$
   2. (2)
@@
 For this reason, recall is always increasing, but precision and overall F1 score are less reliable.
-    $\footnotesize Prec\ ?=\frac{|A\cap R|}{|A|}=\frac{TP\uparrow}{TP\uparrow+FP\uparrow}$ (5)
-    $\footnotesize Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$ (6)
-    $\footnotesize F_{1}\ ?=\frac{2}{Prec\ ?^{-1}+Rec\uparrow^{-1}}$ (7)
+    (4) $Prec\ ?=\frac{|A\cap R|}{|A|}=\frac{TP\uparrow}{TP\uparrow+FP\uparrow}$
+    (5) $Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$
+    (6) $F_{1}\ Score\ ?=\frac{2}{Prec\ ?^{-1}+Rec\uparrow^{-1}}$
     If we define $\Delta TP$ and $\Delta FP$ as the increase in TP and FP for a preprocessing method, then the threshold to increase precision and F1 score is:
-    $\scriptsize\frac{TP+\Delta TP}{TP+\Delta TP+FP+\Delta FP}>\frac{TP}{TP+FP}\Rightarrow\frac{\Delta TP}{\Delta FP}>\frac{TP}{FP}$ (8)
+    (7) $\frac{TP+\Delta TP}{TP+\Delta TP+FP+\Delta FP}>\frac{TP}{TP+FP}\Rightarrow\frac{\Delta TP}{\Delta FP}>\frac{TP}{FP}$
     In our experiments, we actually observe a reduction in precision and overall F1 score.
@@
 So performance does not improve, unless the benefit of each TP is considered more valuable than the disbenefit of each FP.
-This could apply, for example, if we are expecting a post-hoc correction phase where removing FPs is considered to be an easier human task than adding missing TPs.
-V-B Repair Method
-Phase 2 text preprocessing pipeline methods (Stop Words Removal and Stemming/Lemmatisation) have been shown to be less effective in OM tasks, caused mainly by FPs.
-We propose a pipeline repair approach that aims to differentiate FPs and therefore improve Precision and F1 Score.
-As illustrated in Fig.
-9: (1) Phase 1 methods shift Alignment (A) towards Reference (R).
-The number of TPs increases, while FPs decrease.
+This could apply, for example, if we are expecting a post hoc correction phase where removing FPs is considered to be an easier human task than adding missing TPs.
+5.2. Repair Method
+While Phase 2 text preprocessing pipeline methods have been shown to be less effective in syntactic OM mainly caused by FPs, we propose a pipeline repair approach to differentiate FPs from TPs.
+As illustrated in Figure 8: (1) Phase 1 methods shift Alignment (A) towards Reference (R).
+The number of TPs increases, while the number of FPs decreases.
 (2) Phase 2 methods expand Alignment (A).
-The number of TPs increases, but FPs also increase.
-(3) Our context-based pipeline repair approach collapses Alignment (A).
+The number of TPs increases, but the number of FPs also increases.
+(3) Our context-based pipeline repair aims to collapse Alignment (A).
 The number of FPs significantly decreases, with a slight decrease in TPs.
-Figure 9: Mechanism of the context-based pipeline repair.
+Figure 8. Motivation of our context-based pipeline repair.
 One critical step in our approach is to retrieve a reserved word set that may cause FPs after the text preprocessing, and these words will be excluded before the text preprocessing.
@@
 For both $O_{s}$ and $O_{t}$, the algorithm first iterates on all pairs of entities $e_{i},e_{j}$ in each of them.
-For a specific text preprocessing method $f(.)$, if $f(e_{i})=f(e_{j})$, we retrieve the different words between $e_{i}$ and $e_{j}$ and store them in the reserved word set.
+For a specific text preprocessing method $f(\cdot)$, if $f(e_{i})=f(e_{j})$, we retrieve the different words between $e_{i}$ and $e_{j}$ and store them in the reserved word set.
 If a word appears in the reserved set, the text preprocessing pipeline skips the operation for this word.
 To simplify the reserved word set, we also remove the words where $f(w)=w$ from the final set because either skipping or keeping these words in the reserved word set would not change the mapping results.
-Algorithm 1 Finding the reserved words
+Algorithm 1 Finding Reserved Word Set
 0:  Source Ontology $O_{s}$, Target Ontology $O_{t}$, Text Preprocessing Pipeline Method $f(x)$
@@
 An example of generating and using a simple reserved word set is illustrated below.
-(1) Two object properties was_a_member_of and has_members from a single ontology have the same result “member” via the traditional text preprocessing.
+(1) The two object properties was_a_member_of and has_members from a single ontology have the same result “member” via traditional text preprocessing.
 Because there are no duplicate entities within a single ontology, we use a reserved word set in our proposed pipeline repair approach to determine that they are distinguishing entities.
@@
 While a false mapping may occur when they both have the same result “steer committe” after the traditional text preprocessing, using the reserved word set can repair this false mapping.
-As the words “has”, “a”, “was”, and “of” are listed as reserved words, these two named properties are preprocessed as “has a steer committe” and “was a steer committe of”, respectively.
-V-C Evaluation
-We apply our context-based pipeline repair approach to the same OAEI track repositories and alignments as above.
-Fig.
-10 compares with and without context-based pipeline repair in 8 track repositories with 49 distinct alignments.
-The text preprocessing pipeline methods implemented include all the Phase 2 text preprocessing methods: Stop Words Removal (R), Porter Stemmer (SP), Snowball Stemmer (SS), Lancaster Stemmer (SL), Lemmatisation (L), and Lemmatisation + POS Tagging (LT).
-We can see that our context-based repair approach can significantly improve the Precision.
-As a trade-off, it may cause a slight decrease in Recall, but the overall F1 Score is still increasing in the majority of the alignments.
-For example, this repair approach applied to the MaterialInformation-EMMO alignment using Lancaster Stemmer improved Precision and overall F1 score by 8.95% and 2.92%, with only a decrease of 3.17% in Recall.
-Figure 10: Context-based pipeline repair use in Phase 2 text preprocessing methods (the total number of each category can appear to be less than 49 when data points overlap).
-VI Text Preprocessing in the era of LLMs
-With the recent advances in LLMs, we ask whether we still need the classic text processing pipeline.
-In this section, we demonstrate the non-deterministic nature of LLMs in OM text preprocessing and showcase the benefits of LLMs in a repair function inside the OM text preprocessing pipeline.
-VI-A LLMs Used in OM Text Preprocessing
-We observe that some LLMs can perform syntactic matching without preprocessing, but their robustness remains questionable.
-We conduct a study on the use of LLMs for syntactic matching with and without text preprocessing.
-The experiment settings are described as follows:
-  1. (1)
-    We select the same example described in Section IV-A.
-The compound word “art gallery” can have different naming conventions in different ontologies.
-For example, the Pascal case “ArtGallery” used in YAGO and the Snake case “art_gallery” used in Wikidata.
+As the words “has”, “a”, “was”, and “of” are listed as reserved words, these two named properties are processed as “has a steer committe” and “was a steer committe of”, respectively.
+5.3. Evaluation
+We apply our context-based pipeline repair approach to the same OAEI datasets as above.
+Figure 9 compares with and without context-based pipeline repair in 8 tracks with 49 distinct alignments.
+The text preprocessing pipeline methods implemented include all Phase 2 text preprocessing methods: Stop Words Removal (R), Porter Stemmer (SP), Snowball Stemmer (SS), Lancaster Stemmer (SL), Lemmatisation (L), and Lemmatisation + POS Tagging (LT).
+(a) Precision: Most data points are above the equivalent line, indicating that context-based pipeline repair significantly improves the matching correctness.
+(b) Recall: Most data points are located on the equivalent line, and only a few data points are below the equivalent line, indicating that context-based pipeline repair slightly reduces matching completeness.
+(c) F1 Score: Most of the data points are above the equivalent line, indicating that context-based pipeline repair also improves overall matching performance.
+Experimentally, the wider ellipse around the data points indicates that the matching performance improvement ranking in Phase 2 text preprocessing methods is Stemming (S) > Lemmatisation (L) > Stop Words Removal (R); the matching performance improvement ranking in different stemmers is Lancaster Stemmer (SL) > Porter Stemmer (SP) = Snowball Stemmer (SS); and the matching performance improvement ranking in lemmatisation is Lemmatisation (L) $\approx$ Lemmatisation + POS Tagging (LT).
+Figure 9.
+Context-based pipeline repair use in Phase 2 text preprocessing: Stop Words Removal (R), Porter Stemmer (SP), Snowball Stemmer (SS), Lancaster Stemmer (SL), Lemmatisation (L), and Lemmatisation + POS Tagging (LT).
+6. Text Preprocessing Pipeline & LLMs
+In this section, we discuss the pros and cons of LLMs in performing syntactic OM and demonstrate the benefits of integrating the classical text preprocessing pipeline with modern LLMs.
+6.1. Motivation
+In modern LLMs, text preprocessing can be performed implicitly within the models.
+For example, LLMs may understand that “art gallery” and “ArtGallery” are true mappings even if they are in different formats.
+However, it is questionable whether this behaviour is stable across different LLMs and prompt templates and whether LLMs outperform the classical text processing pipeline.
+On the other hand, with a strong background knowledge, LLMs may be able to facilitate mapping repair together with the text preprocessing pipeline.
+For example, LLMs can classify “Bile” and “Bilateral” as false mappings, whereas the classical text preprocessing method stemming cannot.
+6.2. Experiment Setup
   1. (1)
     We choose 10 LLMs from 4 different families.
-These include two OpenAI models (gpt-3.5-turbo and gpt-4-turbo) [34], three Anthropic Claude models (claude-3-opus, claude-3-sonnet, and claude-3-haiku) [35], three Mistral AI models (mistral-large, mistral-medium, and mistral-small) [36], and two open-source Meta Llama models (llama-3-70b and llama-3-8b) [37].
-We use LangChain [38] to build the chats with OpenAI, Mistral AI, and Anthropic Claude models, and Ollama [39] to access Meta Llama models.
-The versions of LLMs used in the experiment are listed in Table II.
-All the model temperatures are set to 0 to minimise the random results generated by LLMs.
-    Series LLMs Versions GPT gpt-4-turbo gpt-4-turbo-2024-04-09 gpt-3.5-turbo gpt-3.5-turbo-0125 Claude claude-3-opus claude-3-opus-20240229 claude-3-sonnet claude-3-sonnet-20240229 claude-3-haiku claude-3-haiku-20240307 Mistral mistral-large mistral-large-2402 mistral-medium mistral-medium-2312 mistral-small mistral-small-2402 Llama llama-3-70b date accessed: 2024-06-01 llama-3-8b date accessed: 2024-06-01
-    TABLE II: Versions of LLMs used in the experiment.
+These include two OpenAI GPT models (OpenAI, [n.
+d.]) (gpt-4o and gpt-4o-mini), two Anthropic Claude models (Anthropic, [n.
+d.])(claude-3-sonnet and claude-3-haiku), and two versions of Meta Llama (Meta, [n.
+d.]) (llama-3-8b and llama-3.1-8b), two versions of Alibaba Qwen (Alibaba Qwen Team, [n.
+d.]) (qwen-2-7b and qwen-2.5-7b), Google Gemma 2 (Google Gemma Team and Google DeepMind, [n.
+d.]), and ChatGLM 4 (Zeng et al., 2024).
+We use LangChain (LangChain, Inc., [n.
+d.]) to build chats with OpenAI GPT and Anthropic Claude models, and Ollama (Ollama, [n.
+d.]) to access open-source models.
+Table 3 lists the details of selected LLMs.
+All model temperatures are set to 0 to minimise the random results generated by LLMs.
+    Table 3. Details of selected LLMs (retrieved from  (Qiang et al., 2023)).
+    Family Model Size Version GPT gpt-4o N/A gpt-4o-2024-05-13 gpt-4o-mini N/A gpt-4o-mini-2024-07-18 Claude claude-3-sonnet N/A claude-3-sonnet-20240229 claude-3-haiku N/A claude-3-haiku-20240307 Llama llama-3-8b* 4.7 GB Ollama Model ID: 365c0bd3c000 llama-3.1-8b* 4.9 GB Ollama Model ID: 46e0c10c039e Qwen qwen-2-7b* 4.4 GB Ollama Model ID: dd314f039b9d qwen-2.5-7b* 4.7 GB Ollama Model ID: 845dbda0ea48 Gemma gemma-2-9b* 5.4 GB Ollama Model ID: ff02c3702f32 GLM glm-4-9b* 5.5 GB Ollama Model ID: 5b699761eca5 * Open-source LLM (retrieved December 1, 2024).
   2. (2)
-    We generate the prompt template as “Is ArtGallery <KEYWORD> to/with art_galley?” (without text preprocessing) and “Is art galley <KEYWORD> to/with art galley?” (with text preprocessing).
-We use only T and N to preprocess the text because R and S/L are shown to be less effective in the main content of the paper.
-Furthermore, R and S/L needs are not evident in the original naming of the ontology entities.
-The <KEYWORD> is a placeholder for a collection of words that can be used to describe an equivalence relationship.
-In this study, we experiment with three common words namely “identical”, “interchangeable”, and “equivalent”.
+    We select 4 different LLM prompt strategies.
+(a) Basic: The simplest prompt template to compare two entities using “Is Entity1 equivalent to Entity2?
+Answer yes or no.” (b) Few-Shot: Adding an example in the prompt, such as “Example: Hair_root is equivalent to Hair_Root.” (c) Self-Reflection: Before LLMs answer the question, ask LLMs to rethink the output and give a short explanation.
+We add a sentence “Write a short explanation.” to invoke this function.
+(d) Few-Shot + Self-Reflection: A combination of (b) and (c).
+Table 4 lists the details of the selected prompt strategies and their templates.
+    Table 4. Selected prompt templates. The few-shot parts are coloured blue, and the self-reflection parts are orange.
+    No.
+Prompt Strategy Prompt Template PT1 Basic Is Entity1 equivalent to Entity2?
+Answer yes or no.
+PT2 Few-Shot Example: Hair_root is equivalent to Hair_Root.
+Is Entity1 equivalent to Entity2?
+Answer yes or no.
+PT3 Self-Reflection Is Entity1 equivalent to Entity2?
+Answer yes or no.
+Write a short explanation.
+PT4 Few-Shot Example: Hair_root is equivalent to Hair_Root.
++ Is Entity1 equivalent to Entity2?
+Answer yes or no.
+Self-Reflection Write a short explanation.
   3. (3)
-    Considering the complexity of prompts, we also test an additional case where the prompts have a self-reflection phase (i.e.
-add “Write a short explanation” to the prompt).
-We use parentheses to indicate that the different responses generated by the prompts have a self-reflection phase.
-Table III shows the results of using LLMs for syntactic matching without text preprocessing.
-We expect the LLMs to generate “Yes” answers for this TP sample.
-However, the majority of the LLMs output the “No” answers (marked with the red colour in the table) without text preprocessing.
-LLMs <KEYWORD> “identical” “interchangeable” “equivalent” gpt-4-turbo No No No gpt-3.5-turbo No No No claude-3-opus No Yes Yes claude-3-sonnet No No No claude-3-haiku No No No mistral-large No No Yes mistral-medium No No No mistral-small No Yes Yes (No) llama-3-70b No No No llama-3-8b Yes Yes Yes
-TABLE III: Using LLMs for syntactic matching without text preprocessing. Prompt: “Is ArtGallery <KEYWORD> to/with art_gallery? Answer yes or no. (Write a short explanation.)”
-Table IV shows the results of using LLMs for syntactic matching with text preprocessing.
-We can see that the error rate can be significantly reduced, but some LLMs still continue to produce incorrect results.
-The result also shows that LLMs with a larger number of parameters (e.g.
-llama-3-70b) and prompts with a self-reflection phase (labelled in parentheses) offer “too much of a good thing”.
-They do not help with syntactic matching and may even produce more errors in OM tasks.
-Although this preliminary study does not cover all possible triggers used in prompt engineering, it is adequate to highlight the importance of proper preprocessing before fitting the text into the LLMs.
-We believe that the classical programming-based text preprocessing pipeline is more stable and reliable than using LLM’s implicit text preprocessing or prompt-based text preprocessing for OM tasks.
-LLMs <KEYWORD> “identical” “interchangeable” “equivalent” gpt-4-turbo Yes Yes Yes gpt-3.5-turbo Yes Yes Yes claude-3-opus Yes Yes Yes claude-3-sonnet Yes (No) No Yes (No) claude-3-haiku No No Yes (No) mistral-large Yes Yes Yes mistral-medium No Yes Yes mistral-small Yes Yes Yes llama-3-70b Yes Yes (No) Yes (No) llama-3-8b Yes Yes Yes
-TABLE IV: Using LLMs for syntactic matching with text preprocessing. Prompt: “Is art gallery <KEYWORD> to/with art gallery? Answer yes or no. (Write a short explanation.)”
-VI-B LLMs Used in OM Text Preprocessing Pipeline Repair
-Using LLMs for OM text preprocessing is not without merit.
-With strong background knowledge, they can be used to facilitate several subtasks in OM.
-For example, LLMs can be used as a repair tool for the text preprocessing pipeline, as an alternative to the approach we proposed in Section V-B.
-The slight difference is that the context-based pipeline repair approach is an ad hoc repair prior to syntactic matching, while the LLM-based approach is a post hoc repair after syntactic matching.
-We conduct a preliminary study on LLMs used in OM text preprocessing pipeline repair.
-The experiment settings are described as follows:
+    We run the experiment on the Largebio Track that contains pairwise alignments between FMA, NCI, and SNOMED.
+The true mappings (TPs) are retrieved from the reference (R).
+The false mappings (FPs) are retrieved from the alignment (A) generated by the text preprocessing pipeline but exclude the true mappings (TPs) listed in the reference (R).
   4. (4)
-    The experiment is set up on the largebio track repository, a biomedical track repository requires higher Precision with low rates of FPs.
-The Lancaster Stemmer always shows lower Precision across 6 pairs of alignments, meaning that the results contain a significant number of FPs.
+    We report our results over a single run.
+There could be slight differences in the results over multiple trails, but these differences are not significant with respect to the ranking and trend of different LLMs and prompt templates.
+Figure 10. Discovery rate of TPs using LLMs on the Largebio Track.
+Figure 11. The discovery rate of FPs using LLMs on the Largebio Track.
+6.3. Detecting True Mappings
+LLMs are expected to answer “yes” for TPs.
+The discovery rate is defined as the number of “yes” answers compared to the total number of TPs.
+For LLMs, we directly prompt LLMs to determine whether two entities are identical or not.
+For the text preprocessing pipeline, we employ the pipeline via function calling (a.k.a tool calling).
+This means that the LLMs act as an interface, and mapping judgement is based on programme logic defined in the text preprocessing method.
+Two entities are considered to be matched if they have the same results after the text preprocessing.
+Figure 10 shows the discovery rate of TPs using LLMs on the Largebio Track.
+Here we use only Tokenisation (T) and Normalisation (N) to preprocess the text because Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are shown to be less effective from our analysis in Section 4.
+Although some API-accessed commercial LLMs have shown the capability to detect TPs without a text preprocessing step, this behaviour is non-deterministic across different LLMs and prompt templates.
+For different LLMs, open-source LLMs perform better than commercial API-accessed LLMs.
+It could be problematic for some commercial API-accessed LLMs (e.g.
+claude-3haiku) to find two identical entities if text preprocessing is not applied prior to the comparison.
+They may consider entities with different formatting that are used in different contexts and therefore have different meanings.
+For different prompt templates, few-shot (PT2) significantly improves correctness, but self-reflection (PT3) may even hamper performance.
+The combination of few-shot and self-reflection (PT4) may even perform worse than the simple basic prompt (PT1).
+We believe that this phenomenon is mainly caused by LLM “overthinking”, a type of LLM hallucinations in which LLMs could over-analyse user input and generate unnecessarily complex details.
+We also need to note here that the example generated in the few-shot prompt is based on our observation of the existing TPs.
+Finding a good example in real-world applications is not easy because ontology entities can have different formatting by nature.
+On the other hand, using the classical text preprocessing pipeline via function calling could help LLMs produce a stable result with a high discovery rate (almost 100% from our observation) across different models and prompt templates.
+6.4. Detecting False Mappings
+LLMs are expected to answer “no” for FPs.
+The discovery rate is defined as the number of “no” answers compared to the total number of FPs.
+Specifically, these FPs are non-trivial mappings that the classical text preprocessing methods cannot detect.
+Figure 11 shows the discovery rate of FPs using LLMs on the Largebio Track.
+Here we use LLMs to detect FPs generated by the Lancaster Stemmer (SL).
+SL has shown a lower precision compared to other text preprocessing methods, which is caused by the method producing a significant number of FPs.
+LLMs have shown a strong capability to discover the FPs produced by the classical text preprocessing pipeline.
+For different models, API-accessed commercial LLMs are better than open-source LLMs and reach a higher discovery rate of more than 80%.
+For different prompt templates, both few-shot (PT2), self-reflection (PT3), and a combination of the above (PT4) are better than a simple basic prompt (PT4).
+While mapping repair is usually non-trivial and requires domain expertise, LLMs have shown potential for performing a high-quality post hoc correction without human intervention.
+6.5. Implications
+LLMs were originally developed and trained for question answering (QA).
+OM is not only a QA task, but also a knowledge-intensive task that requires a comprehensive understanding and reasoning of domain knowledge.
+The studies in this section demonstrate the limitations and opportunities of integrating the text preprocessing pipeline with LLMs in syntactic OM.
+We highlight two pathways for this integration.
+One pathway is to use function calling to invoke the classical text preprocessing pipeline within LLMs, which packages the programming-based text preprocessing pipeline as a tool for LLMs to use.
+This provides stable results for text preprocessing in syntactic OM.
+The other pathway is to use LLMs as a repair tool after text preprocessing, which could be a cost-effective solution for post hoc correction.
+This can also be extended using human-in-the-loop to validate the repaired results generated by LLMs and feed the validated data back into LLMs so that they can better understand the context of the task (i.e.
+fine-tuning LLMs used in syntactic OM).
+7. Discussion
+OM aims to bridge the gap by capturing similar concepts that occur in different ontologies.
+OM is a complex process that combines syntactic matching, lexical matching, and semantic matching.
+Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching.
+This study concentrates on a common practice using the text preprocessing pipeline in syntactic OM.
+While generic text preprocessing pipeline methods in syntactic OM are usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
   1. (1)
-    We choose the same 10 LLMs used in the previous experiment in Appendix VI-A.
-  3. (3)
-    We use the prompt “Is X equivalent to Y?” and preprocess the entity names X and Y with N and T before fitting them into the LLMs.
-The previous experiment in Appendix VI-A shows these two settings are more suitable for OM tasks.
-  4. (4)
-    Considering the complexity of prompts, we also test an additional case where the prompts have a self-reflection phase (i.e. add “Write a short explanation” to the prompt).
-Figure 11: The discovery rate of FPs using LLMs on the largebio track repository.
-Fig.
-11 shows the discovery rate of FPs using 10 different LLMs.
-We expect the LLMs to generate “No” answers for these FP samples.
-The GPT model gpt-4-tubo, the Claude model claude-3-haiku, the Mistral model mistral-medium, and the Llama model llama-3-70b achieve better performance in detecting FPs.
-For the GPT and Llama series, LLMs with a larger number of parameters perform better than those with a relatively smaller number of parameters.
-For the Claude and Mistral series, LLMs with larger parameters do not improve performance.
-In the test cases we analysed, they are even less effective than those with a relatively smaller number of parameters.
-Unlike detecting TPs, we can see a self-reflection phase improves the overall performance of detecting FPs.
-VI-C Implications
-LLMs were originally developed for the purpose of question-answering (QA).
-OM is not only a QA task, but also a knowledge-intensive task that relates to a comprehensive understanding and reasoning of domain knowledge.
-The studies in this section demonstrate the limitations and opportunities of using LLMs for OM tasks.
-From our experience, we highlight two pathways to customise LLM for a new domain task.
-One pathway is function calling (FC), which refers to teaching LLMs using external tools.
-In LLM-based text preprocessing, we could package the programming-based text preprocessing pipeline as a tool for LLMs to use.
-The other pathway is fine-tuning (FT), where LLMs are customised with local data.
-In LLM-based text preprocessing repair, we could use human-in-the-loop to validate the repaired results generated by LLMs and feed the validated data back into LLMs, so that they can better understand the context of the task.
-VII Discussion
-Ontologies provide formalised conceptualisations of knowledge graphs (KGs).
-Incorporating ontologies for KG-related NLP tasks can enhance the capability to handle complex concepts and relations, thereby improving the quality of KGs built from text.
-Ontologies also assist interoperability of text generated from KGs.
-However, aligning and integrating heterogeneous ontologies remains a challenge when using them for NLP tasks.
-OM aims to bridge the gap by capturing similar concepts that occur in different ontologies.
-OM is a complex process that combines syntactic, lexical, and semantic matching.
-Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching.
-This paper concentrates on text preprocessing, a common practice in syntactic matching.
-While generic text preprocessing pipeline methods in syntactic OM are usually applied on the basis of intuition or extrapolation from other experience, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
-  1. (1)
-    Our study shows (i) whether to use, or not use, and (ii) how to use these text preprocessing methods.
-Such experimental results will benefit decision making when selecting appropriate text preprocessing methods for syntactic OM.
+    Our study shows (a) whether to use, or not use, and (b) how to use these text preprocessing methods.
+Such experimental results will benefit decision-making when selecting appropriate text preprocessing methods for syntactic OM.
 For large-scale OM, it can significantly reduce unnecessary trial costs.
@@
 Its broader value is showing how to maximise true mappings and minimise false mappings throughout the text preprocessing pipeline.
-From a statistical perspective, it is a local optimisation for syntactic matching that will benefit the global optimisation for OM, as syntactic matching provides “anchor mappings” for lexical and semantic matching.
+From a statistical perspective, it is a local optimisation for syntactic matching that will benefit the global optimisation for OM, as syntactic matching provides “anchor mappings” for lexical matching and semantic matching.
   3. (3)
     The use of LLMs is not a one-size-fits-all solution for OM tasks.
-OM is a complex multi-level matching mixed with syntactic, lexical, and semantic matching.
+OM is a complex multi-level matching mixed with different matching techniques.
 LLMs may not outperform traditional methods in performing these sub-matching tasks.
+We suggest “rules of thumb” for syntactic OM.
+If two ontology entities can be matched using the intuitive text preprocessing pipeline, there is no need to leverage LLMs for the same purpose.
+  4. (4)
     The development and implementation of an LLM-based framework that incorporates different matching techniques is a new research direction for OM, but it is beyond the scope of this paper.
-We refer readers to several recent works in [40, 41, 42, 43, 44].
+We refer the reader to several recent works in (He et al., 2023; Norouzi et al., 2023; Hertling and Paulheim, 2023; Qiang et al., 2023; Giglou et al., 2024; Amini et al., 2024).
+  5. (5)
     This study also demonstrates that the nature of text preprocessing used in NLP applications can be context-based.
-Not all pipeline methods are effective, and some of them may even hamper the overall performance of downstream tasks.
+Not all pipeline methods are effective, and some of them may even hamper the performance of downstream tasks.
 Modifications are required before applying text preprocessing to specific NLP tasks.
-VIII Limitations
-OAEI track repositories are comprehensive but do not cover all real-world scenarios.
-Furthermore, the reference files contained in the OAEI track repositories are not “complete” gold standards and some need further development.
-In this study, we only deal with equivalence mappings between named classes, object properties, and data type properties.
-Disjointness mappings are not considered.
+8. Limitations
+The references contained in the OAEI tracks are not “complete” gold standards and some need further development.
+In this study, we only deal with equivalence mappings between classes and properties.
+Subsumption mappings are not considered.
 Finding widely used baseline data with ground truth matches for subsumptions is empirically difficult.
-Only a few OAEI track repositories contain subsumption mappings, and they are not enough to robustly demonstrate text preprocessing used in subsumption mappings.
-This work has not been entered into the official OAEI campaign because of its intense focus on only the preprocessing step.
-IX Conclusion
+Only a few OAEI tracks contain subsumption mappings, and they are not enough to robustly demonstrate the text preprocessing used in subsumption mappings.
+This work has not been included in the official OAEI campaign because of its intense focus on only text preprocessing in syntactic OM.
+9. Conclusion
 In this paper, we conduct a comprehensive study on the effect of the text preprocessing pipeline on syntactic OM.
-8 OAEI track repositories with 49 distinct alignments are evaluated.
-Despite the importance of text preprocessing in syntactic OM, our experimental results indicate that the text preprocessing pipeline is currently ill-equipped to handle OM tasks.
+8 OAEI tracks with 49 distinct alignments are evaluated.
+Despite the importance of text preprocessing in syntactic OM, our experimental results indicate that the text preprocessing pipeline is currently ill-equipped.
 (1) We find that Phase 1 text preprocessing methods (Tokenisation and Normalisation) help with both matching completeness (i.e.
@@
 F1 score).
-(3) We also argue the traditional text preprocessing pipeline is not obsolete in the era of LLMs, but an integral part for LLMs used in the OM tasks.
-Our future work will focus on handling class axioms and complex relationships to evaluate the text preprocessing pipeline for OM tasks.
+(3) We also argue that the traditional text preprocessing pipeline is not obsolete in the era of LLMs.
+It is an integral part of the LLMs used in OM tasks.
+Integrating the classical text preprocessing pipeline can significantly improve the robustness of purely prompt-based approaches.
+Our future work will focus on handling class axioms and complex relationships to evaluate the text preprocessing pipeline.
 We will also study the pipeline and pipeline repair approach working with both knowledge-based OM systems and LLM-based OM systems.
-Supplemental Material Statement
-The source code for this study is available from https://github.com/qzc438-research/ontology-nlp.
-The OAEI track repositories used in this study can be downloaded from OAEI MELT at https://dwslab.github.io/melt/track-repository (date accessed: 2024-06-01).
-The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.
-Acknowledgment
-The authors would like to thank Alice Richardson of the Statistical Support Network, Australian National University (ANU), for helpful advice on the statistical analysis in this paper.
-The authors also thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.
+Acknowledgements.
+The authors thank Sven Hertling for curating the Ontology Alignment Evaluation Initiative (OAEI) datasets.
+The authors thank Ernesto Jimenez-Ruiz for helpful advice on reproducing the OAEI Largebio Track.
+The authors thank Alice Richardson of the Statistical Support Network, Australian National University (ANU), for helpful advice on the statistical analysis in this paper.
+The authors thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.

### v5

@@
-How Does A Text Preprocessing Pipeline Affect Syntactic Ontology Matching?
+How Does A Text Preprocessing Pipeline Affect Syntactic Ontology Matching? [Experiment, Analysis & Benchmark]
 Abstract.
-The generic text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many systems for syntactic ontology matching (OM).
+The classic text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many systems for syntactic ontology matching (OM).
 However, the lack of standardisation in text preprocessing creates diversity in mapping results.
-In this paper, we investigate the effect of the text preprocessing pipeline on syntactic OM in 8 Ontology Alignment Evaluation Initiative (OAEI) tracks with 49 distinct alignments.
-We find that Phase 1 text preprocessing (Tokenisation and Normalisation) is currently more effective than Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation).
-To repair the less effective Phase 2 text preprocessing caused by unwanted false mappings, we propose a novel context-based pipeline repair approach that employs an ad hoc check to find common words that cause false mappings.
+In this paper we investigate the effect of the text preprocessing pipeline on syntactic OM in 8 Ontology Alignment Evaluation Initiative (OAEI) tracks with 49 distinct alignments.
+We find that Phase 1 text preprocessing (Tokenisation and Normalisation) is more effective than Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation).
+To repair the unwanted false mappings caused by Phase 2 text preprocessing, we propose a novel context-based pipeline repair approach that employs a post hoc check to find common words that cause false mappings.
 These words are stored in a reserved word set and applied in text preprocessing.
 The experimental results show that our approach improves the matching correctness and the overall matching performance.
-We also discuss the integration of the classical text preprocessing pipeline with modern large language models (LLMs).
-We recommend that LLMs inject the text preprocessing pipeline via function calling to avoid the tendency towards unstable true mappings produced by prompt-based LLM approaches, and use LLMs to repair false mappings generated by the text preprocessing pipeline.
-Ontology Matching, Text Preprocessing
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 29th International Conference on Extending Database Technology; March 24–27, 2026; Tampere, Finlandisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Ontology engineering
+We then consider the broader integration of the classic text preprocessing pipeline with modern large language models (LLMs) for OM.
+We recommend that (1) the text preprocessing pipeline be injected via function calling into LLMs to avoid the tendency towards unstable true mappings produced by LLM prompting; or (2) LLMs be used to repair non-existent and counter-intuitive false mappings generated by the text preprocessing pipeline.
+PVLDB Reference Format:
+PVLDB, 19(1): XXX-XXX, 2025.
+doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
+Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
+For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
+Copyright is held by the owner/author(s).
+Publication rights licensed to the VLDB Endowment.
+Proceedings of the VLDB Endowment, Vol. 19, No. 1 ISSN 2150-8097.
+doi:XX.XX/XXX.XX
+PVLDB Artifact Availability:
+The source code, data, and/or other artifacts have been made available at https://github.com/qzc438-research/ontology-nlp.
 1. Introduction
@@
 BERT (Devlin et al., 2019)).
-Despite the diversity of the models used, they all apply the text preprocessing pipeline to clean the text data before fitting them into the model.
+Despite the diversity of the models used, they all apply the text preprocessing pipeline to clean the text data before using the model.
 The pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L) (Manning et al., 2008).
-Table 1 illustrates examples of using the text preprocessing pipeline to find mappings in syntactic OM.
+Table 1 shows examples of using the text preprocessing pipeline to find mappings in syntactic OM.
 T is the process of breaking the text into the smallest units (i.e.
@@
 S/L is used to deal with the grammatical variation of words, applying rules to find the simplest common form of the word.
-Pipeline Ontology Entity Name Output T cmt ProgramCommittee Program Committee sigkdd Program_Committee N cmt ProgramCommittee program committee conference Program_committee R conference has_a_review review ekaw hasReview S/L conference reviewes review edas isReviewing
+Pipeline Ontology Entity Name Output T cmt ProgramCommittee Program Committee sigkdd Program_Committee N cmt ProgramCommittee program committee conference Program_committee R conference has_a_review review ekaw hasReview S/L conference reviewes review edas (is)Reviewing
 Table 1.
 Examples of using the text preprocessing pipeline in syntactic OM: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
-In the examples, we use whitespace to split the tokens, normalise all the words into lowercase, remove the common stop words “has” and “a”, and apply S/L to find the simplest word “review” from “reviewes” and “reviewing”.
-While a number of OM systems use the text preprocessing pipeline for syntactic OM, few studies explain why a specific text preprocessing method is selected.
-Our study tackles this challenge.
-Firstly, we conduct a comprehensive experimental analysis of the text preprocessing pipeline in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline in syntactic OM.
+In the examples, we use whitespace to split the tokens, normalise all the words into lowercase, remove the common stop words “has” and “a”, and apply S/L to find the simplest word form.
+While a number of OM systems use the text preprocessing pipeline for syntactic OM, few studies question its role in optimising the matching performance.
+Our study fills this gap.
+Firstly, we conduct a comprehensive experimental analysis of the text preprocessing pipeline in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline.
 In each step, text preprocessing methods are evaluated for correctness and completeness.
 Secondly, we propose a novel context-based pipeline repair approach for syntactic OM.
-The method offers a customised way to fine-tune the text preprocessing pipeline for each domain-specific OM and shows promising results for repairing less effective text preprocessing methods.
-Thirdly, we discuss the non-deterministic nature of modern large language models (LLMs) for text preprocessing and highlight the importance of integrating the classical text preprocessing pipeline with LLMs.
+The method offers a customised way to fine-tune the text preprocessing pipeline for each domain-specific OM.
+This method shows promising results for repairing less effective text preprocessing methods.
+Thirdly, we discuss the non-deterministic nature of modern large language models (LLMs) for text preprocessing and highlight the importance of integrating the classic text preprocessing pipeline with LLM-based OM.
 Specifically, this paper makes the following contributions:
-  • (Section 4) We compare the performance of the text preprocessing pipeline in 8 Ontology Alignment Evaluation Initiative (OAEI) (OAEI Community, [n.
-d.]) datasets with 49 distinct alignments.
+  • (Section 4) We compare the performance of the text preprocessing pipeline on 8 Ontology Alignment Evaluation Initiative (OAEI) (OAEI Community, [n.d.]) datasets with 49 distinct alignments.
 While there is a significant improvement using Tokenisation (T) and Normalisation (N) in syntactic OM, Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are currently less effective.
@@
 This method is evaluated on the same OAEI datasets we analysed, showing promising results in improving the matching correctness and the overall matching performance of syntactic OM.
-  • (Section 6) We discuss the integration of the classical text preprocessing pipeline with modern LLMs.
-We recommend using text preprocessing via function calling to overcome the non-deterministic of purely prompt-based approaches in finding true mappings, while LLMs show incredible potential to repair false mappings generated by the text preprocessing pipeline.
+  • (Section 6) We discuss the integration of the classic text preprocessing pipeline with modern LLMs.
+We recommend using text preprocessing via function calling to overcome the non-determinism of purely prompt-based approaches to finding true mappings, while LLMs show great potential to repair false mappings generated by the text preprocessing pipeline.
 The remainder of the paper is organised as follows.
@@
 Section 5 proposes the context-based pipeline repair approach and experimentally validates its performance.
-Section 6 demonstrates the integration of the classical text preprocessing pipeline with modern LLMs in syntactic OM.
+Section 6 demonstrates the integration of the classic text preprocessing pipeline with modern LLMs in syntactic OM.
 Section 7 and 8 discuss the findings and limitations.
-Section 9 concludes the paper.
+Section 9 concludes.
 2. Related Work
-Syntactic OM considers only the meaning of the entity’s name or label, ignoring its lexical and structural context in an ontology.
+Syntactic OM considers only the meaning of an entity’s name or label, ignoring its lexical and structural context in an ontology.
 Correct syntactic matches are often implicit and usually require extra human observation and domain knowledge (Euzenat and Shvaiko, 2013).
@@
 DeepOnto (He et al., 2024) extends normalisation to axioms using EL embedding models (Kulmanov et al., 2019).
-The ML extension of LogMap (Chen et al., 2021) reuses the seed mappings of the traditional system, where each entity is split into its component word and the mapping is based on their similarity.
-However, to the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have been conducted to evaluate the effect of the entire text preprocessing pipeline on syntactic OM.
+The ML extension of LogMap (Chen et al., 2021) reuses the seed mappings of traditional systems, where each entity is split into its component words and the mapping is based on pairwise component similarities.
+However, to the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have evaluated the effect of the entire text preprocessing pipeline on syntactic OM.
 3. Preliminaries: OM Task Formulation
 Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), OM establishes mappings between pairs of entities drawn from each of the two ontologies.
-A correspondence (i.e.
-one instance of mappings) is defined as a 4-tuple $(e_{1},e_{2},r,c)$, where $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$.
-$r$ is the relationship between two matched entities $e_{1}$ and $e_{2}$, and $c\in[0,1]$ is the confidence for each correspondence.
+A correspondence (a.k.a.
+a mapping) is defined as a 4-tuple $(e_{1},e_{2},r,c)$, where $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$.
+$r$ is the relationship between two matched entities $e_{1}$ and $e_{2}$, and $c\in[0,1]$ is the confidence of the correspondence (or may be absent).
 The relationship $r$ in OM tasks can be equivalence ($\equiv$), subsumption ($\subseteq$), or other more complex relationships.
@@
 We address only equivalence because subsumption is typically dealt with in a later semantic and structural matching phase of OM.
-An alignment (A) is a set of candidate correspondences generated by tools, while a reference (R) is a set of gold standard correspondences verified by domain experts (i.e.
+An alignment (A) is a set of candidate correspondences, while a reference (R) is a set of gold-standard correspondences verified by domain experts (i.e.
 the ground truth alignment).
-4. Analysis of Text Preprocessing Pipeline
-In this section, we discuss the role of the text preprocessing pipeline in syntactic OM and report our experiments that study the effect of different text preprocessing options over the OAEI datasets.
+4. Text Preprocessing Pipeline Analysis
+In this section, we discuss the role of the text preprocessing pipeline in syntactic OM and report our experiments that study the effect of different text preprocessing options on the OAEI datasets.
 4.1. Experiment Setup
@@
 segmentation) because text retrieved from the entity’s name or label is commonly short.
-(b) Normalisation (N): includes lowercasing, HTML tags removal, separator formatting, and punctuation removal.
+(b) Normalisation (N): includes lowercasing, HTML tag removal, separator formatting, and punctuation removal.
 Other methods that may potentially change the word semantics are excluded (e.g.
@@
 In the context of OM, the evaluation compares an alignment (A) returned by the OM system with a gold standard reference (R) validated by domain experts.
-Figure 3 illustrates that the four primitive measures in OM can be interpreted as $TP=A\cap R$, $FP=A-R$, $FN=R-A$, and $TN=(C\times C^{\prime})-(A\cup R)$, where $C\times C^{\prime}$ refers to all possible correspondences $\in\{O_{s},O_{t}\}$.
+Figure 3 illustrates that the four primitive measures in OM can be interpreted as $TP=A\cap R$, $FP=A-R$, $FN=R-A$, and $TN=(C\times C^{\prime})-(A\cup R)$, where $C\times C^{\prime}$ refers to all possible correspondences $\{(e_{1},e_{2},\equiv)|e_{1}\in O_{s},e_{2}\in O_{t}\}$.
 Accuracy (Acc), Specificity (Spec), Precision (Prec), Recall (Rec), and $F_{\beta}$ Score ($F_{\beta}$) are the most common evaluation measures based on TP, FP, FN, and TN.
@@
 4.2.1. Comparison of Text Preprocessing Pipeline
-Figure 4 shows the comparison of the text preprocessing pipeline.
-These include Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
+Figure 4 compares the stages of the text preprocessing pipeline.
+Beginning with $\varnothing$ (no text preprocessing), we see the performance of subsequent Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
 The methods are always applied sequentially in the pipeline.
-For T and N, most of the data points above the equivalent line in precision, recall, and F1 score indicate that they can help syntactic OM.
-For R and S/L, some data points are above the equivalent line in recall, but most data points are below the equivalent line in precision and F1 score indicate that they do not play a prime role.
+For T and N, most of the data points above the equivalence line in precision, recall, and F1 score indicate that they can help syntactic OM.
+For R and S/L, some data points are above the equivalence line in recall, but most data points are below the equivalence line in precision and F1 score indicating that they do not play a positive role.
 Three horizontal lines inside each violin plot show three quartiles: Q1, median, and Q3.
@@
 4.2.2. Stemming vs. Lemmatisation
-Figure 5 shows the comparison of Stemming (S) and Lemmatisation (L) after Tokenisation (T) and Normalisation (N) have been applied.
+Figure 5 compares Stemming (S) and Lemmatisation (L) after Tokenisation (T) and Normalisation (N) have been applied.
 Most data points above the L = S line in precision and F1 score indicate that L is better than S in syntactic OM (assuming that post hoc correction is excluded).
@@
 4.2.3. Different Stemming Methods
-Figure 6 shows the comparison of different stemming methods: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL).
-For SP and SS, all data points located on the equivalent line indicate that there is no difference between SP and SS.
-For SP/SS and SL, most data points above the equivalent line in precision and F1 score indicate that SP/SS is more effective than SL.
+Figure 6 compares different stemming methods: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL).
+For SP and SS, all data points located on the equivalence line indicate that there is no difference between SP and SS.
+For SP/SS and SL, most data points above the equivalence line in precision and F1 score indicate that SP/SS is more effective than SL.
 In the violin plots: (a) Precision: The median number of SP and SS is greater than that of SL, and there is no difference between SP and SS.
@@
 4.2.4. Lemmatisation with/without POS Tagging
-Figure 7 shows the comparison of Lemmatisation (L) and Lemmatisation + POS Tagging (LT).
-All data points located on the equivalent line indicate that using POS Tagging in Lemmatisation does not help syntactic OM.
+Figure 7 compares Lemmatisation (L) and Lemmatisation + POS Tagging (LT).
+All data points located on the equivalence line indicate that using POS Tagging in Lemmatisation does not help syntactic OM.
 In the violin plots, for each of (a) Precision, (b) Recall, and (c) F1 Score, the median number and the shape of the distribution are identical for L and LT.
 Figure 7. Lemmatisation (L) without/with POS Tagging (LT).
-4.3. Discussion
+4.3. Implications
 For syntactic OM, the text preprocessing pipeline can be categorised into two phases.
@@
 Word tokenisation (i.e.
-breaking text into single words) could be rewritten to detect the abbreviations that are common practice in ontology concept names and to tackle the conflicting use of camel case and snake case.
+breaking text into single words) could be rewritten to detect the abbreviations that are common practice in ontology concept names and to tackle the conflicting use of Pascal case and snake case.
   2. (2)
@@
     1. (a)
-      Stop Words Removal (R): “AND”, “OR”, and “NOT” are stop words in English and usually carry little useful information, whereas these words express logical operations in ontology entities and may carry important semantics.
+      Stop Words Removal (R): Logical expressions “and”, “or”, and “not” are stop words in English and usually carry little useful information, whereas these words express logical operations in ontology entities and may carry important semantics.
+Some single stop word characters (e.g.
+“d”, “i”, “m”, “o”, “s”, “t”, and “y”) should be removed from the stop word set because contractions (e.g.
+I’d, I’m, o’clock, it’s, isn’t, y’know) are unlikely to appear in ontology entities, and yet those single characters are likely to carry important meaning if they do occur (e.g.
+“Chromosome_Y”).
     2. (b)
@@
 Lemmatisation (L): While L tends to avoid generating FPs, it may also miss some implicit TPs.
-On the other hand, S is more aggressive in finding TPs, but aggression can also lead to more FPs being found.
+On the other hand, S is more eager to find TPs, but eagerness can also lead to more FPs being found.
 Based on our experiments with no post hoc correction applied, L is better than S.
@@
       Different stemming methods: Porter Stemmer (PS) and Snowball Stemmer (SS) have been found to perform better than Lancaster Stemmer (LS), and we cannot see a significant difference between Porter and Snowball.
-Although the Lancaster Stemmer is more aggressive in detecting more TPs, it does not lead to performance improvement as more FPs are matched.
+Although the Lancaster Stemmer is more eager to detect more TPs, it does not lead to performance improvement as more FPs are also matched.
     4. (d)
@@
 For example, isReviewing is equivalent to is_reviewing.
-This means that applying Phase 1 methods only helps to detect TPs, while the number of FPs remains unchanged.
+This means that applying Phase 1 methods only help to detect TPs, while the number of FPs remains unchanged.
 For this reason, Phase 1 methods always have a positive effect on precision, recall, and overall F1 score.
@@
 Moving through the text preprocessing pipeline tends to detect more TPs and FPs in the derived alignment (A).
-For example, isReviewing and isReviewedBy may be object properties with distinctly different meaning, but removing the common stop words “is” and “by”, and using Stemming/Lemmatisation to retrieve the same root word “Review”, could cause a false match.
-For this reason, recall is always increasing, but precision and overall F1 score are less reliable.
+For example, isReviewing and isReviewedBy may be object properties with distinctly different meanings, but removing the common stop words “is” and “by”, and using Stemming/Lemmatisation to retrieve the same root word “Review”, could cause a false match.
+For this reason, recall is always increasing, but precision and the overall F1 score are less reliable.
     (4) $Prec\ ?=\frac{|A\cap R|}{|A|}=\frac{TP\uparrow}{TP\uparrow+FP\uparrow}$
@@
     (6) $F_{1}\ Score\ ?=\frac{2}{Prec\ ?^{-1}+Rec\uparrow^{-1}}$
-    If we define $\Delta TP$ and $\Delta FP$ as the increase in TP and FP for a preprocessing method, then the threshold to increase precision and F1 score is:
+    If we define $\Delta TP$ and $\Delta FP$ as the increase in TP and FP for a preprocessing method, then the condition to increase precision and F1 score is:
     (7) $\frac{TP+\Delta TP}{TP+\Delta TP+FP+\Delta FP}>\frac{TP}{TP+FP}\Rightarrow\frac{\Delta TP}{\Delta FP}>\frac{TP}{FP}$
@@
 So performance does not improve, unless the benefit of each TP is considered more valuable than the disbenefit of each FP.
-This could apply, for example, if we are expecting a post hoc correction phase where removing FPs is considered to be an easier human task than adding missing TPs.
+This could apply, for example, if we are expecting a post hoc correction phase in which removing FPs is considered to be an easier human task than adding missing TPs.
+This idea inspires our novel pipeline repair method.
 5.2. Repair Method
-While Phase 2 text preprocessing pipeline methods have been shown to be less effective in syntactic OM mainly caused by FPs, we propose a pipeline repair approach to differentiate FPs from TPs.
+While Phase 2 text preprocessing methods have been shown to be less effective in syntactic OM by introducing FPs, we propose a pipeline repair approach to differentiate FPs from TPs.
 As illustrated in Figure 8: (1) Phase 1 methods shift Alignment (A) towards Reference (R).
@@
 (3) Our context-based pipeline repair aims to collapse Alignment (A).
-The number of FPs significantly decreases, with a slight decrease in TPs.
+The number of FPs decreases significantly, with a slight decrease in TPs.
 Figure 8. Motivation of our context-based pipeline repair.
-One critical step in our approach is to retrieve a reserved word set that may cause FPs after the text preprocessing, and these words will be excluded before the text preprocessing.
-The selection criteria are based on two widely agreed assumptions: (1) there are no duplicate entities within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies (we call our approach context-based here because pairs of words may have the same or different meanings in different contexts).
-Based on these two assumptions, we propose a simple and intuitive Algorithm 1 to retrieve the reserved word set for context-based pipeline repair.
-For both $O_{s}$ and $O_{t}$, the algorithm first iterates on all pairs of entities $e_{i},e_{j}$ in each of them.
+One critical step in our approach is to assemble a set of reserved words that may cause FPs in text preprocessing, and to exclude these words from text preprocessing.
+The selection criteria are based on two widely agreed assumptions: (1) there are no duplicate entities within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies.
+We call our approach context-based here because pairs of words may have the same or different meanings in different contexts.
+Based on these two assumptions, we propose a simple Algorithm 1 to assemble the reserved word set.
+For both $O_{s}$ and $O_{t}$, the algorithm first iterates on all pairs of entities $e_{i},e_{j}$ in the ontology, respectively.
 For a specific text preprocessing method $f(\cdot)$, if $f(e_{i})=f(e_{j})$, we retrieve the different words between $e_{i}$ and $e_{j}$ and store them in the reserved word set.
 If a word appears in the reserved set, the text preprocessing pipeline skips the operation for this word.
-To simplify the reserved word set, we also remove the words where $f(w)=w$ from the final set because either skipping or keeping these words in the reserved word set would not change the mapping results.
+To simplify the reserved word set, we remove the words where $f(w)=w$ because their occurrence in the reserved word set has no effect on mapping results.
 Algorithm 1 Finding Reserved Word Set
@@
      if $f(e_{i})=f(e_{j})$ then
+        /* Pick up every different word between $e_{i}$ and $e_{j}$*/
         $Reserved\_Word\_Set\leftarrow differ(e_{i},e_{j})$
@@
   /* Same procedure applies …*/
-  /* Phase 2: Find duplicates in $Reserved\_Word\_Set$/
+  /* Phase 2: Find immutable words in $Reserved\_Word\_Set$/
   for $Word\ w\in Reserved\_Word\_Set$ do
@@
 The revision step (i.e.
-Phase 2 of Algorithm 1) is to check whether there are duplicates in the reserved set.
-We can observe that the word “member” is a duplicate because it is the same before and after text preprocessing.
+Phase 2 of Algorithm 1) is to check whether there are immutable words in the reserved set.
+We can observe that the word “member” is immutable because it is the same before and after text preprocessing.
 Removing this word from the reserved word set still makes the two object properties different.
@@
 The text preprocessing pipeline methods implemented include all Phase 2 text preprocessing methods: Stop Words Removal (R), Porter Stemmer (SP), Snowball Stemmer (SS), Lancaster Stemmer (SL), Lemmatisation (L), and Lemmatisation + POS Tagging (LT).
-(a) Precision: Most data points are above the equivalent line, indicating that context-based pipeline repair significantly improves the matching correctness.
-(b) Recall: Most data points are located on the equivalent line, and only a few data points are below the equivalent line, indicating that context-based pipeline repair slightly reduces matching completeness.
-(c) F1 Score: Most of the data points are above the equivalent line, indicating that context-based pipeline repair also improves overall matching performance.
-Experimentally, the wider ellipse around the data points indicates that the matching performance improvement ranking in Phase 2 text preprocessing methods is Stemming (S) > Lemmatisation (L) > Stop Words Removal (R); the matching performance improvement ranking in different stemmers is Lancaster Stemmer (SL) > Porter Stemmer (SP) = Snowball Stemmer (SS); and the matching performance improvement ranking in lemmatisation is Lemmatisation (L) $\approx$ Lemmatisation + POS Tagging (LT).
 Figure 9.
 Context-based pipeline repair use in Phase 2 text preprocessing: Stop Words Removal (R), Porter Stemmer (SP), Snowball Stemmer (SS), Lancaster Stemmer (SL), Lemmatisation (L), and Lemmatisation + POS Tagging (LT).
+  1. (1)
+    Precision: Most data points are above the equivalence line, indicating that context-based pipeline repair significantly improves matching correctness.
+Recall: Most data points are located on the equivalence line, and only a few data points are below the equivalence line, indicating that context-based pipeline repair slightly reduces matching completeness.
+F1 Score: Most of the data points are above the equivalence line, indicating that context-based pipeline repair also improves overall matching performance.
+  2. (2)
+    Experimentally, the wider ellipse around the data points indicates that the matching performance improvement ranking in Phase 2 text preprocessing methods is Stemming (S) > Lemmatisation (L) > Stop Words Removal (R); the matching performance improvement ranking in different stemmers is Lancaster Stemmer (SL) > Porter Stemmer (SP) = Snowball Stemmer (SS); and the matching performance improvement ranking in lemmatisation is Lemmatisation (L) $\approx$ Lemmatisation + POS Tagging (LT).
 6. Text Preprocessing Pipeline & LLMs
-In this section, we discuss the pros and cons of LLMs in performing syntactic OM and demonstrate the benefits of integrating the classical text preprocessing pipeline with modern LLMs.
+In this section, we discuss the pros and cons of LLMs in performing syntactic OM and demonstrate the benefits of integrating the classic text preprocessing pipeline with modern LLMs.
 6.1. Motivation
@@
 For example, LLMs may understand that “art gallery” and “ArtGallery” are true mappings even if they are in different formats.
-However, it is questionable whether this behaviour is stable across different LLMs and prompt templates and whether LLMs outperform the classical text processing pipeline.
+However, it is questionable whether this behaviour is stable across different LLMs and prompt templates and whether LLMs outperform the classic text processing pipeline.
 On the other hand, with a strong background knowledge, LLMs may be able to facilitate mapping repair together with the text preprocessing pipeline.
-For example, LLMs can classify “Bile” and “Bilateral” as false mappings, whereas the classical text preprocessing method stemming cannot.
+For example, LLMs can classify “Bile” and “Bilateral” as false mappings, whereas the classic text preprocessing method stemming cannot.
 6.2. Experiment Setup
@@
     We choose 10 LLMs from 4 different families.
-These include two OpenAI GPT models (OpenAI, [n.
-d.]) (gpt-4o and gpt-4o-mini), two Anthropic Claude models (Anthropic, [n.
-d.])(claude-3-sonnet and claude-3-haiku), and two versions of Meta Llama (Meta, [n.
-d.]) (llama-3-8b and llama-3.1-8b), two versions of Alibaba Qwen (Alibaba Qwen Team, [n.
-d.]) (qwen-2-7b and qwen-2.5-7b), Google Gemma 2 (Google Gemma Team and Google DeepMind, [n.
-d.]), and ChatGLM 4 (Zeng et al., 2024).
-We use LangChain (LangChain, Inc., [n.
-d.]) to build chats with OpenAI GPT and Anthropic Claude models, and Ollama (Ollama, [n.
-d.]) to access open-source models.
+These include two OpenAI GPT models (OpenAI, [n.d.]) (gpt-4o and gpt-4o-mini), two Anthropic Claude models (Anthropic, [n.d.])(claude-3-sonnet and claude-3-haiku), and two versions of Meta Llama (Meta, [n.d.]) (llama-3-8b and llama-3.1-8b), two versions of Alibaba Qwen (Alibaba Qwen Team, [n.d.]) (qwen-2-7b and qwen-2.5-7b), Google Gemma 2 (Google Gemma Team and Google DeepMind, [n.d.]), and ChatGLM 4 (Zeng et al., 2024).
+We use LangChain (LangChain, Inc., [n.d.]) to build chats with OpenAI GPT and Anthropic Claude models, and Ollama (Ollama, [n.d.]) to access open-source models.
 Table 3 lists the details of selected LLMs.
 All model temperatures are set to 0 to minimise the random results generated by LLMs.
-    Table 3. Details of selected LLMs (retrieved from  (Qiang et al., 2023)).
-    Family Model Size Version GPT gpt-4o N/A gpt-4o-2024-05-13 gpt-4o-mini N/A gpt-4o-mini-2024-07-18 Claude claude-3-sonnet N/A claude-3-sonnet-20240229 claude-3-haiku N/A claude-3-haiku-20240307 Llama llama-3-8b* 4.7 GB Ollama Model ID: 365c0bd3c000 llama-3.1-8b* 4.9 GB Ollama Model ID: 46e0c10c039e Qwen qwen-2-7b* 4.4 GB Ollama Model ID: dd314f039b9d qwen-2.5-7b* 4.7 GB Ollama Model ID: 845dbda0ea48 Gemma gemma-2-9b* 5.4 GB Ollama Model ID: ff02c3702f32 GLM glm-4-9b* 5.5 GB Ollama Model ID: 5b699761eca5 * Open-source LLM (retrieved December 1, 2024).
+    Table 3. Details of selected LLMs (retrieved from (Qiang et al., 2023)). API-accessed models are shown as triangles ($\blacktriangle$) and open-source models as circles ($\bullet$).
+    Family Model Version Type GPT gpt-4o gpt-4o-2024-05-13 $\blacktriangle$ gpt-4o-mini gpt-4o-mini-2024-07-18 $\blacktriangle$ Claude claude-3-sonnet claude-3-sonnet-20240229 $\blacktriangle$ claude-3-haiku claude-3-haiku-20240307 $\blacktriangle$ Llama llama-3-8b Ollama Model ID: 365c0bd3c000 $\bullet$ llama-3.1-8b Ollama Model ID: 46e0c10c039e $\bullet$ Qwen qwen-2-7b Ollama Model ID: dd314f039b9d $\bullet$ qwen-2.5-7b Ollama Model ID: 845dbda0ea48 $\bullet$ Gemma gemma-2-9b Ollama Model ID: ff02c3702f32 $\bullet$ GLM glm-4-9b Ollama Model ID: 5b699761eca5 $\bullet$
   2. (2)
@@
     We run the experiment on the Largebio Track that contains pairwise alignments between FMA, NCI, and SNOMED.
-The true mappings (TPs) are retrieved from the reference (R).
-The false mappings (FPs) are retrieved from the alignment (A) generated by the text preprocessing pipeline but exclude the true mappings (TPs) listed in the reference (R).
+The true mappings (TPs) are retrieved from the reference (R) where the mappings can be found by syntactic OM.
+The false mappings (FPs) are retrieved from the alignment (A) generated by the classic text preprocessing pipeline but exclude TPs above.
   4. (4)
     We report our results over a single run.
-There could be slight differences in the results over multiple trails, but these differences are not significant with respect to the ranking and trend of different LLMs and prompt templates.
+There could be slight differences in the results over multiple trials, but these differences are not significant with respect to the ranking and trend of different LLMs and prompt templates.
 Figure 10. Discovery rate of TPs using LLMs on the Largebio Track.
@@
 For LLMs, we directly prompt LLMs to determine whether two entities are identical or not.
-For the text preprocessing pipeline, we employ the pipeline via function calling (a.k.a tool calling).
+For the text preprocessing pipeline, we employ the pipeline via function calling (OpenAI, 2023) (a.k.a.
+tool calling) in LLMs.
 This means that the LLMs act as an interface, and mapping judgement is based on programme logic defined in the text preprocessing method.
@@
 Figure 10 shows the discovery rate of TPs using LLMs on the Largebio Track.
-Here we use only Tokenisation (T) and Normalisation (N) to preprocess the text because Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are shown to be less effective from our analysis in Section 4.
-Although some API-accessed commercial LLMs have shown the capability to detect TPs without a text preprocessing step, this behaviour is non-deterministic across different LLMs and prompt templates.
+Here we use only Tokenisation (T) and Normalisation (N) to preprocess the text because Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are shown to be less effective in our analysis in Section 4.
+Although some API-accessed commercial LLMs have shown the capability to detect TPs without a text preprocessing step, this behaviour varies across different LLMs and prompt templates.
 For different LLMs, open-source LLMs perform better than commercial API-accessed LLMs.
 It could be problematic for some commercial API-accessed LLMs (e.g.
-claude-3haiku) to find two identical entities if text preprocessing is not applied prior to the comparison.
+claude-3-haiku) to find two identical entities if text preprocessing is not applied prior to the comparison.
 They may consider entities with different formatting that are used in different contexts and therefore have different meanings.
@@
 The combination of few-shot and self-reflection (PT4) may even perform worse than the simple basic prompt (PT1).
-We believe that this phenomenon is mainly caused by LLM “overthinking”, a type of LLM hallucinations in which LLMs could over-analyse user input and generate unnecessarily complex details.
+We believe that this phenomenon is mainly caused by “overthinking”, a type of LLM hallucination in which LLMs could over-analyse user input and generate unnecessarily complex details (Chen et al., 2025).
 We also need to note here that the example generated in the few-shot prompt is based on our observation of the existing TPs.
 Finding a good example in real-world applications is not easy because ontology entities can have different formatting by nature.
-On the other hand, using the classical text preprocessing pipeline via function calling could help LLMs produce a stable result with a high discovery rate (almost 100% from our observation) across different models and prompt templates.
+On the other hand, using the classic text preprocessing pipeline via function calling could help LLMs produce a stable result with a high discovery rate (almost 100% from our observation) across different models and prompt templates.
 6.4. Detecting False Mappings
@@
 The discovery rate is defined as the number of “no” answers compared to the total number of FPs.
-Specifically, these FPs are non-trivial mappings that the classical text preprocessing methods cannot detect.
+Specifically, these FPs are non-trivial mappings that cannot be detected by the classic text preprocessing method.
 Figure 11 shows the discovery rate of FPs using LLMs on the Largebio Track.
@@
 SL has shown a lower precision compared to other text preprocessing methods, which is caused by the method producing a significant number of FPs.
-LLMs have shown a strong capability to discover the FPs produced by the classical text preprocessing pipeline.
+LLMs have shown a strong capability to discover the FPs produced by the classic text preprocessing pipeline.
 For different models, API-accessed commercial LLMs are better than open-source LLMs and reach a higher discovery rate of more than 80%.
@@
 OM is not only a QA task, but also a knowledge-intensive task that requires a comprehensive understanding and reasoning of domain knowledge.
-The studies in this section demonstrate the limitations and opportunities of integrating the text preprocessing pipeline with LLMs in syntactic OM.
+The studies in this section demonstrate the limitations and opportunities for integrating the text preprocessing pipeline with LLMs in syntactic OM.
 We highlight two pathways for this integration.
-One pathway is to use function calling to invoke the classical text preprocessing pipeline within LLMs, which packages the programming-based text preprocessing pipeline as a tool for LLMs to use.
+One pathway is to use function calling to invoke the classic text preprocessing pipeline within LLMs, which packages the programming-based text preprocessing pipeline as a tool for LLMs to use.
 This provides stable results for text preprocessing in syntactic OM.
-The other pathway is to use LLMs as a repair tool after text preprocessing, which could be a cost-effective solution for post hoc correction.
-This can also be extended using human-in-the-loop to validate the repaired results generated by LLMs and feed the validated data back into LLMs so that they can better understand the context of the task (i.e.
-fine-tuning LLMs used in syntactic OM).
+The other pathway is to use LLMs as a repair tool after text preprocessing, which could be a cost-effective solution for post hoc correction to reduce the number of FPs.
+This can also be extended using human-in-the-loop to validate the repaired results generated by LLMs and to feed the validated data back into LLMs so that they can better understand the context of the task (i.e.
+fine-tuning LLMs for syntactic OM).
 7. Discussion
@@
 Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching.
-This study concentrates on a common practice using the text preprocessing pipeline in syntactic OM.
+This study concentrates on the common practice of using the text preprocessing pipeline in syntactic OM.
 While generic text preprocessing pipeline methods in syntactic OM are usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
@@
     Our study shows (a) whether to use, or not use, and (b) how to use these text preprocessing methods.
-Such experimental results will benefit decision-making when selecting appropriate text preprocessing methods for syntactic OM.
+Such experimental results will benefit the selection of appropriate text preprocessing methods for syntactic OM.
 For large-scale OM, it can significantly reduce unnecessary trial costs.
@@
 If two ontology entities can be matched using the intuitive text preprocessing pipeline, there is no need to leverage LLMs for the same purpose.
-  4. (4)
 The development and implementation of an LLM-based framework that incorporates different matching techniques is a new research direction for OM, but it is beyond the scope of this paper.
@@
   4. (4)
-    This study also demonstrates that the nature of text preprocessing used in NLP applications can be context-based.
+    This study demonstrates that the nature of text preprocessing used in NLP applications can be context-based.
 Not all pipeline methods are effective, and some of them may even hamper the performance of downstream tasks.
@@
 Only a few OAEI tracks contain subsumption mappings, and they are not enough to robustly demonstrate the text preprocessing used in subsumption mappings.
-This work has not been included in the official OAEI campaign because of its intense focus on only text preprocessing in syntactic OM.
+This work has not been included in the official OAEI campaign because of its intense focus on text preprocessing in syntactic OM.
 9. Conclusion
@@
 F1 score).
-(3) We also argue that the traditional text preprocessing pipeline is not obsolete in the era of LLMs.
+(3) We argue that the traditional text preprocessing pipeline is not obsolete in the era of LLMs.
 It is an integral part of the LLMs used in OM tasks.
-Integrating the classical text preprocessing pipeline can significantly improve the robustness of purely prompt-based approaches.
+Integrating the classic text preprocessing pipeline can significantly improve the robustness of purely prompt-based approaches.
 Our future work will focus on handling class axioms and complex relationships to evaluate the text preprocessing pipeline.
@@
 The authors thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.
+According to the OAEI data policy (retrieved January 1, 2025), “OAEI results and datasets, are publicly available, but subject to a use policy similar to the one defined by NIST for TREC.
+These rules apply to anyone using these data.” Please find more details from the official website: https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.

### v6

@@
-How Does A Text Preprocessing Pipeline Affect Syntactic Ontology Matching? [Experiment, Analysis & Benchmark]
+How Does A Text Preprocessing Pipeline Affect Syntactic Ontology Matching?
 Abstract.
@@
 However, the lack of standardisation in text preprocessing creates diversity in mapping results.
-In this paper we investigate the effect of the text preprocessing pipeline on syntactic OM in 8 Ontology Alignment Evaluation Initiative (OAEI) tracks with 49 distinct alignments.
+In this paper, we investigate the effect of the text preprocessing pipeline on syntactic OM in 8 Ontology Alignment Evaluation Initiative (OAEI) tracks with 49 distinct alignments.
 We find that Phase 1 text preprocessing (Tokenisation and Normalisation) is more effective than Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation).
-To repair the unwanted false mappings caused by Phase 2 text preprocessing, we propose a novel context-based pipeline repair approach that employs a post hoc check to find common words that cause false mappings.
-These words are stored in a reserved word set and applied in text preprocessing.
-The experimental results show that our approach improves the matching correctness and the overall matching performance.
-We then consider the broader integration of the classic text preprocessing pipeline with modern large language models (LLMs) for OM.
-We recommend that (1) the text preprocessing pipeline be injected via function calling into LLMs to avoid the tendency towards unstable true mappings produced by LLM prompting; or (2) LLMs be used to repair non-existent and counter-intuitive false mappings generated by the text preprocessing pipeline.
-PVLDB Reference Format:
-PVLDB, 19(1): XXX-XXX, 2025.
-doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
-Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
-For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
-Copyright is held by the owner/author(s).
-Publication rights licensed to the VLDB Endowment.
-Proceedings of the VLDB Endowment, Vol. 19, No. 1 ISSN 2150-8097.
-doi:XX.XX/XXX.XX
-PVLDB Artifact Availability:
-The source code, data, and/or other artifacts have been made available at https://github.com/qzc438-research/ontology-nlp.
+We propose two novel approaches to repair unwanted false mappings caused by Phase 2 text preprocessing.
+One is an ad hoc logic-based repair approach that employs an ontology-specific check to find common words that cause false mappings.
+These words are stored in a reserved word set and applied before the text preprocessing.
+By leveraging the power of large language models (LLMs), we also propose a post hoc LLM-based repair approach.
+This approach utilises the strong background knowledge provided by LLMs to repair non-existent and counter-intuitive false mappings after the text preprocessing.
+It also overcomes the tendency towards unstable true mappings by injecting the classical text preprocessing pipeline via function calling.
+The experimental results show that these two approaches can improve the matching correctness and the overall matching performance.
+Ontology Matching, Text Preprocessing
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 29th International Conference on Extending Database Technology; March 24–27, 2026; Tampere, Finlandisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Ontology engineering
 1. Introduction
@@
 Despite the diversity of the models used, they all apply the text preprocessing pipeline to clean the text data before using the model.
-The pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L) (Manning et al., 2008).
-Table 1 shows examples of using the text preprocessing pipeline to find mappings in syntactic OM.
+The classical text preprocessing pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L) (Manning et al., 2008).
 T is the process of breaking the text into the smallest units (i.e.
@@
 S/L is used to deal with the grammatical variation of words, applying rules to find the simplest common form of the word.
-Pipeline Ontology Entity Name Output T cmt ProgramCommittee Program Committee sigkdd Program_Committee N cmt ProgramCommittee program committee conference Program_committee R conference has_a_review review ekaw hasReview S/L conference reviewes review edas (is)Reviewing
-Table 1.
-Examples of using the text preprocessing pipeline in syntactic OM: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
-In the examples, we use whitespace to split the tokens, normalise all the words into lowercase, remove the common stop words “has” and “a”, and apply S/L to find the simplest word form.
-While a number of OM systems use the text preprocessing pipeline for syntactic OM, few studies question its role in optimising the matching performance.
+Table 1 shows examples of using the text preprocessing pipeline to find mappings in syntactic OM.
+In the examples, we use whitespace to split the tokens, normalise all the words into lowercase, remove the common stop words “has”, “a” and “is”, and apply stemming/lemmatisation to find the simplest word form “review” from “reviews” and “reviewing”.
+No.
+Entity Name Matching Output Preprocessing 1 ProgramCommittee Program Committee T Program_Committee 2 ProgramCommittee program committee T + N Program_committee 3 has_a_review review T + N + R hasReview 4 reviews review T + N + R + S/L isReviewing
+Table 1. Examples of using the classical text preprocessing pipeline in syntactic ontology matching: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
+A number of OM systems use the text preprocessing pipeline for syntactic OM, but few studies question its role in optimising the matching performance.
 Our study fills this gap.
-Firstly, we conduct a comprehensive experimental analysis of the text preprocessing pipeline in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline.
+We first conduct a comprehensive experimental analysis of the text preprocessing pipeline used in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline.
 In each step, text preprocessing methods are evaluated for correctness and completeness.
-Secondly, we propose a novel context-based pipeline repair approach for syntactic OM.
-The method offers a customised way to fine-tune the text preprocessing pipeline for each domain-specific OM.
-This method shows promising results for repairing less effective text preprocessing methods.
-Thirdly, we discuss the non-deterministic nature of modern large language models (LLMs) for text preprocessing and highlight the importance of integrating the classic text preprocessing pipeline with LLM-based OM.
+While several steps of the text preprocessing pipeline are currently less effective, we propose two novel pipeline repair approaches.
+These two methods utilise the logic of ontological representations and the power of large language models (LLMs), showing promising results for repairing the less effective text preprocessing pipeline.
 Specifically, this paper makes the following contributions:
-  • (Section 4) We compare the performance of the text preprocessing pipeline on 8 Ontology Alignment Evaluation Initiative (OAEI) (OAEI Community, [n.d.]) datasets with 49 distinct alignments.
+  • We compare the performance of the text preprocessing pipeline on 8 Ontology Alignment Evaluation Initiative (OAEI) (OAEI Community, [n.
+d.]) datasets with 49 distinct alignments.
 While there is a significant improvement using Tokenisation (T) and Normalisation (N) in syntactic OM, Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are currently less effective.
-We find inappropriate stop words removal, over-stemming, and over-lemmatisation.
-We also compare the performance of several text preprocessing methods: (1) Stemming and Lemmatisation, (2) different stemming methods (Porter, Snowball, and Lancaster), and (3) Lemmatisation with and without Part-of-Speech (POS) Tagging.
-  • (Section 5) We propose a simple and intuitive context-based pipeline repair method using a reserved word set to capture certain words that may lead to false mappings.
-This method is evaluated on the same OAEI datasets we analysed, showing promising results in improving the matching correctness and the overall matching performance of syntactic OM.
-  • (Section 6) We discuss the integration of the classic text preprocessing pipeline with modern LLMs.
-We recommend using text preprocessing via function calling to overcome the non-determinism of purely prompt-based approaches to finding true mappings, while LLMs show great potential to repair false mappings generated by the text preprocessing pipeline.
+We find common inappropriate stop words removal, over-stemming, and over-lemmatisation.
+  • We propose two approaches to repair the less effective text preprocessing pipeline in syntactic OM: (1) ad hoc logic-based repair and (2) post hoc LLM-based repair.
+Experimental results show that these two methods can improve the matching correctness and the overall matching performance of syntactic OM.
 The remainder of the paper is organised as follows.
-Section 2 and 3 review the related work and the preliminaries of OM task formulation.
+Section 2 and 3 review the related work and the preliminaries of OM and syntactic matching.
 Section 4 analyses the text preprocessing pipeline used in syntactic OM.
-Section 5 proposes the context-based pipeline repair approach and experimentally validates its performance.
-Section 6 demonstrates the integration of the classic text preprocessing pipeline with modern LLMs in syntactic OM.
-Section 7 and 8 discuss the findings and limitations.
-Section 9 concludes.
+Section 5 proposes text preprocessing pipeline repair approaches and experimentally validates their performance.
+Section 6 highlights the contributions of this work and concludes the paper.
 2. Related Work
@@
 However, to the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have evaluated the effect of the entire text preprocessing pipeline on syntactic OM.
-3. Preliminaries: OM Task Formulation
+3. Preliminaries
 Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), OM establishes mappings between pairs of entities drawn from each of the two ontologies.
@@
     Table 2. Selected OAEI tracks and the number of alignments.
-    Name Domain Number of Alignments Anatomy Human and Mouse Anatomy 1 Biodiv Biodiversity and Ecology 9 CommonKG Common Knowledge Graphs 3 Conference Conference 24 Food Food Nutritional Composition 1 Largebio Biomedical 6 MSE Materials Science & Engineering 3 Phenotype Disease and Phenotype 2
-    Compound words are frequently used in ontology naming conventions.
-For example, compound words such as “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO (Suchanek et al., 2007) or the Snake case “art_gallery” used in Wikidata (Vrandečić and Krötzsch, 2014).
-Figure 2 shows the proportion of compound words in the selected OAEI tracks.
-For comparisons between entities with compound words, one approach is to use their tokens, where “XY” is equivalent to “YX” because they share the same concepts X and Y.
-However, our comparison considers both concepts and their order, so that “XY” is not equivalent to “YX”.
-Although concurrent “XY” and “YX” are rare in the matching process, we assume that it could happen and so we take account of the order of concepts in compound word names to ensure a fair comparison in our experiments.
-    Figure 2. Proportion of compound words.
+    Name Domain Alignments Anatomy Human and Mouse Anatomy 1 Biodiv Biodiversity and Ecology 9 CommonKG Common Knowledge Graphs 3 Conference Conference 24 Food Food Nutritional Composition 1 Largebio Biomedical 6 MSE Materials Science & Engineering 3 Phenotype Disease and Phenotype 2
   2. (2)
@@
 In the context of OM, the evaluation compares an alignment (A) returned by the OM system with a gold standard reference (R) validated by domain experts.
-Figure 3 illustrates that the four primitive measures in OM can be interpreted as $TP=A\cap R$, $FP=A-R$, $FN=R-A$, and $TN=(C\times C^{\prime})-(A\cup R)$, where $C\times C^{\prime}$ refers to all possible correspondences $\{(e_{1},e_{2},\equiv)|e_{1}\in O_{s},e_{2}\in O_{t}\}$.
+Figure 2 illustrates that the four primitive measures in OM can be interpreted as $TP=A\cap R$, $FP=A-R$, $FN=R-A$, and $TN=(C\times C^{\prime})-(A\cup R)$, where $C\times C^{\prime}$ refers to all possible correspondences.
 Accuracy (Acc), Specificity (Spec), Precision (Prec), Recall (Rec), and $F_{\beta}$ Score ($F_{\beta}$) are the most common evaluation measures based on TP, FP, FN, and TN.
@@
     $F_{1}\ Score=\frac{2}{Prec^{-1}+Rec^{-1}}$
+Compound words are frequently used in ontology naming conventions.
+For example, compound words such as “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO (Suchanek et al., 2007) or the Snake case “art_gallery” used in Wikidata (Vrandečić and Krötzsch, 2014).
+Figure 3 shows the proportion of compound words in the selected OAEI tracks.
+For comparisons between entities with compound words, one approach is to use their tokens, where “XY” is equivalent to “YX” because they share the same concepts X and Y.
+However, our comparison considers both concepts and their order, so that “XY” is not equivalent to “YX”.
+Although concurrent “XY” and “YX” are rare in the matching process, we assume that it could happen and so we take account of the order of concepts in compound word names to ensure a fair comparison in our experiments.
+Figure 3. Proportion of compound words. We exclude entities with more than 15 number of words because their proportion is less than 1%.
 4.2. Results
-4.2.1. Comparison of Text Preprocessing Pipeline
-Figure 4 compares the stages of the text preprocessing pipeline.
+Figure 4 compares each step of the text preprocessing pipeline.
+These steps are always applied sequentially in the pipeline.
 Beginning with $\varnothing$ (no text preprocessing), we see the performance of subsequent Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
-The methods are always applied sequentially in the pipeline.
 For T and N, most of the data points above the equivalence line in precision, recall, and F1 score indicate that they can help syntactic OM.
 For R and S/L, some data points are above the equivalence line in recall, but most data points are below the equivalence line in precision and F1 score indicating that they do not play a positive role.
-Three horizontal lines inside each violin plot show three quartiles: Q1, median, and Q3.
-The extension of the curves beyond 0% and 100% is an artefact of the Seaborn (Waskom, 2021) package.
 In the violin plots: (a) Precision: The median increases with T and N but decreases with R and S/L.
@@
 Comparison of text preprocessing pipeline: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
-The symbol $\varnothing$ represents the baseline without text preprocessing.
-4.2.2. Stemming vs. Lemmatisation
+In the violin plots, three horizontal lines inside each violin plot show three quartiles: Q1, median, and Q3 (an artefact of the Seaborn (Waskom, 2021) package).
+  1. (1)
     Figure 5 compares Stemming (S) and Lemmatisation (L) after Tokenisation (T) and Normalisation (N) have been applied.
@@
     Figure 5. Stemming (S) vs. Lemmatisation (L).
-4.2.3. Different Stemming Methods
+  2. (2)
     Figure 6 compares different stemming methods: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL).
@@
     Figure 6. Comparison of different stemmers: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL).
-4.2.4. Lemmatisation with/without POS Tagging
+  3. (3)
     Figure 7 compares Lemmatisation (L) and Lemmatisation + POS Tagging (LT).
@@
 singularity, positive names, nouns for classes and verbs for properties (Schober et al., 2007; Taylor et al., 2015).
-5. Context-Based Pipeline Repair
-The experimental results in Section 4 demonstrate that only Phase 1 text processing (Tokenisation and Normalisation) helps with syntactic OM.
+5. Text Preprocessing Pipeline Repair
+The experimental results in Section 4 demonstrate that only Phase 1 text preprocessing (Tokenisation and Normalisation) helps with syntactic OM.
 The use of Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation) does not improve performance and may even have negative impacts.
-In this section, we explore the underlying cause of this behaviour and introduce an intuitive context-based pipeline repair approach.
+In this section, we explore the underlying cause of this behaviour and introduce our novel pipeline repair approaches.
 5.1. Causal Analysis
@@
 This idea inspires our novel pipeline repair method.
-5.2. Repair Method
+5.2. Motivation
 While Phase 2 text preprocessing methods have been shown to be less effective in syntactic OM by introducing FPs, we propose a pipeline repair approach to differentiate FPs from TPs.
@@
 The number of TPs increases, but the number of FPs also increases.
-(3) Our context-based pipeline repair aims to collapse Alignment (A).
+(3) Our text preprocessing pipeline repair aims to collapse Alignment (A).
 The number of FPs decreases significantly, with a slight decrease in TPs.
 Figure 8. Motivation of our context-based pipeline repair.
+5.3. Ad Hoc Logic-Based Repair
+The ad hoc logic-based repair applies before the text preprocessing.
 One critical step in our approach is to assemble a set of reserved words that may cause FPs in text preprocessing, and to exclude these words from text preprocessing.
-The selection criteria are based on two widely agreed assumptions: (1) there are no duplicate entities within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies.
-We call our approach context-based here because pairs of words may have the same or different meanings in different contexts.
+The selection criteria are based on two widely agreed assumptions: (1) there are no duplicate entities within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies (OBO Foundry, [n.
+d.]).
+This approach is context-based because pairs of words may have the same or different meanings in different contexts.
 Based on these two assumptions, we propose a simple Algorithm 1 to assemble the reserved word set.
@@
 As the words “has”, “a”, “was”, and “of” are listed as reserved words, these two named properties are processed as “has a steer committe” and “was a steer committe of”, respectively.
-5.3. Evaluation
-We apply our context-based pipeline repair approach to the same OAEI datasets as above.
-Figure 9 compares with and without context-based pipeline repair in 8 tracks with 49 distinct alignments.
-The text preprocessing pipeline methods implemented include all Phase 2 text preprocessing methods: Stop Words Removal (R), Porter Stemmer (SP), Snowball Stemmer (SS), Lancaster Stemmer (SL), Lemmatisation (L), and Lemmatisation + POS Tagging (LT).
-Figure 9.
-Context-based pipeline repair use in Phase 2 text preprocessing: Stop Words Removal (R), Porter Stemmer (SP), Snowball Stemmer (SS), Lancaster Stemmer (SL), Lemmatisation (L), and Lemmatisation + POS Tagging (LT).
+5.4. Post Hoc LLM-Based Repair
+With strong background knowledge, LLMs can be helpful in detecting false mappings after text preprocessing.
+However, there is a wide range of LLMs with different prompting strategies that can be used in LLM-based repair.
+We aim to find a generic pipeline repair approach and the best practice to integrate LLMs in the repair phase.
+The experiment settings are as follows.
   1. (1)
-    Precision: Most data points are above the equivalence line, indicating that context-based pipeline repair significantly improves matching correctness.
-Recall: Most data points are located on the equivalence line, and only a few data points are below the equivalence line, indicating that context-based pipeline repair slightly reduces matching completeness.
-F1 Score: Most of the data points are above the equivalence line, indicating that context-based pipeline repair also improves overall matching performance.
-  2. (2)
-    Experimentally, the wider ellipse around the data points indicates that the matching performance improvement ranking in Phase 2 text preprocessing methods is Stemming (S) > Lemmatisation (L) > Stop Words Removal (R); the matching performance improvement ranking in different stemmers is Lancaster Stemmer (SL) > Porter Stemmer (SP) = Snowball Stemmer (SS); and the matching performance improvement ranking in lemmatisation is Lemmatisation (L) $\approx$ Lemmatisation + POS Tagging (LT).
-6. Text Preprocessing Pipeline & LLMs
-In this section, we discuss the pros and cons of LLMs in performing syntactic OM and demonstrate the benefits of integrating the classic text preprocessing pipeline with modern LLMs.
-6.1. Motivation
-In modern LLMs, text preprocessing can be performed implicitly within the models.
-For example, LLMs may understand that “art gallery” and “ArtGallery” are true mappings even if they are in different formats.
-However, it is questionable whether this behaviour is stable across different LLMs and prompt templates and whether LLMs outperform the classic text processing pipeline.
-On the other hand, with a strong background knowledge, LLMs may be able to facilitate mapping repair together with the text preprocessing pipeline.
-For example, LLMs can classify “Bile” and “Bilateral” as false mappings, whereas the classic text preprocessing method stemming cannot.
-6.2. Experiment Setup
+    We run the experiment on the Largebio Track that contains pairwise alignments between FMA, NCI, and SNOMED. We choose three datasets: WHOLE-FMA-NCI, WHOLE-FMA-SNOMED, and WHOLE-SNOMED-NCI.
   2. (2)
     We choose 10 LLMs from 4 different families.
-These include two OpenAI GPT models (OpenAI, [n.d.]) (gpt-4o and gpt-4o-mini), two Anthropic Claude models (Anthropic, [n.d.])(claude-3-sonnet and claude-3-haiku), and two versions of Meta Llama (Meta, [n.d.]) (llama-3-8b and llama-3.1-8b), two versions of Alibaba Qwen (Alibaba Qwen Team, [n.d.]) (qwen-2-7b and qwen-2.5-7b), Google Gemma 2 (Google Gemma Team and Google DeepMind, [n.d.]), and ChatGLM 4 (Zeng et al., 2024).
-We use LangChain (LangChain, Inc., [n.d.]) to build chats with OpenAI GPT and Anthropic Claude models, and Ollama (Ollama, [n.d.]) to access open-source models.
+These include two OpenAI GPT models (OpenAI, [n.
+d.]) (gpt-4o and gpt-4o-mini), two Anthropic Claude models (Anthropic, [n.
+d.])(claude-3-sonnet and claude-3-haiku), and two versions of Meta Llama (Meta, [n.
+d.]) (llama-3-8b and llama-3.1-8b), two versions of Alibaba Qwen (Alibaba Qwen Team, [n.
+d.]) (qwen-2-7b and qwen-2.5-7b), Google Gemma 2 (Google Gemma Team and Google DeepMind, [n.
+d.]), and ChatGLM 4 (Zeng et al., 2024).
+We use LangChain (LangChain, Inc., [n.
+d.]) to build chats with OpenAI GPT and Anthropic Claude models, and Ollama (Ollama, [n.
+d.]) to access open-source models.
 Table 3 lists the details of selected LLMs.
@@
     Table 3. Details of selected LLMs (retrieved from (Qiang et al., 2023)). API-accessed models are shown as triangles ($\blacktriangle$) and open-source models as circles ($\bullet$).
-    Family Model Version Type GPT gpt-4o gpt-4o-2024-05-13 $\blacktriangle$ gpt-4o-mini gpt-4o-mini-2024-07-18 $\blacktriangle$ Claude claude-3-sonnet claude-3-sonnet-20240229 $\blacktriangle$ claude-3-haiku claude-3-haiku-20240307 $\blacktriangle$ Llama llama-3-8b Ollama Model ID: 365c0bd3c000 $\bullet$ llama-3.1-8b Ollama Model ID: 46e0c10c039e $\bullet$ Qwen qwen-2-7b Ollama Model ID: dd314f039b9d $\bullet$ qwen-2.5-7b Ollama Model ID: 845dbda0ea48 $\bullet$ Gemma gemma-2-9b Ollama Model ID: ff02c3702f32 $\bullet$ GLM glm-4-9b Ollama Model ID: 5b699761eca5 $\bullet$
+    Family Model Version GPT $\blacktriangle$ gpt-4o gpt-4o-2024-05-13 $\blacktriangle$ gpt-4o-mini gpt-4o-mini-2024-07-18 Claude $\blacktriangle$ claude-3-sonnet claude-3-sonnet-20240229 $\blacktriangle$ claude-3-haiku claude-3-haiku-20240307 Llama $\bullet$  llama-3-8b Ollama Model ID: 365c0bd3c000 $\bullet$  llama-3.1-8b Ollama Model ID: 46e0c10c039e Qwen $\bullet$  qwen-2-7b Ollama Model ID: dd314f039b9d $\bullet$  qwen-2.5-7b Ollama Model ID: 845dbda0ea48 Gemma $\bullet$  gemma-2-9b Ollama Model ID: ff02c3702f32 GLM $\bullet$  glm-4-9b Ollama Model ID: 5b699761eca5
   3. (3)
@@
 (a) Basic: The simplest prompt template to compare two entities using “Is Entity1 equivalent to Entity2?
-Answer yes or no.” (b) Few-Shot: Adding an example in the prompt, such as “Example: Hair_root is equivalent to Hair_Root.” (c) Self-Reflection: Before LLMs answer the question, ask LLMs to rethink the output and give a short explanation.
+Answer yes or no.” (b) Few-Shot: Adding an example in the prompt.
+We use the example “Example: Hair_root is equivalent to Hair_Root.” (c) Self-Reflection: Before LLMs answer the question, ask LLMs to rethink the output and give a short explanation.
 We add a sentence “Write a short explanation.” to invoke this function.
@@
 Self-Reflection Write a short explanation.
+    Figure 9. Discovery rate of TPs using LLMs on the Largebio Track.
+    Figure 10. The discovery rate of FPs using LLMs on the Largebio Track.
   4. (4)
-    We run the experiment on the Largebio Track that contains pairwise alignments between FMA, NCI, and SNOMED.
+    We compare LLMs with the classical text preprocessing pipeline in detecting TPs and FPs, which are two main contributors to precision and recall in matching performance.
 The true mappings (TPs) are retrieved from the reference (R) where the mappings can be found by syntactic OM.
@@
 There could be slight differences in the results over multiple trials, but these differences are not significant with respect to the ranking and trend of different LLMs and prompt templates.
-Figure 10. Discovery rate of TPs using LLMs on the Largebio Track.
-Figure 11. The discovery rate of FPs using LLMs on the Largebio Track.
-6.3. Detecting True Mappings
-LLMs are expected to answer “yes” for TPs.
+Figure 9 compares LLMs with the classical text preprocessing pipeline in detecting TPs.
+LLMs and the text preprocessing pipeline are expected to answer “yes” for TPs.
 The discovery rate is defined as the number of “yes” answers compared to the total number of TPs.
 For LLMs, we directly prompt LLMs to determine whether two entities are identical or not.
-For the text preprocessing pipeline, we employ the pipeline via function calling (OpenAI, 2023) (a.k.a.
-tool calling) in LLMs.
-This means that the LLMs act as an interface, and mapping judgement is based on programme logic defined in the text preprocessing method.
-Two entities are considered to be matched if they have the same results after the text preprocessing.
-Figure 10 shows the discovery rate of TPs using LLMs on the Largebio Track.
+For the text preprocessing pipeline, two entities are considered to be matched if they have the same results after the text preprocessing.
 Here we use only Tokenisation (T) and Normalisation (N) to preprocess the text because Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are shown to be less effective in our analysis in Section 4.
@@
 The combination of few-shot and self-reflection (PT4) may even perform worse than the simple basic prompt (PT1).
-We believe that this phenomenon is mainly caused by “overthinking”, a type of LLM hallucination in which LLMs could over-analyse user input and generate unnecessarily complex details (Chen et al., 2025).
+We believe that this phenomenon is mainly attributed to “overthinking”, a type of LLM hallucination in which LLMs could over-analyse user input and generate unnecessarily complex details (Chen et al., 2025).
 We also need to note here that the example generated in the few-shot prompt is based on our observation of the existing TPs.
@@
 On the other hand, using the classic text preprocessing pipeline via function calling could help LLMs produce a stable result with a high discovery rate (almost 100% from our observation) across different models and prompt templates.
-6.4. Detecting False Mappings
-LLMs are expected to answer “no” for FPs.
+Figure 10 compares LLMs with the classical text preprocessing pipeline in detecting FPs.
+LLMs and the text preprocessing pipeline are expected to answer “no” for FPs.
 The discovery rate is defined as the number of “no” answers compared to the total number of FPs.
-Specifically, these FPs are non-trivial mappings that cannot be detected by the classic text preprocessing method.
-Figure 11 shows the discovery rate of FPs using LLMs on the Largebio Track.
+Specifically, these FPs evaluated are non-trivial mappings that cannot be detected by the classic text preprocessing method.
 Here we use LLMs to detect FPs generated by the Lancaster Stemmer (SL).
@@
 While mapping repair is usually non-trivial and requires domain expertise, LLMs have shown potential for performing a high-quality post hoc correction without human intervention.
-6.5. Implications
-LLMs were originally developed and trained for question answering (QA).
-OM is not only a QA task, but also a knowledge-intensive task that requires a comprehensive understanding and reasoning of domain knowledge.
-The studies in this section demonstrate the limitations and opportunities for integrating the text preprocessing pipeline with LLMs in syntactic OM.
-We highlight two pathways for this integration.
-One pathway is to use function calling to invoke the classic text preprocessing pipeline within LLMs, which packages the programming-based text preprocessing pipeline as a tool for LLMs to use.
-This provides stable results for text preprocessing in syntactic OM.
-The other pathway is to use LLMs as a repair tool after text preprocessing, which could be a cost-effective solution for post hoc correction to reduce the number of FPs.
-This can also be extended using human-in-the-loop to validate the repaired results generated by LLMs and to feed the validated data back into LLMs so that they can better understand the context of the task (i.e.
-fine-tuning LLMs for syntactic OM).
-7. Discussion
-OM aims to bridge the gap by capturing similar concepts that occur in different ontologies.
-OM is a complex process that combines syntactic matching, lexical matching, and semantic matching.
-Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching.
-This study concentrates on the common practice of using the text preprocessing pipeline in syntactic OM.
-While generic text preprocessing pipeline methods in syntactic OM are usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
-  1. (1)
-    Our study shows (a) whether to use, or not use, and (b) how to use these text preprocessing methods.
-Such experimental results will benefit the selection of appropriate text preprocessing methods for syntactic OM.
-For large-scale OM, it can significantly reduce unnecessary trial costs.
-  2. (2)
-    Our context-based pipeline repair approach is proposed to repair the less effective Phase 2 text preprocessing methods.
-Its broader value is showing how to maximise true mappings and minimise false mappings throughout the text preprocessing pipeline.
-From a statistical perspective, it is a local optimisation for syntactic matching that will benefit the global optimisation for OM, as syntactic matching provides “anchor mappings” for lexical matching and semantic matching.
-  3. (3)
-    The use of LLMs is not a one-size-fits-all solution for OM tasks.
-OM is a complex multi-level matching mixed with different matching techniques.
-LLMs may not outperform traditional methods in performing these sub-matching tasks.
-We suggest “rules of thumb” for syntactic OM.
-If two ontology entities can be matched using the intuitive text preprocessing pipeline, there is no need to leverage LLMs for the same purpose.
-The development and implementation of an LLM-based framework that incorporates different matching techniques is a new research direction for OM, but it is beyond the scope of this paper.
-We refer the reader to several recent works in (He et al., 2023; Norouzi et al., 2023; Hertling and Paulheim, 2023; Qiang et al., 2023; Giglou et al., 2024; Amini et al., 2024).
-  4. (4)
-    This study demonstrates that the nature of text preprocessing used in NLP applications can be context-based.
-Not all pipeline methods are effective, and some of them may even hamper the performance of downstream tasks.
-Modifications are required before applying text preprocessing to specific NLP tasks.
-8. Limitations
-The references contained in the OAEI tracks are not “complete” gold standards and some need further development.
-In this study, we only deal with equivalence mappings between classes and properties.
-Subsumption mappings are not considered.
-Finding widely used baseline data with ground truth matches for subsumptions is empirically difficult.
-Only a few OAEI tracks contain subsumption mappings, and they are not enough to robustly demonstrate the text preprocessing used in subsumption mappings.
-This work has not been included in the official OAEI campaign because of its intense focus on text preprocessing in syntactic OM.
+For syntactic OM, we can see that LLMs are not good at detecting true mappings, but they are good at detecting false mappings.
+Therefore, we propose our generic post hoc LLM-based repair with (1) employing the classical text preprocessing pipeline via function calling (OpenAI, 2023) (a.k.a.
+tool calling) to detect TPs.
+This means that the LLMs act as an interface, and the mapping judgement is based on the programme logic defined in the text preprocessing method.
+This provides stable results for text preprocessing when finding FPs.
+LLM prompting is not suitable for this task, even with complex prompting strategies.
+(2) Using LLM prompting to detect FPs.
+The prompt template can be simple like “Is A equivalent to B?”, as we cannot see significant improvement using complex prompting strategies (except Meta Llama models).
+5.5. Evaluation & Ablation Study
+We apply our pipeline repair approaches on the Phase 2 text preprocessing method Lancaster Stemmer (SL), and evaluate their performance on the three large datasets in the Largebio Track: WHOLE-FMA-NCI, WHOLE-FMA-SNOMED, and WHOLE-SNOMED-NCI.
+Figure 11 compares the matching performance with/without using the pipeline approaches: ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of the above (CR).
+We can see that ad hoc logic-based repair (AR) and post hoc LLM-based repair (PR) can significantly improve precision, with only a slight decrease in recall, but the overall F1 score also increases.
+Although there is no restriction on the use of both ad hoc and post hoc repairs, the combination does not improve precision, recall, or F1 score.
+For syntactic OM, its main purpose is to provide “seed mappings” for lexical and semantic matching.
+These mappings are meant to be precise and accurate.
+Both of our pipeline repair approaches have shown high precision and are the best tool suites.
+(a) WHOLE-FMA-NCI
+(b) WHOLE-FMA-SNOMED
+(c) WHOLE-SNOMED-NCI
+Figure 11.
+Evaluation of pipeline repair approaches: Ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of above (CR).
+The symbol $\varnothing$ represents the baseline without text preprocessing.
+Figure 12 compares post hoc LLM-based repair (PR) with two alternative repair approaches.
+The results show that PR achieves almost the same precision with text preprocessing only repair (ToR) and LLM only repair (LoR), but its recall and F1 score are better than the other two alternatives.
+For syntactic OM, this means that our approach could detect more ”seed mappings”.
+(a) WHOLE-FMA-NCI
+(b) WHOLE-FMA-SNOMED
+(c) WHOLE-SNOMED-NCI
+Figure 12. Comparison of post hoc LLM-based repair (PR) with two alternative repair approaches: text preprocessing only repair (ToR) and LLM only repair (LoR).
 6. Conclusion
@@
 precision) is relatively low.
-(2) We propose a novel context-based pipeline repair approach to repair the less effective Phase 2 text preprocessing methods.
-By using a reserved word set to reduce false positive samples detected, our approach outperforms the traditional text preprocessing pipeline, in particular, the matching correctness (i.e.
+(2) We propose two novel pipeline repair methods to repair the less effective Phase 2 text preprocessings.
+Both approaches outperform the traditional text preprocessing pipeline without repair, in particular, the matching correctness (i.e.
 precision) and overall matching performance (i.e.
 F1 score).
-(3) We argue that the traditional text preprocessing pipeline is not obsolete in the era of LLMs.
-It is an integral part of the LLMs used in OM tasks.
-Integrating the classic text preprocessing pipeline can significantly improve the robustness of purely prompt-based approaches.
+OM aims to bridge the gap by capturing similar concepts that occur in different ontologies.
+OM is a complex process that combines syntactic matching, lexical matching, and semantic matching.
+Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching.
+This study concentrates on the common practice of using the text preprocessing pipeline in syntactic OM.
+While generic text preprocessing pipeline methods in syntactic OM are usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
+  1. (1)
+    Our study shows (a) whether to use, or not use, and (b) how to use these text preprocessing methods.
+Such experimental results will benefit the selection of appropriate text preprocessing methods for syntactic OM.
+For large-scale OM, it can significantly reduce unnecessary trial costs.
+  2. (2)
+    Our pipeline repair approaches can be used to repair the less effective Phase 2 text preprocessing methods.
+Its broader value is showing how to maximise true mappings and minimise false mappings throughout the text preprocessing pipeline.
+From a statistical perspective, it is a local optimisation for syntactic matching that will benefit the global optimisation for OM, as syntactic matching provides “anchor mappings” for lexical and semantic matching.
+  3. (3)
+    This study demonstrates that the nature of text preprocessing used in NLP applications can be context-based.
+Not all pipeline methods are effective, and some of them may even hamper the performance of downstream tasks.
+Modifications are required before applying text preprocessing to specific NLP tasks.
     Our future work will focus on handling class axioms and complex relationships to evaluate the text preprocessing pipeline.
@@
 The authors thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.
-According to the OAEI data policy (retrieved January 1, 2025), “OAEI results and datasets, are publicly available, but subject to a use policy similar to the one defined by NIST for TREC.
-These rules apply to anyone using these data.” Please find more details from the official website: https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.
+Limitations
+The references contained in the OAEI tracks are not “complete” gold standards and some need further development.
+In this study, we only deal with equivalence mappings between classes and properties.
+Subsumption mappings are not considered.
+Finding widely used baseline data with ground truth matches for subsumptions is empirically difficult.
+Only a few OAEI tracks contain subsumption mappings, and they are not enough to robustly demonstrate the text preprocessing used in subsumption mappings.
+This work has not been included in the official OAEI campaign because of its intense focus on text preprocessing in syntactic OM.
+Artifacts
+The source code is available from Github at https://github.com/qzc438-research/ontology-nlp.
+The OAEI datasets are available from the OAEI Matching EvaLuation Toolkit (MELT) at https://dwslab.github.io/melt/track-repository (retrieved January 1, 2025).
+The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.

### v7

@@
 This approach utilises the strong background knowledge provided by LLMs to repair non-existent and counter-intuitive false mappings after the text preprocessing.
-It also overcomes the tendency towards unstable true mappings by injecting the classical text preprocessing pipeline via function calling.
+It also overcomes the tendency towards unstable true mappings by injecting the classic text preprocessing pipeline via function calling.
 The experimental results show that these two approaches can improve the matching correctness and the overall matching performance.
@@
 An OM process usually takes two ontologies as input, discovers mappings between entities, and produces a set of correspondences (Euzenat and Shvaiko, 2013).
-A classical OM system usually contains syntactic matching, lexical matching, and semantic matching.
+A classic OM system usually contains syntactic matching, lexical matching, and semantic matching.
 This multilayer architecture has been implemented in several successful OM systems, such as LogMap (Jiménez-Ruiz and Cuenca Grau, 2011; Jiménez-Ruiz et al., 2011), AgreementMakerLight (AML) (Faria et al., 2013, 2014), and FCA-Map (Zhao et al., 2018; Li et al., 2021).
-While syntactic matching captures “anchor mappings” and provides the foundation for the latter lexical and semantic matching, there are many strategies for extracting syntactic information from an ontology entity.
+Syntactic matching captures “anchor mappings” and provides the foundation for the latter lexical and semantic matching.
+There are many strategies for extracting syntactic information from an ontology entity.
 These include older approaches Bag-of-Words (e.g.
@@
 BERT (Devlin et al., 2019)).
-Despite the diversity of the models used, they all apply the text preprocessing pipeline to clean the text data before using the model.
-The classical text preprocessing pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L) (Manning et al., 2008).
+Despite the diversity of the strategies used, they all apply the text preprocessing pipeline to clean the text data before using the model.
+The classic text preprocessing pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L) (Manning et al., 2008).
 T is the process of breaking the text into the smallest units (i.e.
@@
 Entity Name Matching Output Preprocessing 1 ProgramCommittee Program Committee T Program_Committee 2 ProgramCommittee program committee T + N Program_committee 3 has_a_review review T + N + R hasReview 4 reviews review T + N + R + S/L isReviewing
-Table 1. Examples of using the classical text preprocessing pipeline in syntactic ontology matching: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
+Table 1. Examples of using the classic text preprocessing pipeline in syntactic ontology matching: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
 A number of OM systems use the text preprocessing pipeline for syntactic OM, but few studies question its role in optimising the matching performance.
@@
 While several steps of the text preprocessing pipeline are currently less effective, we propose two novel pipeline repair approaches.
-These two methods utilise the logic of ontological representations and the power of large language models (LLMs), showing promising results for repairing the less effective text preprocessing pipeline.
+These two methods utilise the logic of ontological representations and the power of large language models (LLMs), showing promising results for repairing weaknesses of the text preprocessing pipeline.
 Specifically, this paper makes the following contributions:
-  • We compare the performance of the text preprocessing pipeline on 8 Ontology Alignment Evaluation Initiative (OAEI) (OAEI Community, [n.
-d.]) datasets with 49 distinct alignments.
+  • We compare the performance of the text preprocessing pipeline on 8 datasets of the Ontology Alignment Evaluation Initiative (OAEI) (OAEI Community, [n.
+d.]) with 49 distinct alignments.
 While there is a significant improvement using Tokenisation (T) and Normalisation (N) in syntactic OM, Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are currently less effective.
@@
 The text in each document is preprocessed with tokenisation, stop words removal, and stemming.
-Recently, machine learning (ML) models have emerged for modern OM systems.
+Machine learning (ML) models have emerged for modern OM systems.
 While text preprocessing remains useful, its role is more focused on normalising the text that becomes the input to the model.
@@
 However, to the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have evaluated the effect of the entire text preprocessing pipeline on syntactic OM.
+Early explorations in (He et al., 2023; Norouzi et al., 2023) show the potential of using LLMs for OM, but matching by LLM prompting alone commonly results in poor performance.
+OLaLa (Hertling and Paulheim, 2023) proposes a prototype and states that the LLMs for OM include a number of decisions on model and parameter combinations, and there is no one-model-fits-all solution.
+Works on both Agent-OM (Qiang et al., 2023) and LLM4OM (Giglou et al., 2024) point out that LLMs have limited context length and inherited hallucinations.
+While LLM4OM highlights the importance of integrating the Retrieval-Augmented Generation (RAG) (Lewis et al., 2020), Agent-OM introduces an agent-based framework and enhances the compatibility of LLMs via function calling (OpenAI, 2023) for OM to improve matching performance.
+The authors in (Taboada et al., 2025) state that the use of LLMs for OM has a high computational overhead and apply a prioritised depth-first search (PDFS) strategy on the retrieve-then-prompt pipeline to optimise candidate selection.
+Recent works in (Amini et al., 2024; Zamazal, 2024; Sousa et al., 2025) also utilise LLMs and LLM embeddings for complex OM.
+However, these works focus on building a sophisticated framework for using LLMs for OM, and less attention has been paid to the subtask of using LLMs for syntactic matching.
+For syntactic OM, the challenges of using LLMs are twofold.
+On one hand, LLM hallucinations may produce incorrect “seed mappings”.
+Such mappings are toxic for the subsequent lexical and semantic matching.
+On the other hand, it is difficult to select a generic LLM prompting strategy from a number of LLM prompting templates and strategies.
 3. Preliminaries
+An ontology provides a formalised conceptualisation of domain knowledge (Gruber, 1993).
+The basic component of an ontology is the entity.
+Each entity is assigned a unique identifier (i.e.
+IRI) and represents a unique concept of the domain.
+Entities can be a concept (i.e.
+class), relation (i.e.
+property), or instance (i.e.
+individual) (Noy et al., 2001).
+Ideally, the entity names are not duplicated within a single ontology to minimise confusion and overlap.
+It is recommended that ontologies representing similar domain knowledge use similar terminologies for entity names (OBO Foundry, [n.
+d.]).
+Ontology matching (OM) aims to capture and align similar concepts that occur in different ontologies.
 Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), OM establishes mappings between pairs of entities drawn from each of the two ontologies.
@@
 $r$ is the relationship between two matched entities $e_{1}$ and $e_{2}$, and $c\in[0,1]$ is the confidence of the correspondence (or may be absent).
-The relationship $r$ in OM tasks can be equivalence ($\equiv$), subsumption ($\subseteq$), or other more complex relationships.
-In this study, we focus only on the equivalence relationship ($\equiv$) and evaluate the effect of text preprocessing on syntactic matching to produce the “anchor mappings” on which to base any subsequent lexical and semantic matching.
+The relationship $r$ in OM tasks can be equivalence ($\equiv$), subsumption ($\subseteq$), or other more complex relationships (Euzenat and Shvaiko, 2013).
+OM is a complex process that combines syntactic matching, lexical matching, and semantic matching.
+Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching.
+This study concentrates on the common practice of using the text preprocessing pipeline in syntactic OM.
+We focus only on the equivalence relationship ($\equiv$) and evaluate the effect of text preprocessing on syntactic matching to produce the “anchor mappings” on which to base any subsequent lexical and semantic matching.
 We address only equivalence because subsumption is typically dealt with in a later semantic and structural matching phase of OM.
-An alignment (A) is a set of candidate correspondences, while a reference (R) is a set of gold-standard correspondences verified by domain experts (i.e.
-the ground truth alignment).
 4. Text Preprocessing Pipeline Analysis
@@
 If $f(e_{1})=f(e_{2})$, we store the correspondence in the corresponding alignment file.
+While OM methods will often follow text preprocessing with a more sophisticated strategy to determine equivalence of entities, we define equivalence to mean identical text generated by preprocessing alone.
+This enables us to directly evaluate the preprocessing performance as a standalone contribution to OM.
 We compare the generated alignment (A) with the reference (R) to evaluate the performance of the text preprocessing pipeline on syntactic OM.
 Figure 1. Experiment setup to analyse the text preprocessing pipeline: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
+[…]
+Chart of experiment setup to analyse the text preprocessing pipeline.
   1. (1)
@@
 There are 49 distinct alignments evaluated in this study.
+Figure 2 shows the number of entities in each alignment.
     Table 2. Selected OAEI tracks and the number of alignments.
     Name Domain Alignments Anatomy Human and Mouse Anatomy 1 Biodiv Biodiversity and Ecology 9 CommonKG Common Knowledge Graphs 3 Conference Conference 24 Food Food Nutritional Composition 1 Largebio Biomedical 6 MSE Materials Science & Engineering 3 Phenotype Disease and Phenotype 2
+    Figure 2. Number of entities in each alignment.
+[…]
+    Chart of number of entities in each alignment.
+    Compound words are frequently used in ontology naming conventions.
+For example, compound words such as “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO (Suchanek et al., 2007) or the Snake case “art_gallery” used in Wikidata (Vrandečić and Krötzsch, 2014).
+Figure 3(b) shows the frequency distribution of compound words in the selected OAEI tracks.
+For comparisons between entities with compound words, one approach is to use their tokens, where “XY” is equivalent to “YX” because they share the same concepts X and Y.
+However, our comparison considers both concepts and their order, so that “XY” is not equivalent to “YX”.
+Although concurrent “XY” and “YX” are rare in the matching process, we assume that it could happen and so we take account of the order of concepts in compound word names to ensure a fair comparison in our experiments.
+    Figure 3. Frequency distribution of compound words. We exclude entities with more than 15 compound words because their proportion is less than 1%.
+[…]
+    Chart of frequency distribution of compound words.
   2. (2)
@@
     Figure 4. OM evaluation (Euzenat, 2007).
+[…]
+    Chart of OM evaluation (Euzenat, 2007).
     $Precision(Prec)=\frac{|A\cap R|}{|A|}$
@@
     $F_{1}\ Score=\frac{2}{Prec^{-1}+Rec^{-1}}$
-Compound words are frequently used in ontology naming conventions.
-For example, compound words such as “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO (Suchanek et al., 2007) or the Snake case “art_gallery” used in Wikidata (Vrandečić and Krötzsch, 2014).
-Figure 3 shows the proportion of compound words in the selected OAEI tracks.
-For comparisons between entities with compound words, one approach is to use their tokens, where “XY” is equivalent to “YX” because they share the same concepts X and Y.
-However, our comparison considers both concepts and their order, so that “XY” is not equivalent to “YX”.
-Although concurrent “XY” and “YX” are rare in the matching process, we assume that it could happen and so we take account of the order of concepts in compound word names to ensure a fair comparison in our experiments.
-Figure 3. Proportion of compound words. We exclude entities with more than 15 number of words because their proportion is less than 1%.
 4.2. Results
@@
 In the violin plots, three horizontal lines inside each violin plot show three quartiles: Q1, median, and Q3 (an artefact of the Seaborn (Waskom, 2021) package).
+[…]
+Comparison of text preprocessing pipeline.
   1. (1)
@@
     Figure 6. Stemming (S) vs. Lemmatisation (L).
+[…]
+    Stemming (S) vs. Lemmatisation (L).
   2. (2)
@@
     Figure 7. Comparison of different stemmers: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL).
+[…]
+    Comparison of different stemmers.
   3. (3)
@@
     Figure 8. Lemmatisation (L) without/with POS Tagging (LT).
+[…]
+    Lemmatisation (L) without/with POS Tagging (LT).
 4.3. Implications
@@
 Figure 9. Motivation of our context-based pipeline repair.
+[…]
+Motivation of our context-based pipeline repair.
 5.3. Ad Hoc Logic-Based Repair
@@
 One critical step in our approach is to assemble a set of reserved words that may cause FPs in text preprocessing, and to exclude these words from text preprocessing.
-The selection criteria are based on two widely agreed assumptions: (1) there are no duplicate entities within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies (OBO Foundry, [n.
-d.]).
+The selection criteria are based on two ontology features: (1) entity names (class or property) are not duplicated within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies (see Section 3).
 This approach is context-based because pairs of words may have the same or different meanings in different contexts.
-Based on these two assumptions, we propose a simple Algorithm 1 to assemble the reserved word set.
+Based on these two features, we propose a simple Algorithm 1 to assemble the reserved word set.
 For both $O_{s}$ and $O_{t}$, the algorithm first iterates on all pairs of entities $e_{i},e_{j}$ in the ontology, respectively.
@@
 Algorithm 1 Finding Reserved Word Set
-0:  Source Ontology $O_{s}$, Target Ontology $O_{t}$, Text Preprocessing Pipeline Method $f(x)$
-0:  $Reserved\_Word\_Set$
+0:  Source Ontology $O_{s}$, Target Ontology $O_{t}$, Text Preprocessing Pipeline Method $f(\cdot)$
+0:  Reserved_Word_Set $S$
   /* Phase 1: Find duplicates in $O_{s}$ and $O_{t}$/
-  /* Phase 1.1: Find duplicates in $O_{s}$/
+  /* Find duplicates in $O_{s}$/
   for $Entity\ e_{i},e_{j}\in O_{s}$ do
@@
         /* Pick up every different word between $e_{i}$ and $e_{j}$*/
-        $Reserved\_Word\_Set\leftarrow differ(e_{i},e_{j})$
+        $diff(e_{i},e_{j})=(e_{i}\setminus e_{j})\cup(e_{j}\setminus e_{i})$
+        $S=S+diff(e_{i},e_{j})$
      end if
   end for
-  /* Phase 1.2: Find duplicates in $O_{t}$/
+  /* Find duplicates in $O_{t}$/
   /* Same procedure applies …*/
-  /* Phase 2: Find immutable words in $Reserved\_Word\_Set$/
-  for $Word\ w\in Reserved\_Word\_Set$ do
+  /* Phase 2: Find immutable words in $S$/
+  for $Word\ w\in S$ do
      if $f(w)=w$ then
-        $Reserved\_Word\_Set\rightarrow w$
+        $S=S-w$
      end if
   end for
-  return  $Reserved\_Word\_Set$
+  return  Reserved_Word_Set $S$
 An example of generating and using a simple reserved word set is illustrated below.
@@
 Self-Reflection Write a short explanation.
-    Figure 9. Discovery rate of TPs using LLMs on the Largebio Track.
-    Figure 10. The discovery rate of FPs using LLMs on the Largebio Track.
   4. (4)
-    We compare LLMs with the classical text preprocessing pipeline in detecting TPs and FPs, which are two main contributors to precision and recall in matching performance.
+    We compare LLMs with the classic text preprocessing pipeline in detecting TPs and FPs, which are two main contributors to precision and recall in matching performance.
 The true mappings (TPs) are retrieved from the reference (R) where the mappings can be found by syntactic OM.
@@
 There could be slight differences in the results over multiple trials, but these differences are not significant with respect to the ranking and trend of different LLMs and prompt templates.
-Figure 9 compares LLMs with the classical text preprocessing pipeline in detecting TPs.
+Figure 10(l) compares LLMs with the classic text preprocessing pipeline in detecting TPs.
 LLMs and the text preprocessing pipeline are expected to answer “yes” for TPs.
@@
 On the other hand, using the classic text preprocessing pipeline via function calling could help LLMs produce a stable result with a high discovery rate (almost 100% from our observation) across different models and prompt templates.
-Figure 10 compares LLMs with the classical text preprocessing pipeline in detecting FPs.
+Figure 10. Discovery rate of TPs using LLMs on the Largebio Track.
+[…]
+Discovery rate of TPs using LLMs on the Largebio Track.
+Figure 11. The discovery rate of FPs using LLMs on the Largebio Track.
+[…]
+Discovery rate of FPs using LLMs on the Largebio Track.
+Figure 11(l) compares LLMs with the classic text preprocessing pipeline in detecting FPs.
 LLMs and the text preprocessing pipeline are expected to answer “no” for FPs.
@@
 While mapping repair is usually non-trivial and requires domain expertise, LLMs have shown potential for performing a high-quality post hoc correction without human intervention.
-For syntactic OM, we can see that LLMs are not good at detecting true mappings, but they are good at detecting false mappings.
-Therefore, we propose our generic post hoc LLM-based repair with (1) employing the classical text preprocessing pipeline via function calling (OpenAI, 2023) (a.k.a.
+For syntactic OM, we can see that LLMs are not good at detecting true mappings, but they are good at detecting false mappings. Therefore, we propose our generic post hoc LLM-based repair with
+  1. (1)
+    Employ the classic text preprocessing pipeline via function calling (a.k.a.
 tool calling) to detect TPs.
@@
 LLM prompting is not suitable for this task, even with complex prompting strategies.
-(2) Using LLM prompting to detect FPs.
+  2. (2)
+    Use LLM prompting to detect FPs.
 The prompt template can be simple like “Is A equivalent to B?”, as we cannot see significant improvement using complex prompting strategies (except Meta Llama models).
-5.5. Evaluation & Ablation Study
+5.5. Evaluation
 We apply our pipeline repair approaches on the Phase 2 text preprocessing method Lancaster Stemmer (SL), and evaluate their performance on the three large datasets in the Largebio Track: WHOLE-FMA-NCI, WHOLE-FMA-SNOMED, and WHOLE-SNOMED-NCI.
-Figure 11 compares the matching performance with/without using the pipeline approaches: ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of the above (CR).
+Figure 12(c) compares the matching performance with/without using the pipeline approaches: ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of the above (CR).
 We can see that ad hoc logic-based repair (AR) and post hoc LLM-based repair (PR) can significantly improve precision, with only a slight decrease in recall, but the overall F1 score also increases.
@@
 Evaluation of pipeline repair approaches: Ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of above (CR).
-The symbol $\varnothing$ represents the baseline without text preprocessing.
-Figure 12 compares post hoc LLM-based repair (PR) with two alternative repair approaches.
+The symbol $\varnothing$ represents the baseline without text preprocessing repair.
+[…]
+Evaluation of pipeline repair approaches.
+Figure 13 evaluates the scalability of ad hoc logic-based repair (AR).
+The x-axis indicates the number of entities in each alignment, and the y-axis shows the number of reserved words per 100 entities.
+For all the alignments we tested, the value of reserved words @100 is less than 10.
+With the number of entities increasing (i.e.
+the ontology size increases), the value of reserved words @100 does not increase but decreases.
+This feature of our approach is scalable in large-scale syntactic OM.
+Figure 13. Comparison of reserved words per 100 entities (@100) with the number of entities in each alignment.
+[…]
+Comparison of reserved words per 100 entities (@100) with the number of entities in each alignment.
+Figure 14(c) compares post hoc LLM-based repair (PR) with two alternative repair approaches.
 The results show that PR achieves almost the same precision with text preprocessing only repair (TOR) and LLM only repair (LOR), but its recall and F1 score are better than the other two alternatives.
@@
 Figure 14. Comparison of post hoc LLM-based repair (PR) with two alternative repair approaches: text preprocessing only repair (TOR) and LLM only repair (LOR).
+[…]
+Comparison of post hoc LLM-based repair (PR) with two alternative repair approaches.
 6. Conclusion
@@
 F1 score).
-OM aims to bridge the gap by capturing similar concepts that occur in different ontologies.
-OM is a complex process that combines syntactic matching, lexical matching, and semantic matching.
-Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching.
-This study concentrates on the common practice of using the text preprocessing pipeline in syntactic OM.
-While generic text preprocessing pipeline methods in syntactic OM are usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
+While the classic text preprocessing pipeline in syntactic OM is usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
   1. (1)
@@
   3. (3)
-    This study demonstrates that the nature of text preprocessing used in NLP applications can be context-based.
+    We showcase the limitations and opportunities for integrating the text preprocessing pipeline with LLMs in syntactic OM.
+We find that LLMs are good tools to find false mappings, but their outputs for finding true mappings are non-deterministic.
+The best practice is to integrate the classic text preprocessing pipelines with the LLMs.
+Using function calling to invoke the classic text preprocessing pipeline within LLMs, our approach could produce more stable and robust “seed mappings” than traditional text preprocessing methods and purely prompt-based approaches.
+  4. (4)
+    Our study also demonstrates that the nature of text preprocessing used in NLP applications can be context-based.
 Not all pipeline methods are effective, and some of them may even hamper the performance of downstream tasks.
@@
 The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.
+Appendix A Missing References in the LargeBio Track
+As mentioned in the Limitations Section, the OAEI tracks are undergoing continuous development and may miss some reference alignments.
+Here we list the missing reference alignments found by the text preprocessing pipeline in the folder “csvs”.
+These mappings have outstanding possibilities of being correct, and we consider these mappings to be TPs in our study.

### v8

@@
 Abstract.
-The classic text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many systems for syntactic ontology matching (OM).
+The classical text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many systems for syntactic ontology matching (OM).
 However, the lack of standardisation in text preprocessing creates diversity in mapping results.
@@
 This approach utilises the strong background knowledge provided by LLMs to repair non-existent and counter-intuitive false mappings after the text preprocessing.
-It also overcomes the tendency towards unstable true mappings by injecting the classic text preprocessing pipeline via function calling.
+It also overcomes the tendency towards unstable true mappings by injecting the classical text preprocessing pipeline via function calling.
 The experimental results show that these two approaches can improve the matching correctness and the overall matching performance.
@@
 An OM process usually takes two ontologies as input, discovers mappings between entities, and produces a set of correspondences (Euzenat and Shvaiko, 2013).
-A classic OM system usually contains syntactic matching, lexical matching, and semantic matching.
+A classical OM system usually contains syntactic matching, lexical matching, and semantic matching.
 This multilayer architecture has been implemented in several successful OM systems, such as LogMap (Jiménez-Ruiz and Cuenca Grau, 2011; Jiménez-Ruiz et al., 2011), AgreementMakerLight (AML) (Faria et al., 2013, 2014), and FCA-Map (Zhao et al., 2018; Li et al., 2021).
@@
 Despite the diversity of the strategies used, they all apply the text preprocessing pipeline to clean the text data before using the model.
-The classic text preprocessing pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L) (Manning et al., 2008).
+The classical text preprocessing pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L) (Manning et al., 2008).
 T is the process of breaking the text into the smallest units (i.e.
@@
 Entity Name Matching Output Preprocessing 1 ProgramCommittee Program Committee T Program_Committee 2 ProgramCommittee program committee T + N Program_committee 3 has_a_review review T + N + R hasReview 4 reviews review T + N + R + S/L isReviewing
-Table 1. Examples of using the classic text preprocessing pipeline in syntactic ontology matching: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
+Table 1. Examples of using the classical text preprocessing pipeline in syntactic ontology matching: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
 A number of OM systems use the text preprocessing pipeline for syntactic OM, but few studies question its role in optimising the matching performance.
@@
 The shape of the distribution is identical.
-    Figure 6. Stemming (S) vs. Lemmatisation (L).
+    Figure 6. Stemming (S) vs Lemmatisation (L).
 […]
-    Stemming (S) vs. Lemmatisation (L).
+    Stemming (S) vs Lemmatisation (L).
   2. (2)
@@
     2. (b)
-      Stemming (S) vs.
-Lemmatisation (L): While L tends to avoid generating FPs, it may also miss some implicit TPs.
+      Stemming (S) vs Lemmatisation (L): While L tends to avoid generating FPs, it may also miss some implicit TPs.
 On the other hand, S is more eager to find TPs, but eagerness can also lead to more FPs being found.
@@
     4. (d)
-      Lemmatisation (L) vs.
-Lemmatisation + POS Tagging (LT): LT is generally expected to have better results than L alone because tagging can help detect more precise root words.
+      Lemmatisation (L) vs Lemmatisation + POS Tagging (LT): LT is generally expected to have better results than L alone because tagging can help detect more precise root words.
 However, we do not observe such a performance improvement when using POS tagging in our study.
@@
 For a specific text preprocessing method $f(\cdot)$, if $f(e_{i})=f(e_{j})$, we retrieve the different words between $e_{i}$ and $e_{j}$ and store them in the reserved word set.
+We define $S_{i}=\{i_{1},..,p,..,i_{n}\}$ is the string set of $e_{i}$ and $S_{j}=\{j_{1},...,p,...j_{m}\}$ is the string set of $e_{j}$.
+Let $p$ be the longest common substrings in both $S_{i}$ and $S_{j}$, then $S_{e}\div S_{j}=\{i_{1},..,i_{n},j_{1},...,j_{m}\}$.
 If a word appears in the reserved set, the text preprocessing pipeline skips the operation for this word.
@@
 0:  Source Ontology $O_{s}$, Target Ontology $O_{t}$, Text Preprocessing Pipeline Method $f(\cdot)$
-0:  Reserved_Word_Set $S$
+0:  Reserved_Word_Set $S_{r}$
   /* Phase 1: Find duplicates in $O_{s}$ and $O_{t}$/
@@
      if $f(e_{i})=f(e_{j})$ then
-        /* Pick up every different word between $e_{i}$ and $e_{j}$*/
-        $diff(e_{i},e_{j})=(e_{i}\setminus e_{j})\cup(e_{j}\setminus e_{i})$
-        $S=S+diff(e_{i},e_{j})$
+        $S_{r}=S_{r}\cup(S_{e}\div S_{j})$
      end if
@@
      if $f(w)=w$ then
-        $S=S-w$
+        $S_{r}=S_{r}\setminus\{w\}$
      end if
   end for
-  return  Reserved_Word_Set $S$
+  return  Reserved_Word_Set $S_{r}$
 An example of generating and using a simple reserved word set is illustrated below.
@@
   4. (4)
-    We compare LLMs with the classic text preprocessing pipeline in detecting TPs and FPs, which are two main contributors to precision and recall in matching performance.
+    We compare LLMs with the classical text preprocessing pipeline in detecting TPs and FPs, which are two main contributors to precision and recall in matching performance.
 The true mappings (TPs) are retrieved from the reference (R) where the mappings can be found by syntactic OM.
-The false mappings (FPs) are retrieved from the alignment (A) generated by the classic text preprocessing pipeline but exclude TPs above.
+The false mappings (FPs) are retrieved from the alignment (A) generated by the classical text preprocessing pipeline but exclude TPs above.
   5. (5)
@@
 There could be slight differences in the results over multiple trials, but these differences are not significant with respect to the ranking and trend of different LLMs and prompt templates.
-Figure 10(l) compares LLMs with the classic text preprocessing pipeline in detecting TPs.
+Figure 10(l) compares LLMs with the classical text preprocessing pipeline in detecting TPs.
 LLMs and the text preprocessing pipeline are expected to answer “yes” for TPs.
@@
 Finding a good example in real-world applications is not easy because ontology entities can have different formatting by nature.
-On the other hand, using the classic text preprocessing pipeline via function calling could help LLMs produce a stable result with a high discovery rate (almost 100% from our observation) across different models and prompt templates.
+On the other hand, using the classical text preprocessing pipeline via function calling could help LLMs produce a stable result with a high discovery rate (almost 100% from our observation) across different models and prompt templates.
 Figure 10. Discovery rate of TPs using LLMs on the Largebio Track.
@@
 Discovery rate of FPs using LLMs on the Largebio Track.
-Figure 11(l) compares LLMs with the classic text preprocessing pipeline in detecting FPs.
+Figure 11(l) compares LLMs with the classical text preprocessing pipeline in detecting FPs.
 LLMs and the text preprocessing pipeline are expected to answer “no” for FPs.
 The discovery rate is defined as the number of “no” answers compared to the total number of FPs.
-Specifically, these FPs evaluated are non-trivial mappings that cannot be detected by the classic text preprocessing method.
+Specifically, these FPs evaluated are non-trivial mappings that cannot be detected by the classical text preprocessing method.
 Here we use LLMs to detect FPs generated by the Lancaster Stemmer (SL).
 SL has shown a lower precision compared to other text preprocessing methods, which is caused by the method producing a significant number of FPs.
-LLMs have shown a strong capability to discover the FPs produced by the classic text preprocessing pipeline.
+LLMs have shown a strong capability to discover the FPs produced by the classical text preprocessing pipeline.
 For different models, API-accessed commercial LLMs are better than open-source LLMs and reach a higher discovery rate of more than 80%.
@@
   1. (1)
-    Employ the classic text preprocessing pipeline via function calling (a.k.a.
+    Employ the classical text preprocessing pipeline via function calling (a.k.a.
 tool calling) to detect TPs.
@@
 The x-axis indicates the number of entities in each alignment, and the y-axis shows the number of reserved words per 100 entities.
-For all the alignments we tested, the value of reserved words @100 is less than 10.
+For all the alignments we tested, the value of reserved words per 100 entities is less than 10.
 With the number of entities increasing (i.e.
-the ontology size increases), the value of reserved words @100 does not increase but decreases.
-This feature of our approach is scalable in large-scale syntactic OM.
-Figure 13. Comparison of reserved words per 100 entities (@100) with the number of entities in each alignment.
+the ontology size increases), the number of reserved words per 100 entities decreases.
+This may be because reserved words such as“has”, “is”, and “a” are commonly used in many ontologies and in multiple distinct entities within those ontologies.
+Conventionally there are a limited number of such words.
+Therefore the reserved word feature of our approach is scalable in large-scale OM.
+Figure 13. Reserved words per 100 entities vs the number of entities in each alignment.
 […]
@@
 Comparison of post hoc LLM-based repair (PR) with two alternative repair approaches.
+6. Limitations
+  1. (1)
+    The references contained in the OAEI tracks are not perfect gold standards and some need further development. Therefore at present 100% accuracy is not necessarily desirable.
+  2. (2)
+    In this study we deal only with equivalence mappings between classes and properties.
+Subsumption mappings are not considered.
+Finding widely used baseline data with ground truth matches for subsumptions is empirically difficult.
+Only a few OAEI tracks contain subsumption mappings, and they are not enough to robustly demonstrate the text preprocessing used in subsumption mappings.
+  3. (3)
+    This work has not been included in the official OAEI campaign because of its intense focus on text preprocessing in syntactic OM.
 7. Conclusion
@@
 F1 score).
-While the classic text preprocessing pipeline in syntactic OM is usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
+While the classical text preprocessing pipeline in syntactic OM is usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
   1. (1)
@@
 We find that LLMs are good tools to find false mappings, but their outputs for finding true mappings are non-deterministic.
-The best practice is to integrate the classic text preprocessing pipelines with the LLMs.
-Using function calling to invoke the classic text preprocessing pipeline within LLMs, our approach could produce more stable and robust “seed mappings” than traditional text preprocessing methods and purely prompt-based approaches.
+The best practice is to integrate the classical text preprocessing pipelines with the LLMs.
+Using function calling to invoke the classical text preprocessing pipeline within LLMs, our approach could produce more stable and robust “seed mappings” than traditional text preprocessing methods and purely prompt-based approaches.
   4. (4)
@@
 The authors thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.
-Limitations
-The references contained in the OAEI tracks are not “complete” gold standards and some need further development.
-In this study, we only deal with equivalence mappings between classes and properties.
-Subsumption mappings are not considered.
-Finding widely used baseline data with ground truth matches for subsumptions is empirically difficult.
-Only a few OAEI tracks contain subsumption mappings, and they are not enough to robustly demonstrate the text preprocessing used in subsumption mappings.
-This work has not been included in the official OAEI campaign because of its intense focus on text preprocessing in syntactic OM.
 Artifacts
@@
 Appendix A Missing References in the LargeBio Track
-As mentioned in the Limitations Section, the OAEI tracks are undergoing continuous development and may miss some reference alignments.
-Here we list the missing reference alignments found by the text preprocessing pipeline in the folder “csvs”.
-These mappings have outstanding possibilities of being correct, and we consider these mappings to be TPs in our study.
+As mentioned in the Limitations Section, the OAEI tracks are undergoing continuous development and may be missing some reference alignments.
+Here we list the missing reference alignments found by our text preprocessing pipeline in the folder “csvs”.
+These mappings would appear to be correct and we treat these mappings as TPs in our study.

### v10

@@
-How Does A Text Preprocessing Pipeline Affect Syntactic Ontology Matching?
-Abstract.
-The classical text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many systems for syntactic ontology matching (OM).
-However, the lack of standardisation in text preprocessing creates diversity in mapping results.
-In this paper, we investigate the effect of the text preprocessing pipeline on syntactic OM in 8 Ontology Alignment Evaluation Initiative (OAEI) tracks with 49 distinct alignments.
-We find that Phase 1 text preprocessing (Tokenisation and Normalisation) is more effective than Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation).
-We propose two novel approaches to repair unwanted false mappings caused by Phase 2 text preprocessing.
-One is an ad hoc logic-based repair approach that employs an ontology-specific check to find common words that cause false mappings.
-These words are stored in a reserved word set and applied before the text preprocessing.
-By leveraging the power of large language models (LLMs), we also propose a post hoc LLM-based repair approach.
-This approach utilises the strong background knowledge provided by LLMs to repair non-existent and counter-intuitive false mappings after the text preprocessing.
-It also overcomes the tendency towards unstable true mappings by injecting the classical text preprocessing pipeline via function calling.
-The experimental results show that these two approaches can improve the matching correctness and the overall matching performance.
-Ontology Matching, Text Preprocessing
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 29th International Conference on Extending Database Technology; March 24–27, 2026; Tampere, Finlandisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Ontology engineering
-1. Introduction
-Ontology matching (OM) is essential for enabling interoperability between heterogeneous ontologies.
-An OM process usually takes two ontologies as input, discovers mappings between entities, and produces a set of correspondences (Euzenat and Shvaiko, 2013).
-A classical OM system usually contains syntactic matching, lexical matching, and semantic matching.
-This multilayer architecture has been implemented in several successful OM systems, such as LogMap (Jiménez-Ruiz and Cuenca Grau, 2011; Jiménez-Ruiz et al., 2011), AgreementMakerLight (AML) (Faria et al., 2013, 2014), and FCA-Map (Zhao et al., 2018; Li et al., 2021).
-Syntactic matching captures “anchor mappings” and provides the foundation for the latter lexical and semantic matching.
-There are many strategies for extracting syntactic information from an ontology entity.
-These include older approaches Bag-of-Words (e.g.
-TF-IDF (Sammut and Webb, 2010)), popular word embedding models (e.g.
-Word2Vec (Mikolov et al., 2013)), and state-of-the-art language models (e.g.
-BERT (Devlin et al., 2019)).
-Despite the diversity of the strategies used, they all apply the text preprocessing pipeline to clean the text data before using the model.
-The classical text preprocessing pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L) (Manning et al., 2008).
+How Does A Text Preprocessing Pipeline Affect Ontology Matching?
+Abstract
+The classical text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many systems for ontology matching (OM).
+However, the lack of standardisation in text preprocessing creates diversity in the mapping results.
+In this paper, we investigate the effect of the text preprocessing pipeline on 8 Ontology Alignment Evaluation Initiative (OAEI) tracks with 49 distinct alignments.
+We find that Tokenisation and Normalisation (categorised as Phase 1 text preprocessing) are more effective than Stop Words Removal and Stemming/Lemmatisation (categorised as Phase 2 text preprocessing).
+We propose two novel approaches to repair unwanted false mappings that occur in Phase 2 text preprocessing.
+One is an ad hoc logic-based repair approach used before text preprocessing, employing an ontology-specific check to find common words that cause false mappings.
+The other repair approach is the post hoc large language model (LLM)-based approach, used after text preprocessing, which utilises the strong background knowledge provided by LLMs to repair non-existent and counter-intuitive false mappings.
+The experimental results indicate that these two approaches can significantly improve the matching correctness and the overall matching performance.
+I Introduction
+Ontology matching (OM) is crucial for enabling interoperability between heterogeneous ontologies, thereby facilitating the integration of disparate knowledge graphs (KGs) [1].
+An OM process usually takes two ontologies as input, discovers mappings between entities, and produces a set of correspondences [2].
+The classical text preprocessing pipeline is commonly used in OM.
+The pipeline consists of a set of steps to segment, reconstruct, analyse, and process text, namely Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L) [3].
 T is the process of breaking the text into the smallest units (i.e.
 tokens).
-N is the process of transforming these different tokens into a single canonical form.
+N transforms these different tokens into a single canonical form.
 R is the process of removing filler words that usually carry little meaning and can be omitted in most cases.
 S/L is used to deal with the grammatical variation of words, applying rules to find the simplest common form of the word.
-Table 1 shows examples of using the text preprocessing pipeline to find mappings in syntactic OM.
-In the examples, we use whitespace to split the tokens, normalise all the words into lowercase, remove the common stop words “has”, “a” and “is”, and apply stemming/lemmatisation to find the simplest word form “review” from “reviews” and “reviewing”.
-No.
-Entity Name Matching Output Preprocessing 1 ProgramCommittee Program Committee T Program_Committee 2 ProgramCommittee program committee T + N Program_committee 3 has_a_review review T + N + R hasReview 4 reviews review T + N + R + S/L isReviewing
-Table 1. Examples of using the classical text preprocessing pipeline in syntactic ontology matching: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
-A number of OM systems use the text preprocessing pipeline for syntactic OM, but few studies question its role in optimising the matching performance.
+Fig.
+1 shows that two different ontology entities “reviews” and “isReviewing” can be matched via the text preprocessing pipeline as follows.
+After tokenisation, our tokens comprise a sequence of word units which we present as separated by white spaces within a token (“reviews” cf “is Reviewing’).
+We then normalise all words to lowercase (“reviews” cf “is reviewing”), remove the common stop words “is” (“reviews” cf “reviewing”), and finally apply stemming/lemmatisation to find the simplest word form “review” from both “reviews” and “reviewing” (“review” cf “review”).
+Figure 1: An Example of using the text preprocessing pipeline in OM.
+A number of OM systems utilise the text preprocessing pipeline, but few studies have justified the choice of a specific text preprocessing method and systematically studied its role in optimising the matching performance.
 Our study fills this gap.
-We first conduct a comprehensive experimental analysis of the text preprocessing pipeline used in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline.
+We first conduct a comprehensive experimental analysis of the text preprocessing pipeline used in OM across a wide range of domains, aiming to explain the impact on OM of each step in the pipeline.
 In each step, text preprocessing methods are evaluated for correctness and completeness.
-While several steps of the text preprocessing pipeline are currently less effective, we propose two novel pipeline repair approaches.
+We find that some steps of the text preprocessing pipeline are currently less effective.
+Based on this insight, we propose two novel pipeline repair approaches.
 These two methods utilise the logic of ontological representations and the power of large language models (LLMs), showing promising results for repairing weaknesses of the text preprocessing pipeline.
-Specifically, this paper makes the following contributions:
-  • We compare the performance of the text preprocessing pipeline on 8 datasets of the Ontology Alignment Evaluation Initiative (OAEI) (OAEI Community, [n.
-d.]) with 49 distinct alignments.
-While there is a significant improvement using Tokenisation (T) and Normalisation (N) in syntactic OM, Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are currently less effective.
-We find common inappropriate stop words removal, over-stemming, and over-lemmatisation.
-  • We propose two approaches to repair the less effective text preprocessing pipeline in syntactic OM: (1) ad hoc logic-based repair and (2) post hoc LLM-based repair.
-Experimental results show that these two methods can improve the matching correctness and the overall matching performance of syntactic OM.
-The remainder of the paper is organised as follows.
-Section 2 and 3 review the related work and the preliminaries of OM and syntactic matching.
-Section 4 analyses the text preprocessing pipeline used in syntactic OM.
-Section 5 proposes text preprocessing pipeline repair approaches and experimentally validates their performance.
-Section 7 highlights the contributions of this work and concludes the paper.
-2. Related Work
-Syntactic OM considers only the meaning of an entity’s name or label, ignoring its lexical and structural context in an ontology.
-Correct syntactic matches are often implicit and usually require extra human observation and domain knowledge (Euzenat and Shvaiko, 2013).
-Text preprocessing is introduced to automate this process.
-The use of text preprocessing in syntactic OM can be traced back to the early stages of OM systems, which were initially developed and widely used as a basic component to generate linguistic-based mappings.
-AgreementMaker (Cruz et al., 2009) has a normaliser to unify the textual information of the entities.
-SAMBO (Lambrix and Tan, 2006) uses the Porter Stemmer for each word to improve the similarity measure for terms with different prefixes and suffixes.
-In RiMOM (Li et al., 2009), the context information of each entity is viewed as a document.
-The text in each document is preprocessed with tokenisation, stop words removal, and stemming.
-Machine learning (ML) models have emerged for modern OM systems.
-While text preprocessing remains useful, its role is more focused on normalising the text that becomes the input to the model.
-For example, BERTMap (He et al., 2022) uses BERT’s inherent WordPiece tokeniser to build the subword of each entity.
-DeepOnto (He et al., 2024) extends normalisation to axioms using EL embedding models (Kulmanov et al., 2019).
-The ML extension of LogMap (Chen et al., 2021) reuses the seed mappings of traditional systems, where each entity is split into its component words and the mapping is based on pairwise component similarities.
-However, to the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have evaluated the effect of the entire text preprocessing pipeline on syntactic OM.
-Early explorations in (He et al., 2023; Norouzi et al., 2023) show the potential of using LLMs for OM, but matching by LLM prompting alone commonly results in poor performance.
-OLaLa (Hertling and Paulheim, 2023) proposes a prototype and states that the LLMs for OM include a number of decisions on model and parameter combinations, and there is no one-model-fits-all solution.
-Works on both Agent-OM (Qiang et al., 2023) and LLM4OM (Giglou et al., 2024) point out that LLMs have limited context length and inherited hallucinations.
-While LLM4OM highlights the importance of integrating the Retrieval-Augmented Generation (RAG) (Lewis et al., 2020), Agent-OM introduces an agent-based framework and enhances the compatibility of LLMs via function calling (OpenAI, 2023) for OM to improve matching performance.
-The authors in (Taboada et al., 2025) state that the use of LLMs for OM has a high computational overhead and apply a prioritised depth-first search (PDFS) strategy on the retrieve-then-prompt pipeline to optimise candidate selection.
-Recent works in (Amini et al., 2024; Zamazal, 2024; Sousa et al., 2025) also utilise LLMs and LLM embeddings for complex OM.
-However, these works focus on building a sophisticated framework for using LLMs for OM, and less attention has been paid to the subtask of using LLMs for syntactic matching.
-For syntactic OM, the challenges of using LLMs are twofold.
-On one hand, LLM hallucinations may produce incorrect “seed mappings”.
-Such mappings are toxic for the subsequent lexical and semantic matching.
-On the other hand, it is difficult to select a generic LLM prompting strategy from a number of LLM prompting templates and strategies.
-3. Preliminaries
-An ontology provides a formalised conceptualisation of domain knowledge (Gruber, 1993).
+Specifically, this paper makes the following contributions.
+  • We comprehensively evaluate the performance of each step in the text preprocessing pipeline on 8 datasets of the Ontology Alignment Evaluation Initiative (OAEI) [4] with 49 distinct alignments.
+While there is a significant improvement using Tokenisation (T) and Normalisation (N), Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are currently less effective.
+We find the issue is caused by inappropriate stop words removal, over-stemming, and over-lemmatisation.
+We also compare the performance of (1) Stemming and Lemmatisation, (2) different stemming methods including Porter, Snowball, and Lancaster, and (3) Lemmatisation with and without Part-of-Speech (POS) Tagging.
+  • We propose two approaches to repair the less effective text preprocessing pipeline: (1) ad hoc logic-based repair and (2) post hoc LLM-based repair.
+To reduce the number of false mappings, the logic-based repair approach applies an ad hoc check based on ontology-specific settings before text preprocessing, while the LLM-based repair utilises LLMs as a post hoc validator after text preprocessing.
+We examine the performance of OM across different LLMs and prompt templates.
+We find the optimal pipeline to combine the traditional text processing pipeline with modern LLMs for OM.
+The experimental results indicate that these two repair methods can significantly improve the matching correctness and the overall matching performance.
+The paper is organised as follows.
+Sections II and III review the preliminaries and related work of OM and the text preprocessing pipeline.
+Section IV analyses the text preprocessing pipeline used in OM.
+Section V proposes text preprocessing pipeline repair approaches and experimentally validates their performance.
+Section VI discusses the limitations of this work.
+Section VII concludes the paper.
+II Preliminaries
+An ontology provides a formalised conceptualisation of domain knowledge [5].
 The basic component of an ontology is the entity.
@@
 property), or instance (i.e.
-individual) (Noy et al., 2001).
+individual) [6].
 Ideally, the entity names are not duplicated within a single ontology to minimise confusion and overlap.
-It is recommended that ontologies representing similar domain knowledge use similar terminologies for entity names (OBO Foundry, [n.
-d.]).
+It is recommended that ontologies representing similar domain knowledge use similar terminologies for entity names [7].
 Ontology matching (OM) aims to capture and align similar concepts that occur in different ontologies.
@@
 $r$ is the relationship between two matched entities $e_{1}$ and $e_{2}$, and $c\in[0,1]$ is the confidence of the correspondence (or may be absent).
-The relationship $r$ in OM tasks can be equivalence ($\equiv$), subsumption ($\subseteq$), or other more complex relationships (Euzenat and Shvaiko, 2013).
-OM is a complex process that combines syntactic matching, lexical matching, and semantic matching.
-Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching.
-This study concentrates on the common practice of using the text preprocessing pipeline in syntactic OM.
-We focus only on the equivalence relationship ($\equiv$) and evaluate the effect of text preprocessing on syntactic matching to produce the “anchor mappings” on which to base any subsequent lexical and semantic matching.
-We address only equivalence because subsumption is typically dealt with in a later semantic and structural matching phase of OM.
-4. Text Preprocessing Pipeline Analysis
-In this section, we discuss the role of the text preprocessing pipeline in syntactic OM and report our experiments that study the effect of different text preprocessing options on the OAEI datasets.
-4.1. Experiment Setup
-Figure 1 shows the experiment setup to analyse the text preprocessing pipeline.
+The relationship $r$ in OM tasks can be equivalence ($\equiv$), subsumption ($\subseteq$), or other more complex relationships [2].
+An alignment (A) is a set of candidate correspondences, while a reference (R) is a set of gold-standard correspondences verified by domain experts (i.e.
+the ground truth alignment).
+A classical OM system usually contains syntactic matching, lexical matching, and semantic matching.
+This multilayer architecture has been implemented in several successful OM systems, including LogMap [8, 9], AgreementMakerLight (AML) [10, 11], FCA-Map [12, 13], and Agent-OM [14].
+Syntactic matching is often the first step of classical OM.
+Text preprocessing is commonly used for syntactic matching.
+It has proven to correctly match many ontology entities because entities with similar string combinations are likely to be correct, and it is computationally efficient.
+III Related Work
+III-A Text preprocessing in OM
+The use of text preprocessing in OM can be traced back to the early stages of OM systems, which were initially developed and widely used as a basic component to generate linguistic-based mappings.
+AgreementMaker [15] has a normaliser to unify the textual information of the entities.
+SAMBO [16] uses the Porter Stemmer for each word to improve the similarity measure for terms with different prefixes and suffixes.
+In RiMOM [17], the context information of each entity is viewed as a document.
+The text in each document is preprocessed with tokenisation, stop words removal, and stemming.
+Recently, machine learning (ML) models have emerged for modern OM systems.
+While text preprocessing remains useful, its role is more focused on normalising the text that becomes the input to the model.
+For example, BERTMap [18] uses BERT’s inherent WordPiece tokeniser to build the subword of each entity.
+DeepOnto [19] extends normalisation to axioms using EL embedding models [20].
+The ML extension of LogMap [21] reuses the seed mappings of traditional systems, where each entity is split into its component words and the mapping is based on pairwise component similarities.
+To the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have evaluated the effect of the entire text preprocessing pipeline.
+III-B LLMs in OM
+Advances in LLM research raises the question of whether text preprocessing is obsolete for OM.
+Early explorations in [22, 23] show the potential of using LLMs for OM, but matching by LLM prompting alone commonly results in poor performance.
+OLaLa [24] proposes a prototype and states that the LLMs for OM include a number of decisions on model and parameter combinations, and there is no one-model-fits-all solution.
+Works on both Agent-OM [14] and LLM4OM [25] point out that LLMs have limited context length and inherited hallucinations.
+While LLM4OM highlights the importance of integrating the Retrieval-Augmented Generation (RAG) [26], Agent-OM introduces an agent-based framework and enhances the compatibility of LLMs via function calling [27] for OM to improve matching performance.
+The authors in [28] state that the use of LLMs for OM has a high computational overhead and apply a prioritised depth-first search (PDFS) strategy on the retrieve-then-prompt pipeline to optimise candidate selection.
+Recent works in [29, 30, 31] also utilise LLMs and LLM embeddings for complex OM.
+However, these works focus on building a sophisticated framework for using LLMs for OM, and less attention has been paid to the subtask of using LLMs for text preprocessing as used in OM.
+The challenges are twofold.
+On one hand, LLM hallucinations may produce incorrect “seed mappings”.
+Such mappings are toxic for the subsequent lexical and semantic matching.
+On the other hand, it is difficult to select a generic LLM prompting strategy from a number of LLM prompting templates and strategies.
+With strong background knowledge, LLMs can be helpful in detecting false mappings after text preprocessing.
+We discuss LLM-based repair in Section V-D.
+IV Text Preprocessing Pipeline Analysis
+IV-A Experiment Setup
+Fig.
+2 shows the experiment setup to analyse the text preprocessing pipeline.
 Alignment (A) is generated via the text preprocessing pipeline.
@@
 If $f(e_{1})=f(e_{2})$, we store the correspondence in the corresponding alignment file.
-While OM methods will often follow text preprocessing with a more sophisticated strategy to determine equivalence of entities, we define equivalence to mean identical text generated by preprocessing alone.
+While OM systems often follow text preprocessing with a more sophisticated strategy to determine the equivalence of entities, we define equivalence to mean identical text generated by preprocessing alone.
 This enables us to directly evaluate the preprocessing performance as a standalone contribution to OM.
-We compare the generated alignment (A) with the reference (R) to evaluate the performance of the text preprocessing pipeline on syntactic OM.
-Figure 1. Experiment setup to analyse the text preprocessing pipeline: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
-[…]
-Chart of experiment setup to analyse the text preprocessing pipeline.
-  1. (1)
-    Selected OAEI tracks: We select 8 OAEI tracks stored in the Ontology Matching Evaluation Toolkit (MELT) (Hertling et al., 2019) public repository (retrieved January 1, 2025).
-Table 2 shows the details of the selected tracks.
+We compare the generated alignment (A) with the reference (R) to evaluate the performance of the text preprocessing pipeline on OM.
+Figure 2: Experiment setup to analyse the effect of text preprocessing pipeline in OM: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
+We select 8 OAEI tracks stored in the Ontology Matching Evaluation Toolkit (MELT) [32] public repository (retrieved January 1, 2025).
+Table I shows the details of the selected tracks.
 Each track may contain more than one alignment corresponding to different pairs of ontologies.
-For example, the Largebio Track has multiple reference files, pairing FMA (Rosse and Mejino, 2003) and NCI (Golbeck et al., 2003), FMA and SNOMED (Donnelly, 2006), and SNOMED and NCI, respectively.
+For example, the Largebio Track has multiple reference files, pairing FMA [33] and NCI [34], FMA and SNOMED [35], and SNOMED and NCI, respectively.
 The number of alignments for each track is given in the table.
 There are 49 distinct alignments evaluated in this study.
-Figure 2 shows the number of entities in each alignment.
-    Table 2. Selected OAEI tracks and the number of alignments.
-    Name Domain Alignments Anatomy Human and Mouse Anatomy 1 Biodiv Biodiversity and Ecology 9 CommonKG Common Knowledge Graphs 3 Conference Conference 24 Food Food Nutritional Composition 1 Largebio Biomedical 6 MSE Materials Science & Engineering 3 Phenotype Disease and Phenotype 2
-    Figure 2. Number of entities in each alignment.
-[…]
-    Chart of number of entities in each alignment.
+Fig.
+3 shows the number of entities in each alignment across different tracks.
+TABLE I: Selected OAEI tracks and the number of alignments.
+Name Domain # Alignments Anatomy Human and Mouse Anatomy 1 Biodiv Biodiversity and Ecology 9 CommonKG Common Knowledge Graphs 3 Conference Conference 24 Food Food Nutritional Composition 1 Largebio Biomedical 6 MSE Materials Science & Engineering 3 Phenotype Disease and Phenotype 2
+Figure 3: Number of entities in each alignment across different tracks.
 Compound words are frequently used in ontology naming conventions.
-For example, compound words such as “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO (Suchanek et al., 2007) or the Snake case “art_gallery” used in Wikidata (Vrandečić and Krötzsch, 2014).
-Figure 3(b) shows the frequency distribution of compound words in the selected OAEI tracks.
+For example, compound words such as “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO [36] or the Snake case “art_gallery” used in Wikidata [37].
+Fig.
+4 shows the frequency distribution of compound words in the selected OAEI tracks.
 For comparisons between entities with compound words, one approach is to use their tokens, where “XY” is equivalent to “YX” because they share the same concepts X and Y.
@@
 Although concurrent “XY” and “YX” are rare in the matching process, we assume that it could happen and so we take account of the order of concepts in compound word names to ensure a fair comparison in our experiments.
-    Figure 3. Frequency distribution of compound words. We exclude entities with more than 15 compound words because their proportion is less than 1%.
-[…]
-    Chart of frequency distribution of compound words.
-  2. (2)
-    Selected pipeline methods: There are a variety of methods that can be used in a general text preprocessing pipeline, but not all of them are applicable to syntactic OM.
-We select the following methods in each text preprocessing step: (a) Tokenisation (T): We use full word tokenisation to separate the compound words.
-Experiments on several subword tokenisation methods (e.g.
-Byte Pair Encoding (Gage, 1994; Sennrich et al., 2016), WordPiece (Schuster and Nakajima, 2012; Devlin et al., 2019), and Unigram (Kudo, 2018; Kudo and Richardson, 2018)) show the results after de-tokenisation (i.e.
-converting the tokens to a sentence) are generally the same as full word tokenisation.
+(a) Classes
+(b) Properties
+Figure 4: Frequency distribution of compound words. We exclude entities with more than 15 compound words because their proportion is less than 1%.
+There are a variety of methods that can be employed in a general text preprocessing pipeline, but not all of them are applicable to syntactic OM.
+We select the following methods in each text preprocessing step: (a) Tokenisation (T): We use the conventional full word tokenisation method to separate the compound words in an entity name.
+Experiments on several alternative subword tokenisation methods (e.g.
+Byte Pair Encoding [38, 39], WordPiece [40, 41], and Unigram [42, 43]) show that results after concatenating the processed name parts back together are generally the same as for full word tokenisation.
 We do not use sentence tokenisation (a.k.a.
@@
 removal of special characters and numbers).
-(c) Stop Words Removal (R): includes the most common English stop words defined in the Natural Language Toolkit (NLTK) (Bird et al., 2008).
+(c) Stop Words Removal (R): includes the most common English stop words defined in the Natural Language Toolkit (NLTK) [44].
 (d) Stemming/Lemmatisation (S/L): Stemming methods include Porter, Snowball, and Lancaster.
-Lemmatisation uses the NLTK Lemmatiser based on WordNet (Miller, 1995), and the word categorisation uses POS Tagging.
-  3. (3)
-    Selected measures: There are four primitive measures in information retrieval: true positive (TP), false positive (FP), false negative (FN), and true negative (TN).
+Lemmatisation utilises the NLTK Lemmatiser based on WordNet [45], and word categorisation employs POS Tagging.
+There are four primitive measures in information retrieval: true positive (TP), false positive (FP), false negative (FN), and true negative (TN).
 In the context of OM, the evaluation compares an alignment (A) returned by the OM system with a gold standard reference (R) validated by domain experts.
-Figure 4 illustrates that the four primitive measures in OM can be interpreted as $TP=A\cap R$, $FP=A-R$, $FN=R-A$, and $TN=(C\times C^{\prime})-(A\cup R)$, where $C\times C^{\prime}$ refers to all possible correspondences.
+Fig.
+5 illustrates that the four primitive measures in OM can be interpreted as $TP=A\cap R$, FP=ARFP=A-R, FN=RAFN=R-A, and $TN=(C\times C^{\prime})-(A\cup R)$, where $C\times C^{\prime}$ refers to all possible correspondences.
 Accuracy (Acc), Specificity (Spec), Precision (Prec), Recall (Rec), and $F_{\beta}$ Score ($F_{\beta}$) are the most common evaluation measures based on TP, FP, FN, and TN.
@@
 They are defined as follows.
-    Figure 4. OM evaluation (Euzenat, 2007).
-[…]
-    Chart of OM evaluation (Euzenat, 2007).
+Figure 5: OM evaluation [46].
 $Precision(Prec)=\frac{|A\cap R|}{|A|}$
@@
 $F_{1}\ Score=\frac{2}{Prec^{-1}+Rec^{-1}}$
-4.2. Results
-Figure 5 compares each step of the text preprocessing pipeline.
+IV-B Results
+Fig.
+6 compares each step of the text preprocessing pipeline.
 These steps are always applied sequentially in the pipeline.
 Beginning with $\varnothing$ (no text preprocessing), we see the performance of subsequent Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
-For T and N, most of the data points above the equivalence line in precision, recall, and F1 score indicate that they can help syntactic OM.
+For T and N, most of the data points above the equivalence line in precision, recall, and F1 score indicate that they can help OM.
 For R and S/L, some data points are above the equivalence line in recall, but most data points are below the equivalence line in precision and F1 score indicating that they do not play a positive role.
@@
 (c) F1 Score: The median increases with T and N but then decreases with R and S/L.
-The shape of the distribution does not change after N.
-Figure 5.
-Comparison of text preprocessing pipeline: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
-In the violin plots, three horizontal lines inside each violin plot show three quartiles: Q1, median, and Q3 (an artefact of the Seaborn (Waskom, 2021) package).
-[…]
-Comparison of text preprocessing pipeline.
-  1. (1)
-    Figure 6 compares Stemming (S) and Lemmatisation (L) after Tokenisation (T) and Normalisation (N) have been applied.
-Most data points above the L = S line in precision and F1 score indicate that L is better than S in syntactic OM (assuming that post hoc correction is excluded).
+The shape of the distribution remains unchanged after N, which also indicates that R and S/L do not help OM.
+Figure 6: Comparison of text preprocessing pipeline: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
+In the violin plots, three horizontal lines inside each violin plot show three quartiles: Q1, median, and Q3 (an artefact of the Seaborn [47] package).
+Fig.
+7 compares Stemming (S) and Lemmatisation (L) after Tokenisation (T) and Normalisation (N) have been applied.
+Most data points above the L = S line in precision and F1 score indicate that L is better than S in OM (assuming that post hoc correction is excluded).
 In the violin plots: (a) Precision: The median after L is greater than that achieved by S.
@@
 The shape of the distribution is identical.
-    Figure 6. Stemming (S) vs Lemmatisation (L).
-[…]
-    Stemming (S) vs Lemmatisation (L).
-  2. (2)
-    Figure 7 compares different stemming methods: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL).
+Figure 7: Comparison of Stemming (S) and Lemmatisation (L).
+Fig.
+8 compares different stemming methods: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL).
 For SP and SS, all data points located on the equivalence line indicate that there is no difference between SP and SS.
@@
 The shape of the distribution is identical in SP, SS, and SL.
-    Figure 7. Comparison of different stemmers: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL).
-[…]
-    Comparison of different stemmers.
-  3. (3)
-    Figure 8 compares Lemmatisation (L) and Lemmatisation + POS Tagging (LT).
-All data points located on the equivalence line indicate that using POS Tagging in Lemmatisation does not help syntactic OM.
+Figure 8: Comparison of different stemming methods: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL).
+Fig.
+9 compares Lemmatisation (L) and Lemmatisation + POS Tagging (LT).
+All data points located on the equivalence line indicate that using POS Tagging in Lemmatisation does not help OM.
 In the violin plots, for each of (a) Precision, (b) Recall, and (c) F1 Score, the median number and the shape of the distribution are identical for L and LT.
-    Figure 8. Lemmatisation (L) without/with POS Tagging (LT).
-[…]
-    Lemmatisation (L) without/with POS Tagging (LT).
-4.3. Implications
-For syntactic OM, the text preprocessing pipeline can be categorised into two phases.
+Figure 9: Comparison of Lemmatisation (L) without/with POS Tagging (LT).
+IV-C Discussion
+For OM, the text preprocessing pipeline can be categorised into two phases.
 Phase 1 text preprocessing methods contain Normalisation and Tokenisation, whereas Phase 2 preprocessing methods consist of Stop Words Removal and Stemming/Lemmatisation.
@@
 In contrast, Phase 2 text preprocessing methods change words to better capture semantic similarity, such as removing prefixes and suffixes or substituting common etymological roots.
-  1. (1)
-    Phase 1 text preprocessing methods: We observe that matching performance usually increases with each Phase 1 text preprocessing method.
-However, Phase 1 methods could also benefit from customisation in OM tasks.
-Some traditional methods originating from natural language processing (NLP) are not useful for OM and need to be adjusted.
+For Phase 1 text preprocessing methods, we observe that matching performance usually increases with each method.
+In most cases, OM only requires minor preprocessing using Phase 1 text preprocessing methods.
+This is because entity names in the ontology are often compound words that do not occur in natural language, but they are partially formalised by agreement.
+There have been well-defined conventions established over the years, e.g.
+singularity, positive names, nouns for classes and verbs for properties [48, 49].
+However, Phase 1 text preprocessing methods could also benefit from customisation in OM tasks.
+Some traditional methods originating from natural language processing (NLP) are not applicable to OM and require adjustment.
 For example, sentence segmentation (i.e.
@@
 Word tokenisation (i.e.
-breaking text into single words) could be rewritten to detect the abbreviations that are common practice in ontology concept names and to tackle the conflicting use of Pascal case and snake case.
-  2. (2)
-    Phase 2 text preprocessing methods: We observe no benefit in using Phase 2 text preprocessing methods in syntactic OM.
+breaking text into single words) could be rewritten to detect abbreviations commonly used in ontology concept names and to tackle the conflicting use of Pascal case and snake case.
+For Phase 2 text preprocessing methods, we observe no benefit in using these methods in OM.
 In some cases, Stop Words Removal and Stemming/Lemmatisation may even hamper the matching performance.
@@
 On the other hand, S is more eager to find TPs, but eagerness can also lead to more FPs being found.
-Based on our experiments with no post hoc correction applied, L is better than S.
+Based on our experiments, L is superior to S without post hoc correction.
   3. (c)
-      Different stemming methods: Porter Stemmer (PS) and Snowball Stemmer (SS) have been found to perform better than Lancaster Stemmer (LS), and we cannot see a significant difference between Porter and Snowball.
-Although the Lancaster Stemmer is more eager to detect more TPs, it does not lead to performance improvement as more FPs are also matched.
+    Different stemming methods: Porter Stemmer (PS) and Snowball Stemmer (SS) have been found to perform better than Lancaster Stemmer (LS), and there is no significant difference between PS and SS.
+Although LS is more eager to detect more TPs, it does not lead to performance improvement as more FPs are also matched.
   4. (d)
@@
 The reason may be that ontology classes are usually nouns or gerunds, and in such cases, we could expect the simpler grammatical assumption to have similar results.
-  3. (3)
-    In most cases, OM only requires minor preprocessing using Phase 1 text preprocessing methods.
-This is because entity names in the ontology are often compound words that do not occur in natural language, but they are partially formalised by agreement.
-There have been well-defined conventions established over the years, e.g.
-singularity, positive names, nouns for classes and verbs for properties (Schober et al., 2007; Taylor et al., 2015).
-5. Text Preprocessing Pipeline Repair
-The experimental results in Section 4 demonstrate that only Phase 1 text preprocessing (Tokenisation and Normalisation) helps with syntactic OM.
+V Text Preprocessing Pipeline Repair
+The experimental results in Section IV demonstrate that only Phase 1 text preprocessing (Tokenisation and Normalisation) helps with OM.
 The use of Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation) does not improve performance and may even have negative impacts.
-In this section, we explore the underlying cause of this behaviour and introduce our novel pipeline repair approaches.
-5.1. Causal Analysis
-  1. (1)
+In this section, we explore the underlying cause of this behaviour and propose our novel pipeline repair approaches.
+V-A Causal Analysis
 Phase 1 text preprocessing methods (Tokenisation and Normalisation) do not change the text meaning.
@@
 This means that applying Phase 1 methods only help to detect TPs, while the number of FPs remains unchanged.
-For this reason, Phase 1 methods always have a positive effect on precision, recall, and overall F1 score.
-    (1) $Prec\uparrow=\frac{|A\cap R|}{|A|}=\frac{TP}{TP+FP}=1-\frac{FP}{TP\uparrow+FP}$
-    (2) $Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$
-    (3) $F_{1}\ Score\uparrow=\frac{2}{Prec\uparrow^{-1}+Rec\uparrow^{-1}}$
-  2. (2)
+For this reason, Phase 1 methods always have a positive effect on precision, recall, and overall F1 score (Equations 12, and 3).
+$Prec\uparrow=\frac{|A\cap R|}{|A|}=\frac{TP}{TP+FP}=1-\frac{FP}{TP\uparrow+FP}$ (1)
+$Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$ (2)
+$F_{1}\ Score\uparrow=\frac{2}{Prec\uparrow^{-1}+Rec\uparrow^{-1}}$ (3)
 Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are actually relaxations of matching rules in OM tasks.
@@
 For example, isReviewing and isReviewedBy may be object properties with distinctly different meanings, but removing the common stop words “is” and “by”, and using Stemming/Lemmatisation to retrieve the same root word “Review”, could cause a false match.
-For this reason, recall is always increasing, but precision and the overall F1 score are less reliable.
-    (4) $Prec\ ?=\frac{|A\cap R|}{|A|}=\frac{TP\uparrow}{TP\uparrow+FP\uparrow}$
-    (5) $Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$
-    (6) $F_{1}\ Score\ ?=\frac{2}{Prec\ ?^{-1}+Rec\uparrow^{-1}}$
-    If we define $\Delta TP$ and $\Delta FP$ as the increase in TP and FP for a preprocessing method, then the condition to increase precision and F1 score is:
-    (7) $\frac{TP+\Delta TP}{TP+\Delta TP+FP+\Delta FP}>\frac{TP}{TP+FP}\Rightarrow\frac{\Delta TP}{\Delta FP}>\frac{TP}{FP}$
+For this reason, recall is always increasing (Equation 5), but precision and the overall F1 score are less reliable (Equations 4 and 6).
+[…]
+$Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$ (5)
+$F_{1}\ Score\ ?=\frac{2}{Prec\ ?^{-1}+Rec\uparrow^{-1}}$ (6)
+If we define $\Delta TP$ and $\Delta FP$ as the increase in TP and FP for a preprocessing method, then Equation 7 presents the condition to increase precision and F1 score.
+$\frac{TP+\Delta TP}{TP+\Delta TP+FP+\Delta FP}>\frac{TP}{TP+FP}\Rightarrow\frac{\Delta TP}{\Delta FP}>\frac{TP}{FP}$ (7)
 In our experiments, we actually observe a reduction in precision and overall F1 score.
@@
 This idea inspires our novel pipeline repair method.
-5.2. Motivation
-While Phase 2 text preprocessing methods have been shown to be less effective in syntactic OM by introducing FPs, we propose a pipeline repair approach to differentiate FPs from TPs.
-As illustrated in Figure 9: (1) Phase 1 methods shift Alignment (A) towards Reference (R).
+V-B Motivation
+While Phase 2 text preprocessing methods have been shown to be less effective in OM by introducing FPs, we propose a pipeline repair approach to differentiate FPs from TPs.
+As illustrated in Fig.
+10: (1) Phase 1 methods shift Alignment (A) towards Reference (R).
 The number of TPs increases, while the number of FPs decreases.
@@
 The number of FPs decreases significantly, with a slight decrease in TPs.
-Figure 9. Motivation of our context-based pipeline repair.
-[…]
-Motivation of our context-based pipeline repair.
-5.3. Ad Hoc Logic-Based Repair
-The ad hoc logic-based repair applies before the text preprocessing.
+Figure 10: Motivation of our context-based pipeline repair.
+V-C Ad Hoc Logic-Based Repair
+The ad hoc logic-based repair is applied before text preprocessing.
 One critical step in our approach is to assemble a set of reserved words that may cause FPs in text preprocessing, and to exclude these words from text preprocessing.
-The selection criteria are based on two ontology features: (1) entity names (class or property) are not duplicated within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies (see Section 3).
+The selection criteria are based on two ontology features: (1) entity names (class or property) are not duplicated within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies (see Section II).
 This approach is context-based because pairs of words may have the same or different meanings in different contexts.
@@
 0: Reserved_Word_Set $S_{r}$
-  /* Phase 1: Find duplicates in $O_{s}$ and $O_{t}$/
-  /* Find duplicates in $O_{s}$/
+/* Phase 1: Find duplicates in $O_{s}$ and $O_{t}$ */
+ /* Find duplicates in $O_{s}$ */
 for $Entity\ e_{i},e_{j}\in O_{s}$ do
   if $f(e_{i})=f(e_{j})$ then
-        $S_{r}=S_{r}\cup(S_{e}\div S_{j})$
+   $S_{r}=S_{r}\cup(S_{i}\div S_{j})$
   end if
 end for
-  /* Find duplicates in $O_{t}$/
+ /* Find duplicates in $O_{t}$ */
  /* Same procedure applies …*/
-  /* Phase 2: Find immutable words in $S$/
+/* Phase 2: Find immutable words in $S$ */
 for $Word\ w\in S$ do
@@
 return Reserved_Word_Set $S_{r}$
-An example of generating and using a simple reserved word set is illustrated below.
-(1) The two object properties was_a_member_of and has_members from a single ontology have the same result “member” via traditional text preprocessing.
+For example, consider two object properties was_a_member_of and has_members from a single ontology that have the same result “member” via traditional text preprocessing methods.
 Because there are no duplicate entities within a single ontology, we use a reserved word set in our proposed pipeline repair approach to determine that they are distinguishing entities.
@@
 Therefore, the final reserved word set is [“was”, “a”, “of”, “has”, “members”].
-(2) The generated reserved word set can be used to repair false mappings between entities within the same domain context but coming from different ontologies.
+The generated reserved word set can be used to repair false mappings between entities within the same domain context but coming from different ontologies.
 For example, we expect that the two object properties has_a_steering_committee and was_a_steering_committee_of are non-identical.
@@
 As the words “has”, “a”, “was”, and “of” are listed as reserved words, these two named properties are processed as “has a steer committe” and “was a steer committe of”, respectively.
-5.4. Post Hoc LLM-Based Repair
-With strong background knowledge, LLMs can be helpful in detecting false mappings after text preprocessing.
-However, there is a wide range of LLMs with different prompting strategies that can be used in LLM-based repair.
-We aim to find a generic pipeline repair approach and the best practice to integrate LLMs in the repair phase.
-The experiment settings are as follows.
-  1. (1)
-    We run the experiment on the Largebio Track that contains pairwise alignments between FMA, NCI, and SNOMED. We choose three datasets: WHOLE-FMA-NCI, WHOLE-FMA-SNOMED, and WHOLE-SNOMED-NCI.
-  2. (2)
+We apply our context-based pipeline repair approach to the same OAEI datasets as above.
+Fig.
+11 compares with and without context-based pipeline repair in 8 tracks with 49 distinct alignments.
+The text preprocessing pipeline methods implemented include less effective Phase 2 text preprocessing methods: Stop Words Removal (R), Porter Stemmer (SP), Snowball Stemmer (SS), Lancaster Stemmer (SL), Lemmatisation (L), and Lemmatisation + POS Tagging (LT).
+For precision, most data points are above the equivalence line, indicating that context-based pipeline repair significantly improves matching correctness.
+For recall, most data points are located on the equivalence line, and only a few data points are below the equivalence line, indicating that context-based pipeline repair slightly reduces matching completeness.
+Looking at F1 score, most of the data points are above the equivalence line, indicating that context-based pipeline repair also improves overall matching performance.
+Fig.
+11 also shows the relative improvement in matching performance made by Phase 2 preprocessing methods.
+This can be seen from the distance of data points from the equivalence line.
+We see that the matching performance improvement has Stemming (S) > Lemmatisation (L) > Stop Words Removal (R).
+Further, the matching performance improvement with different stemmers has Lancaster Stemmer (SL) > Porter Stemmer (SP) = Snowball Stemmer (SS) and the matching performance improvement for lemmatisation (L) alone is very similar to Lemmatisation + POS Tagging (LT).
+Figure 11: Context-based pipeline repair use in Phase 2 text preprocessing: Stop Words Removal (R), Porter Stemmer (SP), Snowball Stemmer (SS), Lancaster Stemmer (SL), Lemmatisation (L), and Lemmatisation + POS Tagging (LT).
+Fig.
+12 evaluates the scalability of ad hoc logic-based repair.
+The x-axis indicates the number of entities in each alignment, and the y-axis shows the number of reserved words per 100 entities.
+For all the alignments we tested, the value of reserved words per 100 entities is less than 10.
+With the number of entities increasing (i.e.
+the ontology size increases), the number of reserved words per 100 entities decreases.
+This may be because reserved words such as“has”, “is”, and “a” are commonly used in many ontologies and in multiple distinct entities within those ontologies.
+Conventionally, there are a limited number of such words, and therefore the reserved word feature of our approach is scalable in large-scale OM.
+Figure 12: Reserved words per 100 entities compared to the number of entities.
+V-D Post Hoc LLM-Based Repair
+In this study, we aim to develop a generic pipeline repair approach and establish best practices for integrating LLMs in the repair phase by empirical analysis with selected OAEI tracks.
+LLMs vary with respect to prompting strategies that could be used in LLM-based repair.
 We choose 10 LLMs from 4 different families.
-These include two OpenAI GPT models (OpenAI, [n.
-d.]) (gpt-4o and gpt-4o-mini), two Anthropic Claude models (Anthropic, [n.
-d.])(claude-3-sonnet and claude-3-haiku), and two versions of Meta Llama (Meta, [n.
-d.]) (llama-3-8b and llama-3.1-8b), two versions of Alibaba Qwen (Alibaba Qwen Team, [n.
-d.]) (qwen-2-7b and qwen-2.5-7b), Google Gemma 2 (Google Gemma Team and Google DeepMind, [n.
-d.]), and ChatGLM 4 (Zeng et al., 2024).
-We use LangChain (LangChain, Inc., [n.
-d.]) to build chats with OpenAI GPT and Anthropic Claude models, and Ollama (Ollama, [n.
-d.]) to access open-source models.
-Table 3 lists the details of selected LLMs.
+These include two OpenAI GPT models [50] (gpt-4o and gpt-4o-mini), two Anthropic Claude models [51] (claude-3-sonnet and claude-3-haiku), and two versions of Meta Llama [52] (llama-3-8b and llama-3.1-8b), two versions of Alibaba Qwen [53] (qwen-2-7b and qwen-2.5-7b), Google Gemma 2 [54], and ChatGLM 4 [55].
+We use LangChain [56] to build chats with OpenAI GPT and Anthropic Claude models, and Ollama [57] to access open-source models.
+Table II lists the details of selected LLMs (retrieved from [14]).
 All model temperatures are set to 0 to minimise the random results generated by LLMs.
-    Table 3. Details of selected LLMs (retrieved from (Qiang et al., 2023)). API-accessed models are shown as triangles ($\blacktriangle$) and open-source models as circles ($\bullet$).
+TABLE II: Details of selected LLMs. API-accessed models are shown as triangles ($\blacktriangle$) and open-source models as circles ($\bullet$).
 Family Model Version GPT $\blacktriangle$ gpt-4o gpt-4o-2024-05-13 $\blacktriangle$ gpt-4o-mini gpt-4o-mini-2024-07-18 Claude $\blacktriangle$ claude-3-sonnet claude-3-sonnet-20240229 $\blacktriangle$ claude-3-haiku claude-3-haiku-20240307 Llama $\bullet$  llama-3-8b Ollama Model ID: 365c0bd3c000 $\bullet$  llama-3.1-8b Ollama Model ID: 46e0c10c039e Qwen $\bullet$  qwen-2-7b Ollama Model ID: dd314f039b9d $\bullet$  qwen-2.5-7b Ollama Model ID: 845dbda0ea48 Gemma $\bullet$  gemma-2-9b Ollama Model ID: ff02c3702f32 GLM $\bullet$  glm-4-9b Ollama Model ID: 5b699761eca5
-  3. (3)
 We select 4 different LLM prompt strategies.
@@
 (d) Few-Shot + Self-Reflection: A combination of (b) and (c).
-Table 4 lists the details of the selected prompt strategies and their templates.
-    Table 4. Selected prompt templates. The few-shot parts are coloured blue, and the self-reflection parts are orange.
+Table III lists the details of the selected prompt strategies and their templates.
+TABLE III: Selected prompt templates. The few-shot parts are coloured blue, and the self-reflection parts are orange.
 No.
-Prompt Strategy Prompt Template PT1 Basic Is Entity1 equivalent to Entity2?
+Prompt Strategy Prompt Template PT1 Zero-Shot Is Entity1 equivalent to Entity2?
 Answer yes or no.
@@
 Self-Reflection Write a short explanation.
-  4. (4)
+We experiment with the Largebio Track that contains pairwise alignments between FMA, NCI, and SNOMED.
+We choose three datasets: WHOLE-FMA-NCI, WHOLE-FMA-SNOMED, and WHOLE-SNOMED-NCI.
 We compare LLMs with the classical text preprocessing pipeline in detecting TPs and FPs, which are two main contributors to precision and recall in matching performance.
-The true mappings (TPs) are retrieved from the reference (R) where the mappings can be found by syntactic OM.
+The true mappings (TPs) are retrieved from the reference (R) where the mappings can be found by OM.
 The false mappings (FPs) are retrieved from the alignment (A) generated by the classical text preprocessing pipeline but exclude TPs above.
-  5. (5)
-    We report our results over a single run.
-There could be slight differences in the results over multiple trials, but these differences are not significant with respect to the ranking and trend of different LLMs and prompt templates.
-Figure 10(l) compares LLMs with the classical text preprocessing pipeline in detecting TPs.
+In this study, we report our results on a single run.
+There may be slight differences in the results across multiple trials, but these differences are not significant in terms of the ranking and trend of different LLMs and prompt templates.
+Fig.
+13 compares LLMs with the classical text preprocessing pipeline in detecting TPs.
 LLMs and the text preprocessing pipeline are expected to answer “yes” for TPs.
@@
 For the text preprocessing pipeline, two entities are considered to be matched if they have the same results after the text preprocessing.
-Here we use only Tokenisation (T) and Normalisation (N) to preprocess the text because Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are shown to be less effective in our analysis in Section 4.
+Here we use only Tokenisation (T) and Normalisation (N) to preprocess the text because Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are shown to be less effective in our analysis in Section IV.
 Although some API-accessed commercial LLMs have shown the capability to detect TPs without a text preprocessing step, this behaviour varies across different LLMs and prompt templates.
@@
 The combination of few-shot and self-reflection (PT4) may even perform worse than the simple basic prompt (PT1).
-We believe that this phenomenon is mainly attributed to “overthinking”, a type of LLM hallucination in which LLMs could over-analyse user input and generate unnecessarily complex details (Chen et al., 2025).
+We believe that this phenomenon is mainly attributed to “overthinking”, a type of LLM hallucination in which LLMs could over-analyse user input and generate unnecessarily complex details [58].
 We also need to note here that the example generated in the few-shot prompt is based on our observation of the existing TPs.
@@
 On the other hand, using the classical text preprocessing pipeline via function calling could help LLMs produce a stable result with a high discovery rate (almost 100% from our observation) across different models and prompt templates.
-Figure 10. Discovery rate of TPs using LLMs on the Largebio Track.
-[…]
-Discovery rate of TPs using LLMs on the Largebio Track.
-Figure 11. The discovery rate of FPs using LLMs on the Largebio Track.
-[…]
-Discovery rate of FPs using LLMs on the Largebio Track.
-Figure 11(l) compares LLMs with the classical text preprocessing pipeline in detecting FPs.
+(a) PT1: WHOLE-FMA-NCI
+(b) PT1: WHOLE-FMA-SNOMED
+(c) PT1: WHOLE-SNOMED-NCI
+(d) PT2: WHOLE-FMA-NCI
+(e) PT2: WHOLE-FMA-SNOMED
+(f) PT2: WHOLE-SNOMED-NCI
+(g) PT3: WHOLE-FMA-NCI
+(h) PT3: WHOLE-FMA-SNOMED
+(i) PT3: WHOLE-SNOMED-NCI
+(j) PT4: WHOLE-FMA-NCI
+(k) PT4: WHOLE-FMA-SNOMED
+(l) PT4: WHOLE-SNOMED-NCI
+Figure 13: Discovery rate of TPs using LLMs on the Largebio Track.
+Fig.
+14 compares LLMs with the classical text preprocessing pipeline in detecting FPs.
 LLMs and the text preprocessing pipeline are expected to answer “no” for FPs.
@@
 While mapping repair is usually non-trivial and requires domain expertise, LLMs have shown potential for performing a high-quality post hoc correction without human intervention.
-For syntactic OM, we can see that LLMs are not good at detecting true mappings, but they are good at detecting false mappings. Therefore, we propose our generic post hoc LLM-based repair with
-  1. (1)
-    Employ the classical text preprocessing pipeline via function calling (a.k.a.
+(a) PT1: WHOLE-FMA-NCI
+(b) PT1: WHOLE-FMA-SNOMED
+(c) PT1: WHOLE-SNOMED-NCI
+(d) PT2: WHOLE-FMA-NCI
+(e) PT2: WHOLE-FMA-SNOMED
+(f) PT2: WHOLE-SNOMED-NCI
+(g) PT3: WHOLE-FMA-NCI
+(h) PT3: WHOLE-FMA-SNOMED
+(i) PT3: WHOLE-SNOMED-NCI
+(j) PT4: WHOLE-FMA-NCI
+(k) PT4: WHOLE-FMA-SNOMED
+(l) PT4: WHOLE-SNOMED-NCI
+Figure 14: The discovery rate of FPs using LLMs on the Largebio Track.
+For OM, we can see that LLMs are not good at detecting true mappings, but they are good at detecting false mappings.
+Therefore, we propose our post hoc LLM-based repair with two steps.
+(1) Given an alignment, employ the classical text preprocessing pipeline on each mapping via an LLM function calling (a.k.a.
 tool calling) to detect TPs.
-This means that the LLMs act as an interface, and the mapping judgement is based on the programme logic defined in the text preprocessing method.
+This means that the LLMs act as an interface, and the mapping judgement is based on the program logic defined in the text preprocessing method.
 This provides stable results for text preprocessing when finding FPs.
 LLM prompting is not suitable for this task, even with complex prompting strategies.
-  2. (2)
-    Use LLM prompting to detect FPs.
-The prompt template can be simple like “Is A equivalent to B?”, as we cannot see significant improvement using complex prompting strategies (except Meta Llama models).
-5.5. Evaluation
-We apply our pipeline repair approaches on the Phase 2 text preprocessing method Lancaster Stemmer (SL), and evaluate their performance on the three large datasets in the Largebio Track: WHOLE-FMA-NCI, WHOLE-FMA-SNOMED, and WHOLE-SNOMED-NCI.
-Figure 12(c) compares the matching performance with/without using the pipeline approaches: ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of the above (CR).
-We can see that ad hoc logic-based repair (AR) and post hoc LLM-based repair (PR) can significantly improve precision, with only a slight decrease in recall, but the overall F1 score also increases.
-Although there is no restriction on the use of both ad hoc and post hoc repairs, the combination does not improve precision, recall, or F1 score.
-For syntactic OM, its main purpose is to provide “seed mappings” for lexical and semantic matching.
+(2) Given an alignment, use LLM prompting on each mapping to detect FPs.
+The prompt template can be simple like “Is A equivalent to B?”, as we cannot see significant improvement using complex prompting strategies (except for Meta Llama models).
+We apply our pipeline repair approaches on the Phase 2 text preprocessing method Lancaster Stemmer (SL).
+Due to the superior performance and superior cost of the best LLMs, we evaluate only on the three largest datasets in the Largebio Track: WHOLE-FMA-NCI, WHOLE-FMA-SNOMED, and WHOLE-SNOMED-NCI.
+With this choice, we expect the performance increase may justify the cost of the repair step.
+Fig.
+15 compares post hoc LLM-based repair (PR) with two alternatives: text-preprocessing-only repair (TOR) and LLM-only repair (LOR).
+The results show that PR achieves almost the same precision as TOR and LOR, but its recall and F1 score are better than the other two alternatives.
+For OM, this means that our approach could detect more ”seed mappings”.
+Fig.
+16 compares the matching performance with/without using the pipeline repair approaches: ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of the above (CR).
+We can see that AR and PR can significantly improve precision, with only a slight decrease in recall, but the overall F1 score also increases.
+For text preprocessing used in OM, its main purpose is to provide “seed mappings” for lexical and semantic matching.
 These mappings are meant to be precise and accurate.
-Both of our pipeline repair approaches have shown high precision and are the best tool suites.
-(a) WHOLE-FMA-NCI
-(b) WHOLE-FMA-SNOMED
-(c) WHOLE-SNOMED-NCI
-Figure 12.
-Evaluation of pipeline repair approaches: Ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of above (CR).
+Both of our pipeline repair approaches shown with high precision are the best tool suites for OM.
+Although there is no restriction on the use of both ad hoc and post hoc repairs, the combination CR does not improve precision, recall, or F1 score.
+Figure 15: Comparison of post hoc LLM-based repair (PR) with two alternatives: text-preprocessing-only repair (TOR) and LLM-only repair (LOR).
+Figure 16: Evaluation of pipeline repair approaches: Ad hoc logic-based repair (AR), post hoc LLM-based repair (PR), and the combination of above (CR).
 The symbol $\varnothing$ represents the baseline without text preprocessing repair.
-[…]
-Evaluation of pipeline repair approaches.
-Figure 13 evaluates the scalability of ad hoc logic-based repair (AR).
-The x-axis indicates the number of entities in each alignment, and the y-axis shows the number of reserved words per 100 entities.
-For all the alignments we tested, the value of reserved words per 100 entities is less than 10.
-With the number of entities increasing (i.e.
-the ontology size increases), the number of reserved words per 100 entities decreases.
-This may be because reserved words such as“has”, “is”, and “a” are commonly used in many ontologies and in multiple distinct entities within those ontologies.
-Conventionally there are a limited number of such words.
-Therefore the reserved word feature of our approach is scalable in large-scale OM.
-Figure 13. Reserved words per 100 entities vs the number of entities in each alignment.
-[…]
-Comparison of reserved words per 100 entities (@100) with the number of entities in each alignment.
-Figure 14(c) compares post hoc LLM-based repair (PR) with two alternative repair approaches.
-The results show that PR achieves almost the same precision with text preprocessing only repair (TOR) and LLM only repair (LOR), but its recall and F1 score are better than the other two alternatives.
-For syntactic OM, this means that our approach could detect more ”seed mappings”.
-(a) WHOLE-FMA-NCI
-(b) WHOLE-FMA-SNOMED
-(c) WHOLE-SNOMED-NCI
-Figure 14. Comparison of post hoc LLM-based repair (PR) with two alternative repair approaches: text preprocessing only repair (TOR) and LLM only repair (LOR).
-[…]
-Comparison of post hoc LLM-based repair (PR) with two alternative repair approaches.
-6. Limitations
-  1. (1)
-    The references contained in the OAEI tracks are not perfect gold standards and some need further development. Therefore at present 100% accuracy is not necessarily desirable.
-  2. (2)
-    In this study we deal only with equivalence mappings between classes and properties.
+VI Limitations
+The references contained in the OAEI tracks are not perfect gold standards and some need further development.
+Therefore at present 100% accuracy is not necessarily desirable.
+In this study, we deal only with equivalence mappings between classes and properties.
 Subsumption mappings are not considered.
@@
 Only a few OAEI tracks contain subsumption mappings, and they are not enough to robustly demonstrate the text preprocessing used in subsumption mappings.
-  3. (3)
-    This work has not been included in the official OAEI campaign because of its intense focus on text preprocessing in syntactic OM.
-7. Conclusion
-In this paper, we conduct a comprehensive study on the effect of the text preprocessing pipeline on syntactic OM.
+VII Conclusion
+In this paper, we conduct a comprehensive study on the effect of the text preprocessing pipeline on OM.
 8 OAEI tracks with 49 distinct alignments are evaluated.
-Despite the importance of text preprocessing in syntactic OM, our experimental results indicate that the text preprocessing pipeline is currently ill-equipped.
-(1) We find that Phase 1 text preprocessing methods (Tokenisation and Normalisation) help with both matching completeness (i.e.
+Despite the importance of text preprocessing in OM, our experimental results indicate that the text preprocessing pipeline is currently ill-equipped.
+We find that Phase 1 text preprocessing methods (Tokenisation and Normalisation) help with both matching completeness (i.e.
 recall) and correctness (i.e.
@@
 precision) is relatively low.
-(2) We propose two novel pipeline repair methods to repair the less effective Phase 2 text preprocessings.
+We propose two novel pipeline repair methods to repair the less effective Phase 2 text preprocessings.
 Both approaches outperform the traditional text preprocessing pipeline without repair, in particular, the matching correctness (i.e.
@@
 F1 score).
-While the classical text preprocessing pipeline in syntactic OM is usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
-  1. (1)
+We argue that the traditional text preprocessing pipeline is not obsolete in the era of LLMs.
+It is an integral part of the LLMs used in OM tasks.
+Integrating the classical text preprocessing pipeline can significantly improve the robustness of purely prompt-based approaches.
+While the text preprocessing pipeline in OM is usually applied on the basis of intuition or extrapolation from other experiences, this work advances the state of knowledge towards making design decisions objective and supported by evidence.
   • Our study shows (a) whether to use, or not use, and (b) how to use these text preprocessing methods.
-Such experimental results will benefit the selection of appropriate text preprocessing methods for syntactic OM.
+Such experimental results will benefit the selection of appropriate text preprocessing methods for OM.
 For large-scale OM, it can significantly reduce unnecessary trial costs.
-  2. (2)
   • Our pipeline repair approaches can be used to repair the less effective Phase 2 text preprocessing methods.
 Its broader value is showing how to maximise true mappings and minimise false mappings throughout the text preprocessing pipeline.
-From a statistical perspective, it is a local optimisation for syntactic matching that will benefit the global optimisation for OM, as syntactic matching provides “anchor mappings” for lexical and semantic matching.
-  3. (3)
-    We showcase the limitations and opportunities for integrating the text preprocessing pipeline with LLMs in syntactic OM.
+From a statistical perspective, it is a local optimisation for syntactic matching that will benefit the global optimisation for OM.
+  • We showcase the opportunities for integrating the text preprocessing pipeline with LLMs in OM.
 We find that LLMs are good tools to find false mappings, but their outputs for finding true mappings are non-deterministic.
@@
 Using function calling to invoke the classical text preprocessing pipeline within LLMs, our approach could produce more stable and robust “seed mappings” than traditional text preprocessing methods and purely prompt-based approaches.
-  4. (4)
-    Our study also demonstrates that the nature of text preprocessing used in NLP applications can be context-based.
-Not all pipeline methods are effective, and some of them may even hamper the performance of downstream tasks.
-Modifications are required before applying text preprocessing to specific NLP tasks.
+  • Our study demonstrates that the nature of text preprocessing used in downstream applications can be context-based.
+Not all methods are effective, and some of them may even hamper the performance of downstream tasks.
+Modifications are required before applying text preprocessing methods to specific tasks.
 Our future work will focus on handling class axioms and complex relationships to evaluate the text preprocessing pipeline.
 We will also study the pipeline and pipeline repair approach working with both knowledge-based OM systems and LLM-based OM systems.
-Acknowledgements.
+This work has been included in the Agent-OM entry to the 2025 OAEI campaign.
+Artifacts
+The source code, data, and/or other artifacts have been made available at https://github.com/qzc438-research/ontology-nlp.
+The OAEI datasets are available from the OAEI Matching EvaLuation Toolkit (MELT) at https://dwslab.github.io/melt/track-repository (retrieved January 1, 2025).
+The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.
+Acknowledgment
 The authors thank Sven Hertling for curating the Ontology Alignment Evaluation Initiative (OAEI) datasets.
@@
 The authors thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.
-Artifacts
-The source code is available from Github at https://github.com/qzc438-research/ontology-nlp.
-The OAEI datasets are available from the OAEI Matching EvaLuation Toolkit (MELT) at https://dwslab.github.io/melt/track-repository (retrieved January 1, 2025).
-The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.
-Appendix A Missing References in the LargeBio Track
-As mentioned in the Limitations Section, the OAEI tracks are undergoing continuous development and may be missing some reference alignments.
-Here we list the missing reference alignments found by our text preprocessing pipeline in the folder “csvs”.
-These mappings would appear to be correct and we treat these mappings as TPs in our study.